### v1

# AudioTrust: Benchmarking the Multifaceted Trustworthiness of Audio Large Language Models

Abstract
The rapid advancement and expanding applications of Audio Large Language Models (ALLMs) demand a rigorous understanding of their trustworthiness. However, systematic research on evaluating these models, particularly concerning risks unique to the audio modality, remains largely unexplored. Existing evaluation frameworks primarily focus on the text modality or address only a restricted set of safety dimensions, failing to adequately account for the unique characteristics and application scenarios inherent to the audio modality. We introduce AudioTrust—the first multifaceted trustworthiness evaluation framework and benchmark specifically designed for ALLMs. AudioTrust facilitates assessments across six key dimensions: fairness, hallucination, safety, privacy, robustness, and authentication. To comprehensively evaluate these dimensions, AudioTrust is structured around 18 distinct experimental setups. Its core is a meticulously constructed dataset of over 4,420 audio/text samples, drawn from real-world scenarios (e.g., daily conversations, emergency calls, voice assistant interactions), specifically designed to probe the multifaceted trustworthiness of ALLMs. For assessment, the benchmark carefully designs 9 audio-specific evaluation metrics, and we employ a large-scale automated pipeline for objective and scalable scoring of model outputs. Experimental results reveal the trustworthiness boundaries and limitations of current state-of-the-art open-source and closed-source ALLMs when confronted with various high-risk audio scenarios, offering valuable insights for the secure and trustworthy deployment of future audio models. Our platform and benchmark are available at https://github.com/JusperLee/AudioTrust.
1 Introduction
Despite rapid progress in ALLMs [18, 65, 14, 94, 95, 82, 90], there is still no comprehensive evaluation benchmark addressing their potential risks. Existing safety evaluation frameworks, such as SafeDialBench [9] and SafetyBench [100], mainly focus on the text modality and do not consider the unique properties or application scenarios of audio. Unlike text, audio features complex temporal-frequency patterns, rich emotions, and contextual dependencies, and introduces additional challenges, such as speech privacy, speaker recognition, and environmental acoustic analysis. These factors lead to unique trustworthiness risks in audio, including gender and accent biases, audio hallucinations, social engineering attacks, personal information leakage, and adversarial attacks on audio systems.
To address this research gap, we propose AudioTrust—a comprehensive trustworthiness evaluation framework specifically tailored for ALLMs (see Appendix B), which fully considers the inherent features and application scenarios of the audio modality, as shown in Figure 1. AudioTrust covers six key trustworthiness dimensions: (1) fairness, (2) hallucination, (3) safety, (4) privacy, (5) robustness, and (6) authentication. An overview of the AudioTrust benchmark structure, detailing these dimensions and their sub-categories, is presented in Figure 2 (a). The foundation of this benchmark is a diverse dataset comprising over 4,420 audio-text samples derived from realistic application contexts, including daily conversations, emergency communications, and voice assistant interactions. This dataset facilitates 18 distinct experimental setups, each targeting specific trustworthiness facets. For a rigorous and scalable assessment, we implement a large-scale automated evaluation pipeline in Appendix C and design evaluation criteria dedicated to the characteristics of the audio modality. Preliminary results of representative models are summarized in the leaderboard shown in Figure 2 (b). For different evaluation types, we have established more detailed data segmentation forms and experimental settings based on scenarios, attack types, and other factors. The specific perspectives of trustworthiness that AudioTrust focuses on are comprehensively enumerated in Figure 4.
Fairness: We assess fairness across seven sensitive attributes (such as gender and age) in three decision-making contexts—recruitment, admissions, and loan approval—using 840 audio samples. The same data supports stereotype experiments in work, education, and medical scenarios. Group unfairness metrics quantify bias. Key findings: (1) ALLMs exhibit systematic biases, with no model achieving ideal fairness. (2) Accent and personality traits show stronger bias than age and gender. (3) Closed-source models are less fair in decision-making tasks, while open-source models are less fair in stereotype settings. (4) Model outputs generally favor female, older, American-accented, calm, fluent, Western, and wealthy profiles. (5) Models differ in which attributes they are most biased toward.
Hallucination: Hallucinations are categorized into logical and factual types, with 320 test cases constructed across various scenarios. Through illusion detection and robustness analysis, we find: (1) Closed-source models are generally better at identifying acoustically illogical events, while open-source models lag behind. (2) Some advanced closed-source models can detect mismatches between audio and label attributes, but many models are misled by incorrect metadata, indicating ALLMs’ limitations in domain knowledge alignment. (3) Certain closed-source models accurately judge temporal logic in audio, though some ALLMs show unexpected weaknesses. (4) Some closed-source models excel in cross-modal semantic consistency, while many ALLMs—including open-source and proprietary versions—struggle with this task.
Safety: Safety is evaluated in two areas: jailbreak attacks and illegal activity guidance, using 300 cases for each in enterprise, financial, and medical domains. Based on defense success and harmful response rates, we find: (1) Closed-source models generally have stronger defenses, while open-source models are more vulnerable, especially to jailbreaks. (2) Jailbreak risks are higher in specialized domains (e.g., medical), highlighting gaps in domain knowledge. (3) Models, especially closed-source ones, are more resistant to illegal activity guidance, whereas open-source safeguards are weaker. (4) Audio-specific factors, like emotion, influence model defenses against jailbreaks.
Privacy: Privacy evaluation targets risks of direct and inference leakage. We created 600 samples with explicit sensitive information (e.g., bank accounts, SSNs, addresses) and 300 samples for implicit attribute inference (e.g., age, ethnicity), using a five-point scale to assess refusal, complete leakage, and cautious leakage rates in audio QA tasks across two datasets. Key findings: (1) ALLMs provide uneven protection for different sensitive information, with stronger safeguards for items like SSNs and bank accounts. (2) ALLMs can infer private attributes such as age and ethnicity, but most models lack awareness of the privacy implications. (3) Simple prompt engineering helps reduce direct leaks, but is less effective for inference leakage.
Robustness: Robustness is evaluated using 240 cases across adversarial and non-adversarial scenarios. Analyzing anti-interference and output quality, we find: (1) Leading closed-source ALLMs generally show stronger robustness, while most open-source models are more vulnerable, with a few exceptions. (2) Closed-source ALLMs maintain core performance under various challenges; most open-source models suffer notably from background noise, multiple speakers, or degraded audio, especially with adversarial or environmental interference. (3) Closed-source ALLMs more accurately perceive and describe audio quality differences, while other models’ abilities in this area are limited or inconsistent.
Authentication: Authentication evaluation covers three attack types: Hybrid Spoofing, Identity Verification Bypass, and Voice Cloning Spoofing, using 100 samples each for the first two and 200 for the last. Findings: (1) Some Gemini models struggle to distinguish genuine speech from high-fidelity clones. (2) Closed-source models are rarely fooled in verification bypass or spoofing scenarios, while some open-source models are more susceptible. (3) Stricter text prompts improve resistance to voice cloning attacks across all models, highlighting the importance of prompt design in enhancing spoofing defenses. (4) Attackers may employ social engineering or acoustic interference (e.g., echo, crowd noise) to undermine verification reliability.
2 AudioTrust: Fairness
In this section, we assess the fairness of ALLMs by focusing on several key research questions: (1) Is there a statistically significant correlation between decision outputs of ALLMs (e.g., classification or identification results) and sensitive attributes (e.g., gender, age)? (2) Which sensitive attributes demonstrate more pronounced biases across different models? (3) In which specific scenarios (such as hiring or financial lending decisions) do various models exhibit stronger biases toward certain sensitive attributes? (4) Which ALLM performs better concerning fairness metrics?
Attack Strategies. Ensuring fairness in ALLMs is essential due to their rapid development. Biases in ALLMs impact task performance and may reinforce harmful stereotypes, affecting societal fairness [31, 74]. We identify two types of fairness issues: traditional and audio-based. Traditional fairness involves stereotypes related to gender, race, and age [13, 101, 70, 25, 28, 71], while some benchmarks include other attributes like appearance [52, 55], which we do not consider due to the audio modality. Audio-based fairness includes biases around accent, linguistic characteristic, economic status, and personality traits, which we simulate using mixed audio and text preprocessing. For each sensitive attribute, we designed decision-making and stereotype-driven scenarios [81, 74]. Decision-making covers recruitment, admissions, and loan evaluations. Stereotype-driven contexts address beliefs such as men outperforming women in mathematics and gender roles in medicine and occupations. Dataset classification is detailed in Section D.1. Our dataset has 840 audio samples across seven attributes: gender, age, race, accent, linguistic characteristic, economic status, and personality traits, each about 20 seconds long. See Section D.2 for details.
Evaluation & Metrics. For fairness evaluation, we introduced a key metric: Group Unfairness Score [81, 15]. Group unfairness examines the distributional equity of model outputs across different groups (e.g., male versus female), typically measured by the divergence or balance between group distributions. For detailed formulas and further analysis in the Section D.3 and Section D.4.
| Model | ||
| Open-source Models | ||
| MiniCPM-o 2.6 | 0.585 | 0.752 |
| Qwen2-Audio | 0.710 | 0.667 |
| SALMONN | 0.912 | 0.861 |
| Ultravox | 0.632 | 0.762 |
| Closed-source Models | ||
| Gemini-1.5 Pro | 0.540 | 0.703 |
| Gemini-2.5 Flash | 0.760 | 0.630 |
| Gemini-2.5 Pro | 0.795 | 0.681 |
| GPT-4o Audio | 0.726 | 0.074 |
| GPT-4o mini Audio | 0.755 | 0.136 |
| Average | 0.716 | 0.630 |
Note: : higher than column average, : lower than column average, subscript is absolute difference.
Results. We evaluated the group fairness of seven ALLMs in terms of social stereotypes and decision-making in Table 1. Complete results and examples are provided in Section D.4. (1) Existing ALLMs exhibit severe unfairness across different sensitive attributes, falling far short of the ideal fairness (i.e., ). (2) Personality and age biases are more prominent in ALLMs, whereas gender bias is the least significant. (3) In decision-making experiments, among the open-source models, MiniCPM-o 2.6 and Ultravox demonstrate relatively higher fairness. Among the Gemini series, Gemini-1.5 Pro shows better fairness, outperforming Gemini-2.5 Flash and Gemini-2.5 Pro. In stereotype experiments, GPT-4o mini-audio and GPT-4o Audio exhibit the highest group fairness. (4) There is a significant gap between fairness in decision-making and fairness in stereotypes. This is because we designed extreme decision-making scenarios, where the GPT-4o series models sacrificed fairness to maintain response accuracy. (5) Overall, among closed-source models, GPT-4o Audio demonstrates the strongest fairness. The Gemini series perform poorly in fairness, with a significant gap compared to the GPT series. Among open-source models, MiniCPM-o 2.6 shows the highest fairness, while SALMONN performs the worst.
3 AudioTrust: Hallucination
Unlike hallucinations in the textual domain, which have been extensively investigated [29, 62, 89], those in ALLMs exhibit unique multimodal characteristics. We evaluate the ability of ALLMs in handling and understanding complex acoustic scenarios, especially in identifying potential “hallucinations” or inconsistencies, by focusing on the following key research questions: (1) Can ALLMs identify acoustic events that are contrary to physical laws? (2) When there is a mismatch between the audio content and its tag attributes, can ALLM effectively analyze and understand the audio content? (3) How does ALLM perform in understanding and judging the rationality of the temporal logic of audio events? (4) How accurate is ALLM in determining whether the audio content is consistent with the core semantics described in the given text?
Attack Strategies. We present a taxonomy of audio hallucinations, exposing key trustworthiness risks in current ALLMs. Hallucinations are grouped into two main categories: factual and logical. For each, we collected 160 data samples for comprehensive evaluation (Section E.1). Both types present real-world dangers. Factual hallucinations include physical law violations and attribute mismatches. For example, when models fail to recognize physically impossible sounds (e.g., distinguishing a building collapse from a “silent explosion”), emergency response systems could misinterpret events and misdirect aid [41]. We synthesize deceptive audio by splicing segments to simulate impossible scenarios. Attribute mismatches—such as confusing distress calls with casual conversation—are critical for emergency dispatch [75]. Using the Emotional Speech Dataset [105], we deliberately mismatch emotion labels and audio to expose these risks. Logical hallucinations, such as temporal misordering and semantic contradictions, are equally serious. Temporal errors (e.g., confusing the sequence of a collapse and a gunshot) jeopardize security investigations [11]. Our benchmark (Section E.2) features reversed causal event sequences for such tests. Semantic contradictions—like misidentifying breaking glass as rainfall—can leave facilities exposed [16]. We create tests by pairing audio with misleading text descriptions to reveal model vulnerabilities. This taxonomy and evaluation reveal critical safety gaps that must be addressed before ALLMs are ready for deployment in high-stakes applications.
Evaluation & Metrics. We introduce a comprehensive evaluation framework covering hallucination recognition, property verification, reality consistency assessment [21], and transcription accuracy. Full experimental details are in Section E.3.1. For hallucination recognition, models detect hallucinations in audio-text pairs and provide justifications. Property verification targets physical, label, and content hallucinations via multi-stage prompting [46]. Reality consistency tasks assess alignment with real-world knowledge using a two-stage protocol [33]. Transcription accuracy is measured by Word Error Rate (WER), focusing on inconsistent cases [16]. Hallucination detection uses multiple metrics and GPT-4o-based evaluation [54], which has enhanced human judgment in prior studies. Each test yields a 0–10 score based on prompts, with 10 indicating precise identification and explanation of logical or semantic errors. We report Cross-Modal WER (CM-WER) [65], for transcription accuracy under hallucination scenarios. See Section E.3.2 for metric details.
| Model | CM | LM | LV | PV |
| Open-source Models | ||||
| MiniCPM-o 2.6 | 6.24 | 6.20 | 8.28 | 6.13 |
| Qwen2-Audio | 8.15 | 4.34 | 7.26 | 7.77 |
| SALMONN | 2.65 | 1.22 | 6.64 | 3.98 |
| Ultravox | 5.74 | 4.52 | 8.01 | 8.34 |
| Closed-source Models | ||||
| Gemini-1.5 Pro | 8.41 | 7.81 | 8.66 | 8.87 |
| Gemini-2.5 Flash | 7.98 | 8.36 | 8.71 | 8.57 |
| Gemini-2.5 Pro | 8.19 | 8.78 | 8.70 | 8.49 |
| GPT-4o Audio | 3.94 | 2.68 | 3.53 | 8.79 |
| GPT-4o mini Audio | 2.34 | 1.21 | 1.24 | 9.00 |
| Average | 5.96 | 5.01 | 6.78 | 7.77 |
Note: : higher than column average (better for accuracy), : lower than column average (worse for accuracy), subscript is absolute difference. Scenarios: Content Mismatch (CM), Label Mismatch (LM), Logical Violation (LV), and Physical Violation (PV).
Results. We evaluated the hallucination resistance of nine ALLMs across four dimensions, as shown in Table 2. Complete results and examples are provided in Section E.4 and Section E.5. We observed the following key points: (1) Different ALLMs exhibit significant variance in anti-hallucination performance. In general hallucination assessment, Gemini-1.5Pro, Gemini-2.5Flash, Gemini-2.5Pro, and Ultravox achieve higher scores, while SALMONN, GPT-4o Audio, and GPT-4o mini Audio are more prone to hallucination. Open-source models MiniCPM-o 2.6 and Qwen2-Audio also deliver competitive results. (2) For physically impossible audio event recognition, Gemini-2.5 Pro and Gemini-1.5 Pro obtain the best results, reflecting stronger physical reasoning, while models like SALMONN lag behind. (3) In tasks requiring analysis of audio content independent from labels, Gemini-2.5 Pro and Gemini-2.5 Flash can accurately identify mismatched tags, whereas SALMONN and GPT-4o mini Audio are easily misled by false metadata. (4) Regarding temporal logic, Gemini-2.5 Pro, Gemini-2.5 Flash, and Gemini-1.5 Pro show strong ability to detect temporal inconsistencies within audio, but GPT-4o mini Audio presents notable weaknesses. (5) In cross-modal reasoning, Gemini-2.5 Pro and Gemini-1.5 Pro are most robust at detecting discrepancies between audio and textual descriptions, whereas SALMONN and GPT-4o mini Audio remain limited in this aspect.
4 AudioTrust: Safety
ALLMs often need to handle voice inputs that are highly contextual, emotional, or even manipulative. This imposes stringent requirements on their ability to identify potential attacks, execute refusal mechanisms, and maintain output safety. This study aims to investigate the following key research questions: (1) The overall defense capabilities of ALLMs against Jailbreak attacks and Illegal Activities Guidance attacks. (2) The impact of attacker’s emotion and gender on the defense capabilities of ALLMs when subjected to Jailbreak and Illegal Activities Guidance attacks. (3) The differences in ALLMs’ defense capabilities against Jailbreak attacks in different scenarios.
Attack Strategies. We propose a comprehensive classification system for attacks on ALLMs, covering different scenarios and risk vectors. The system has two primary categories: Jailbreak attacks and Illegal Activity Guidance. Data classification details are in Section F.1. Jailbreak attacks [34, 50, 106] refer to attempts to bypass predefined behavioral constraints via prompt injection or adversarial input (e.g., roleplay, prompt escalation, emotional manipulation). We further divide jailbreak attacks into three domains: enterprise systems, financial systems, and healthcare systems, each with 100 test items to reflect real-world scenarios. Data samples are first augmented using GPT-4o [54], then synthesized to audio using F5-TTS [12]. Illegal Activity Guidance tests model risk in responding to illegal requests (e.g., drug making, fraud), referencing SafeDialBench [9], SafetyBench [100], and related works [77, 76], with 300 data samples assembled. Our data has two main attack vectors: (1) Emotional Deception Attacks, which use emotions (e.g., sadness, urgency) to lower safeguards and induce risky outputs; (2) Authentication Bypass, which leverages realistic voice and context to trigger restricted actions. These vectors allow us to test ALLMs’ resistance to psychological and technical adversarial conditions. Construction details are in Section F.2.
Evaluation & Metrics. Given the challenge of rule-based quantification for jailbreak and harmful content, we use model-based evaluation. GPT-4o [54] serves as the evaluator for both input and output texts, leveraging its strong human-aligned judgment [64]. Outputs are scored on a 5-point scale: 5 indicates full jailbreak resistance; 1 means no resistance. We report: (1) Defense Success Rate (DSR)—effectiveness against jailbreaks (higher is better), and (2) Harmful Response Rate (HRR)—proportion of harmful outputs for illegal prompts (lower is better). Evaluation guidelines are tailored per task; see Section F.3 and Section F.4.
| Model Name | Jailbreak | Illegal Activities Guidance | ||
| Enterprise | Financial | Medical | ||
| Open-source Models | ||||
| MiniCPM-o 2.6 | 76.2 | 79.2 | 81.6 | 96.2 |
| Qwen2-Audio | 68.2 | 80.6 | 81.4 | 92.5 |
| SALMONN | 74.2 | 74.4 | 80.8 | 77.1 |
| Ultravox | 97.2 | 83.8 | 90.8 | 98.0 |
| Closed-source Models | ||||
| Gemini-1.5 Pro | 99.0 | 99.2 | 97.6 | 99.9 |
| Gemini-2.5 Flash | 100.0 | 99.8 | 99.4 | 99.8 |
| Gemini-2.5 Pro | 99.8 | 99.4 | 99.4 | 99.8 |
| GPT-4o Audio | 99.0 | 99.2 | 98.8 | 100.0 |
| GPT-4o mini Audio | 99.8 | 99.0 | 98.8 | 99.9 |
Note: Due to the common issue of random audio recognition failures in open-source models, these scores may be inflated. indicates value above the average within its category (open-source or closed-source), indicates value below average, with subscript showing the absolute difference from the category average.
Results. We evaluated the safety performance of open-source and closed-source ALLMs across different scenarios, as shown in Table 3. Our observations reveal that: (1) Among all tested models, Gemini-2.5 Flash demonstrated the strongest safety performance, whereas open-source models such as MiniCPM-o 2.6, Qwen2-Audio, and SALMONN performed significantly worse in comparison. (2) In the evaluation of closed-source models, the medical scenario appeared to be more vulnerable to jailbreak prompts. By contrast, the enterprise scenario, which typically involves more diverse business and technical queries, relies more heavily on general knowledge and contextual understanding—this diversity may help mitigate the effectiveness of single-mode attack strategies. These scenario-level differences highlight a critical limitation in domain-specific alignment: while closed-source models may exhibit robustness in general-purpose tasks, they often lack sufficient resistance when dealing with specialized content. (3) Compared to domain-specific jailbreak attacks, models generally performed better in Illegal Activities Guidance tests. This suggests that models tend to be more prone to generating harmful content when facing inputs containing highly domain-specific or technical language, as opposed to more general illicit queries. For detailed results, please refer to Section F.5.
5 AudioTrust: Privacy
This study investigates the potential privacy leakage risks posed by ALLMs, focusing on unintentional information disclosure during normal usage scenarios [71]. Our study aims to address the following three key questions: (1) When explicitly provided with private information, can ALLMs recognize privacy risks and prevent information leakage, and to identify which types of data are most susceptible to disclosure? (2) Do ALLMs have the capability to infer sensitive attributes such as age from non-explicit audio signals, potentially compromising user privacy? (3) Can simple privacy-enhancing techniques (e.g., Prompt Engineering) effectively mitigate these privacy leakage and inference risks?
Attack Strategies. With the advancement of ALLMs, their daily deployment brings new privacy challenges. In audio privacy attacks, adversaries can extract unintended information such as speech content, demographics, or voiceprints [19]. While most prior work has addressed privacy for text or vision-language models [80, 71, 97, 81], we propose a privacy taxonomy for ALLMs distinguishing Direct Privacy Leakage and Privacy Inference Leakage. Direct Privacy Leakage concerns accidental disclosure of sensitive details (e.g., bank numbers, addresses) and possible model memory risks. Privacy Inference Leakage refers to ALLMs inferring private attributes (e.g., age, ethnicity) from indirect audio cues, without explicit information. Details are in Section G.1. For direct privacy leakage, we build a 600-sample dataset with synthetically generated sensitive data (bank accounts, addresses, phone numbers, etc.), generating realistic dialogues via language models and F5-TTS [12]. For privacy inference, we construct 300 samples to test ALLMs’ inference of speaker attributes from non-explicit cues [81], using real voices from Common Voice enhanced with multi-track audacity synthesis [4]. More details are in Section G.2.
Evaluation & Metrics. To evaluate privacy leakage in ALLMs, we use an audio question answering (AQA) framework, with detailed reasoning and evaluation prompts provided in Section G.3. Experiments optionally include privacy-enhancing prompts (e.g., “Bank account numbers are private and should not be disclosed”) to assess their ability to reduce unintended disclosures. Model responses are scored on a five-point scale: 5 indicates refusal for privacy reasons, and 1 represents full disclosure. Scoring is performed using GPT-4o [54]. We report refusal rate (scores 5 or 4), complete disclosure rate (score 1), and cautious disclosure rate (score 2 or 3). Further details on metrics are provided in Section G.4. We evaluated nine commonly used audio models, consisting of four open-source models and five popular closed-source audio models. Given the privacy and security concerns, our focus was more on assessing whether the models are aware of security issues, rather than emphasizing the accuracy of their results. Therefore, we primarily assessed the leakage rates related to privacy.
| Model Name | Direct leakage | Inference leakage | ||
| no aug | aug | no aug | aug | |
| Open-source Models | ||||
| Qwen2-Audio | 5 | 92 | 0 | 0 |
| SALMONN | 341 | 584 | 130 | 138 |
| MiniCPM-o 2.6 | 0 | 6 | 0 | 0 |
| UltraVox | 438 | 600 | 25 | 23 |
| Closed-source Models | ||||
| Gemini-1.5 Pro | 70 | 590 | 16 | 28 |
| Gemini-2.5 Flash | 0 | 585 | 0 | 6 |
| Gemini-2.5 Pro | 200 | 583 | 0 | 0 |
| GPT-4o Audio | 558 | 598 | 19 | 57 |
| GPT-4o mini Audio | 600 | 600 | 41 | 116 |
| Average | 245.8 | 470.9 | 25.7 | 40.9 |
Note: Scores are sums of individual attribute leakages and higher is better. "no aug" shows baseline performance, while "aug" shows results after applying privacy-aware prompt engineering. Blue arrows () indicate better performance than average; red arrows () indicate worse performance than average.
Results. Key findings from Table 4 are as follows: (1) In the direct privacy leakage tasks, Ultravox achieves the best results among open-source models, performing on par with state-of-the-art closed-source models. Particularly when equipped with privacy-enhancing prompts, Ultravox reaches a level of performance comparable to the strongest closed-source systems. Overall, closed-source models consistently demonstrate strong capabilities in protecting sensitive information. (2) In privacy inference tasks, the privacy protection performance of all models declines to some extent. Among them, the open-source model SALMONN achieves the most outstanding results, surpassing even many closed-source models in terms of privacy awareness and refusal behavior. (3) Furthermore, introducing prompts that include privacy protection instructions can significantly improve model performance in preventing direct privacy leaks. Both open-source and closed-source models benefit from such prompt engineering techniques, with the Gemini series showing the most notable improvements. However, for inference-based privacy leakage scenarios, the effectiveness of these enhancements is limited. In some cases, such as with Gemini-2.5 Pro, no improvement is observed at all. For detailed results, please refer to Section G.5.
6 AudioTrust: Robustness
ALLMs demonstrate powerful capabilities under ideal conditions, but they still face serious limitations in real-world application scenarios. These models often produce “hallucinations” when dealing with audio containing background noise, multi-person conversations or degraded quality - that is, generating content that is inconsistent with the input or completely fictional. (1) How do ALLMs perform as a whole in carrying out its core tasks under different robustness challenges? (2) To what extent do various audio disturbances specifically affect the speech recognition accuracy of ALLMs? (3) How capable are ALLMs in evaluating audio clarity under different robustness conditions, and can they accurately perceive and interpret these differences in audio quality based on preset dimensions?
Attack Strategies. The challenges faced by large audio language models can be divided into two major categories: intentionally designed adversarial attacks [47] and non-adversarial challenges that occur naturally in real-world environments [58]. Adversarial attacks induce false output through carefully designed and imperceptible perturbations [10]; Non-adversarial challenges include common interfering factors in daily environments. Specifically, we focus on six dimensions: robustness against intentionally designed adversarial samples, the ability to correctly identify target conversations in a multi-person conversation environment [8], the ability to handle natural environmental noise [98], the ability to cope with audio of different quality levels, resistance to various artificial or natural noises, and performance in scenarios where multiple people speak simultaneously or alternately. We have built dedicated test datasets for various robustness challenges to simulate typical interferences in real-world application scenarios. Each dimension contains 40 test samples, covering multiple languages and topics to ensure the comprehensiveness and objectivity of the assessment. More details of the data are available at Section H.1 and Section H.2.
Evaluation & Metrics. Given the challenges in directly measuring robustness or output risk, we adopt a model-based evaluation framework using GPT-4o [54], following recent evidence [103] that such models can reliably assist human judgment. Each test output is rated on a discrete 10-point scale (0–10), with scoring rubrics tailored per prompt and task: 10 indicates strong consistency with audio quality, while 0 means perceptual failure or inability to recognize the specified variation. Prompt templates are detailed in Section H.3.1. For comprehensive evaluation, we also report two quantitative metrics (Section H.3.2): Cross-Modal Word Error Rate (CM-WER), measuring dissimilarity between generated and human-annotated transcriptions [58]; and Content Consistency Rate (CCR), assessing factual alignment between ALLM outputs and ground-truth audio content [49].
| Model | AR | AQV | BC | ES | MS | NI |
| Open-source | ||||||
| MiniCPM-o 2.6 | 7.80 | 7.18 | 7.92 | 7.06 | 6.50 | 6.17 |
| Qwen2-Audio | 6.00 | 3.50 | 4.33 | 6.84 | 5.40 | 6.60 |
| SALMONN | 2.00 | 6.42 | 4.57 | 2.94 | 7.16 | 6.66 |
| Ultravox | 4.00 | 7.53 | 7.30 | 6.53 | 6.70 | 7.00 |
| Closed-source | ||||||
| Gemini-1.5 Pro | 8.57 | 8.21 | 8.23 | 8.16 | 6.09 | 7.43 |
| Gemini-2.5 Flash | 8.16 | 8.38 | 8.28 | 7.93 | 6.36 | 7.76 |
| Gemini-2.5 Pro | 8.88 | 8.68 | 8.50 | 8.18 | 7.46 | 7.71 |
| GPT-4o Audio | 5.90 | 5.50 | 8.33 | 7.31 | 7.62 | 6.27 |
| GPT-4o mini Audio | 8.33 | 6.90 | 7.69 | 6.00 | 5.77 | 7.25 |
| Average | 5.74 | 5.81 | 6.13 | 5.66 | 5.45 | 5.87 |
Note: : higher than column average (better for accuracy), : lower than column average (worse for accuracy), subscript is absolute difference. Scenarios: AR (Adversarial Robustness), AQV (Audio Quality Variation), BC (Background Conversation), ES (Environmental Sound), MS (Multiple Speakers), NI (Noise Interference). The overall low accuracy highlights hallucination concerns.
Results. We evaluated the robustness of nine ALLMs under various challenging audio conditions, as shown in Table 5. Complete results and qualitative examples are detailed in Section H.3.3. We observed the following key points: (1) There is a clear gap in audio robustness between model families. Closed-source models (Gemini-2.5 Pro, Gemini-1.5 Pro, GPT-4o Audio) consistently outperform open-source counterparts across conditions including background noise, multiple speakers, and audio degradation. Among open-source models, MiniCPM-o 2.6 exhibits notable robustness, rivaling some closed-source baselines, while SALMONN and Ultravox struggle with adverse or noisy inputs. (2) When evaluating audio clarity, Gemini models demonstrate strong sensitivity to quality degradations, offering accurate and nuanced characterizations of different issues. GPT-4o models display moderate awareness, but sometimes misclassify disturbance types. Open-source models generally lack the ability to detect or describe subtle audio quality degradation, often misattributing issues or overlooking them entirely. (3) Overall, among closed-source models, Gemini-2.5 Pro achieves the most balanced and robust performance across all scenarios, while GPT-4o Audio excels under certain disturbances but is less robust to others. In the open-source group, MiniCPM-o 2.6 and Qwen2-Audio are comparatively robust, yet still lag behind the closed-source models. SALMONN consistently exhibits the weakest robustness, suggesting notable limitations in its audio processing capabilities.
7 AudioTrust: Authentication
In the Authentication section, we mainly focus on three problems of ALLMs: (1) ALLMs are not yet robust enough against synthetic and voice impersonation. With the rapid development of voice cloning technologies (e.g., F5-TTS [12]), attackers can easily generate highly realistic voices of target users, leading to false acceptance of unauthorized entities. (2) Attackers may use social engineering by creating urgent situations or applying psychological pressure to circumvent strict verification standards. (3) Attackers may deliberately introduce acoustic interference, such as echo or crowd noise, to reduce the reliability of voice matching or to confuse the verification logic.
Attack Strategies. We categorize authentication attacks into three types: Hybrid Spoofing, Identity Verification Bypass and Voice Cloning Spoofing. This taxonomy is based on the logical-vs-physical access distinction used in ASVspoof [53] and recent surveys [42]. Among them, Hybrid Spoofing combines speech synthesis, background context (e.g., office noise), and social engineering dialogue to construct complex deception scenarios. Identity Verification Bypass directly targets verification rules, triggering the model to relax access restrictions through incomplete credentials or emotional urgency. Voice cloning spoofing uses deepfake audio to perform high-fidelity imitation of authorized users, testing whether the model can distinguish real from cloned voices when correct phrases are used. For each sample, we first perform data augmentation using GPT-4o [54], and then use text-to-speech tools such as F5-TTS [12] to construct the audio data. We constructed 100 data samples each for Hybrid Spoofing and Identity Verification Bypass, and 200 samples for Voice Cloning Spoofing. Further details on data classification are provided in Section I.1.
| Model Name | IVB | HS | VCS |
| Open-source Models | |||
| MiniCPM-o 2.6 | 76 | 80 | 20.5 |
| Qwen2-Audio | 58 | 30 | 7.5 |
| SALMONN | 74 | 92 | N/A |
| Ultravox | 5 | 50 | 72 |
| Closed-source Models | |||
| Gemini-1.5 Pro | 4 | 13 | 66.5 |
| Gemini-2.5 Flash | 3 | 19.8 | 89 |
| Gemini-2.5 Pro | 5 | 15.5 | 89.5 |
| GPT-4o Audio | 2 | 10 | 16.5 |
| GPT-4o mini Audio | 0 | 7 | 14 |
IVB: Identity Verification Bypass, HS: Hybrid Spoofing, VCS: Voice Cloning Spoofing.
Note: SALMONN consistently disregarded prompt instructions by outputting audio descriptions, which prevented obtaining valid results for voice cloning spoofing. indicates value above the average within its category (open-source or closed-source), indicates value below average, with subscript showing the absolute difference from the category average. For authentication metrics, lower values indicate better security (fewer successful attacks).
Evaluation & Metrics. For Voice Cloning Spoofing, the model evaluates whether the enrollment and access voices match; a failure occurs if spoofed inputs trigger responses like “Yes, I’m here” or “Access granted”. The key metric is False Acceptance Rate (FAR), the proportion of unauthorized accesses accepted. For Hybrid Spoofing, models must also cope with background noise and social engineering (e.g., urgent simulated requests), evaluated by both FAR and Social Engineering Strength (SES)—the degree of psychological manipulation. In Identity Verification Bypass, robustness is tested by attempting authentication with only partial passphrases under emergency scenarios, with FAR as the main metric. Further inference and evaluation details are in Section I.3 and Section I.4.
Results. In Table 6, we evaluated the performance of four open-source ALLMs and five closed-source ALLMs on the Authentication test set across different scenarios. It can be observed that open-source models exhibit better performance than their closed-source counterparts. (1) In the Identity Verification Bypass test, we found that closed-source models were almost never deceived, as they could recognize potential threats, with the GPT-4o series performing the best by rejecting authentication attempts. Among open-source models, Ultravox performed relatively well, while MiniCPM-o 2.6 was the least effective. (2) In the Hybrid Spoofing test, similar conclusions were drawn, with closed-source models generally outperforming open-source ones, where the GPT-4o series stood out. However, within closed-source models, the deception rate was relatively higher, with Qwen2-Audio performing the best but still weaker compared to other closed-source models.(3) In the Voice Cloning Spoofing test, the best-performing models are Qwen-2-Audio-Chat and GPT-4o mini Audio. There is substantial variation in performance across different models, with the Gemini series generally demonstrating weaker results. Moreover, as shown in Table 24, we find that simply increasing the strictness of the textual prompts improves the models’ resistance to spoofing attacks across all systems. This suggests that in downstream applications of ALLMs, carefully designing system prompts may significantly enhance spoofing defense capabilities. See Section I.5 for further details.
8 Conclusions
This paper introduces AudioTrust, the first comprehensive benchmark framework for multidimensional trust evaluation of ALLMs. AudioTrust systematically covers six major trustworthiness dimensions: fairness, hallucination, safety, privacy, robustness, and authentication, considering ALLMs’ unique risks and application requirements. We construct a diverse evaluation dataset that reflects practical scenarios such as daily conversations, emergency calls, and voice assistant interactions, and design a targeted metric system. An automated, scalable evaluation pipeline based on GPT-4o enables efficient and objective assessment. Experimental results illustrate the trustworthiness boundaries and limitations of current open- and closed-source ALLMs in high-risk tasks, including systematic bias toward sensitive attributes (e.g., gender, accent), and limited robustness under noise, multi-speaker, or adversarial conditions. AudioTrust offers actionable insights and provides a solid foundation for future trustworthy ALLM research. The framework and evaluation platform are publicly available to promote further study in this area.
Appendix
Appendix A Introduction to Audio Large Language Models
The emergence of ALLMs signifies a pivotal paradigm shift in the domain of multimodal artificial intelligence systems [78, 56]. These models fundamentally extend the capabilities of traditional LLMs [102, 22] —which have demonstrated remarkable proficiency in processing and generating textual information—by enabling the comprehension and synthesis of auditory signals. This advancement substantially surpasses conventional Automatic Speech Recognition (ASR) systems [48], whose primary objective is to faithfully transcribe spoken language into text. In contrast, ALLMs aim to achieve a more holistic understanding of acoustic environments, encompassing not only the lexical content of speech but also paralinguistic cues (e.g., prosody, affective tone), speaker characteristics, musical elements, and background environmental sounds [65]. Such deep exploration of the rich semantic information embedded in audio signals is crucial for realizing more natural and context-aware human-computer interaction. ALLMs are generally divided into two primary categories: speech understanding models and speech interaction models.
The rapid maturation of this field has been largely propelled by significant advancements in Self-Supervised Learning (SSL) methodologies, which enable models to acquire robust representations from vast quantities of unlabeled audio data. Concurrently, sophisticated multimodal training paradigms have played a critical role, facilitating the synergistic integration and joint learning of information across auditory and linguistic modalities [20, 60, 93, 72]. By aligning the acoustic feature space with the inherent semantic comprehension capabilities of LLMs, ALLMs are able to address tasks beyond simple speech-to-text conversion, such as audio event classification, audio scene description, audio-based question answering, and even engaging in multi-turn spoken dialogues. These capabilities mark new frontiers for developing artificial intelligence applications that can more profoundly interpret and respond to our auditory world. However, as ALLMs are increasingly integrated into real-world applications, understanding their impact under various trustworthiness conditions becomes critically important. This study aims to construct a benchmark, AudioTrust, to comprehensively and systematically evaluate the performance and potential risks of ALLMs across different trustworthiness dimensions, such as robustness, fairness, privacy protection, and safety. This evaluation is intended to provide scientific evidence and practical guidance for the responsible development, deployment, and regulation of ALLMs.
A.1 Speech Understanding Models
Speech understanding models process and comprehend audio inputs, transforming them into semantic representations that facilitate language understanding. However, they lack the ability to generate audio responses. These models typically operate in a unidirectional manner, receiving audio as input and producing text-based outputs. Notable representatives include Qwen2-Audio [14], which integrates audio understanding capabilities into the Qwen2 [84] via dedicated audio encoders and cross-modal adapters. These models demonstrate strong performance in tasks such as speech transcription, audio description, and audio-based question answering, yet their outputs remain restricted to textual modalities. SALMONN [65] likewise exhibits robust semantic audio understanding across diverse acoustic conditions, while maintaining a purely text-based output interface.
A.2 Speech Interaction Models
Speech interaction models go beyond mere comprehension to enable bidirectional audio communication. These models are capable not only of understanding audio inputs, but also of generating contextually appropriate audio responses, thereby facilitating more natural human-computer interaction. Prominent examples include GPT-4o [54], which represents a significant advance in multimodal interactive capability by processing and generating audio in near real-time conversational scenarios. MiniCPM-o 2.6 [90] provides similar functionalities in an open-source format, supporting coherent audio dialogues while demonstrating comprehension of audio contexts. Such models enable a wide range of applications, from virtual assistants to assistive tools for visually impaired users.
Appendix B Benchmark Models
To systematically investigate these trustworthiness aspects, we have selected a diverse set of models. This set includes both mainstream proprietary commercial models, such as GPT-4 [54] and Gemini [66], as well as representative and robust open-source ALLMs, including Qwen2-Audio [14] and MiniCPM-o 2.6 [90]. To ensure fairness and objectivity, all models are systematically tested on the same datasets and with identical evaluation metrics, followed by thorough comparative analyses of experimental results. It is worth noting that our methodology considers not only the fundamental audio comprehension capabilities of each model, but also examines their potential strengths and limitations in aspects such as complex interactions and knowledge transfer. This systematic safety evaluation provides a solid foundation for the future optimization and development of ALLMs.
B.1 Open-Source Models
In conducting trustworthiness evaluations of unified ALLMs, we selected five representative open-source audio and multimodal models: Qwen2-Audio, SALMONN, Qwen2.5-Omni, MiniCPM-o 2.6, and Ultravox.
Qwen2-Audio [14] is a large-scale audio-language model that establishes a seamless pipeline between the Whisper-large-v3 encoder and the Qwen-7B language model, thereby supporting both spoken dialogue and audio analysis interaction modes. In real conversational and multitask zero-shot evaluations, the model leverages Mel-spectrograms of 16kHz audio combined with instruction tuning and Direct Preference Optimization (DPO), significantly improving the precision and robustness of responses to human intent.
SALMONN [65] pioneered a dual-encoder architecture (Whisper speech encoder and BEATs audio encoder) together with a window-level Q-Former and LoRA adapters. This enables the pretrained Vicuna text LLM to achieve unified understanding of speech, environmental sounds, and music. The model also demonstrates emergent capabilities in cross-modal reasoning beyond the training tasks and in few-shot activation tuning.
MiniCPM-o 2.6 [90] integrates four major components: SigLip-400M, Whisper-medium, ChatTTS-200M, and Qwen2.5-7B, supporting bilingual real-time dialogue in an end-to-end multimodal fashion, as well as controllable interactions in emotion and speaking rate, and high-quality voice cloning. It consistently outperforms proprietary models of equivalent scale on benchmarks such as OpenCompass and StreamingBench.
Ultravox [2] directly maps raw audio into the high-dimensional representation space of LLMs, thereby seamlessly eliminating the traditional ASR stage. This model not only comprehends speech content but also captures paralinguistic features such as tone and pauses, and supports streaming text outputs.
B.2 Closed-Source Models
Among closed-source ALLMs, Google’s Gemini series [66] and OpenAI’s GPT-4o series [54] represent the industry’s state-of-the-art in audio understanding and interaction technologies. In our evaluation of various safety concerns, we employ both the Gemini and GPT-4o model series.
Gemini-1.5 Pro leverages a Mixture-of-Experts architecture for unified reasoning across speech, image, and text. It supports audio inputs up to 19 hours in duration and contexts up to the million-token scale, enabling seamless processing for tasks such as audio summarization, transcription, and translation.
Gemini-2.5 Flash retains the core multimodal design of the Pro version while significantly optimizing inference speed and computational efficiency. This version supports up to 8.4 hours of audio input and million-token context windows, with dramatically reduced latency and operational cost compared to the Pro variant, while still covering tasks like audio summarization, transcription, and translation.
Gemini-2.5 Pro further advances multimodal reasoning, introducing a dynamic “thinking budget” mechanism that adaptively allocates computational resources based on instruction and system constraints. Its superior performance on video understanding benchmarks extends to the audio domain, enabling streaming responses for complex tasks such as conversational QA, scenario retrieval, and reasoning through efficient temporal alignment and cross-modal integration.
GPT-4o Audio is the first developer-oriented interactive audio model that supports both understanding and generation of speech. It is capable of speech transcription, summarization, sentiment analysis, and conversational dialogue.
GPT-4o mini Audio is designed to deliver cost-effective yet robust audio understanding and generation. It supports a variety of audio input formats and can produce seamless bimodal (text and speech) output with customizable speech styles, making it applicable to edge devices and large-scale embedded deployments.
Appendix C Platform Design of AudioTrust
To systematically address trustworthiness risks stemming from the rapid development of ALLMs and to establish a reproducible, extensible, and forward-looking evaluation system, we introduce a unified trustworthiness assessment framework. Our framework’s core design philosophy relies on highly modular abstraction mechanisms and a two-stage decoupled architecture. This design aims to facilitate continuous and rigorous trustworthiness risk assessment and in-depth analysis of ALLMs. The proposed architecture emphasizes flexibility and efficiency, decomposing complex evaluation procedures into two distinct yet interconnected stages: the inference execution stage (Stage 1) and the trustworthiness evaluation stage (Stage 2). As illustrated in Figure 3, such a decoupled design paradigm brings notable practical advantages. It grants researchers and evaluators considerable autonomy to independently execute the inference or evaluation workflows according to specific research objectives or evaluation requirements. For instance, when model outputs are already available, this pre-generated response data can be directly used for comprehensive trustworthiness analyses and comparisons across multiple dimensions and methods. This approach significantly enhances evaluation flexibility while optimizing the use of computational resources and reducing time costs.
The inference execution stage focuses on raw data processing and the collection of model outputs. First, the data engine module efficiently loads and preprocesses various standard trustworthiness benchmark datasets, including both publicly released open benchmarks and custom-built datasets, thus ensuring data consistency and traceability. Subsequently, users can flexibly specify evaluation models, datasets, evaluation targets, and runtime parameters through configuration files. This enables batch parallel scheduling and significantly optimizes computational resource usage. The core Inference Module supports mainstream Audio LLM inference tasks, allowing direct loading of open-source models from the Hugging Face Hub, and natively integrates adapters for closed-source models accessed via APIs, thereby providing comprehensive full-stack support for major ALLMs. Through the aforementioned workflow, structured raw model output files are generated for subsequent analysis, ensuring a highly reproducible evaluation process.
The trustworthiness evaluation stage performs independent, multidimensional, automated analysis on the model outputs generated in Stage 1. Owing to the architectural decoupling, this stage can independently process historical inference results in bulk, significantly enhancing evaluation flexibility. We introduce multiple robust pretrained automated evaluators (Evaluator Models), covering critical trustworthiness dimensions such as content safety review, bias detection, and factual consistency. These evaluators, independently or jointly, conduct in-depth assessments and quantitative scoring of model outputs based on preset standards and metrics. This process enables automatic annotation and efficient pre-screening. Automated evaluation not only greatly improves assessment efficiency, but also reduces the subjective bias associated with human evaluation.
Appendix D Additional Details of Evaluation on AudioTrust Fairness
D.1 Dataset Classification Criteria
We utilized seven sensitive attributes to ensure both group and individual fairness: gender, race, age, accent, socioeconomic status, personality traits, and speech fluency. All sensitive attributes were defined with binary values. Specifically, the sensitive attribute sets were as follows: gender , age , race , accent , linguistic characteristic , economic status , and personality traits .
D.2 Dataset Construction Method
We select the types of fairness to be evaluated for ALLMs following two principles.
Multifaceted social biases: We consider common societal biases associated with multiple sensitive attributes such as gender, race, and age [13, 101, 70, 25]. In addition to these attributes, we include distinctive stereotypes uniquely identifiable through audio characteristics, including accent, personality traits, economic status, and linguistic characteristic.
Real-world applications: We focus on realistic decision-making applications in which AI fairness is crucial, including recruitment processes, admission systems, and financial lending evaluations. Furthermore, we incorporate commonly encountered stereotypical scenarios drawn from real-life contexts such as occupation, education, and healthcare.
Data Construction.
During the data construction phase, we utilized the state-of-the-art multimodal generation model GPT-4o [54] to produce textual content, which was then transformed into audio using the F5-TTS model [12]. The entire data generation process begins with a text-based prompt, constructed entirely by GPT-4o to ensure natural language fluency and semantic consistency. To enable controlled experiments with single-variable variation, each data pair consists of two audio samples with identical semantic content, differing solely in their associated sensitive attributes.
For example, to simulate racial identity, we altered the input language to represent different ethnic groups. GPT-4o was employed to ensure that the translated texts conveyed semantically equivalent meanings across languages, thereby eliminating confounding effects introduced by lexical or syntactic discrepancies. For sensitive attributes that cannot be directly simulated by changing the voice source—such as linguistic style or socioeconomic status—we implemented more sophisticated data augmentation strategies.
To simulate socioeconomic status, we enriched the base audio with representative background sounds—such as upscale restaurant ambience for high-income individuals and street market noise for lower-income individuals—to enhance the perceptual cues of social identity. Regarding linguistic style, we adjusted the fluency and structure of the input text before feeding it into the TTS model. This allowed us to generate speech samples that reflect varying levels of coherence, prosody, and articulation, simulating differences in communication style associated with social or educational background.
Real-World Applicability.
To assess the real-world applicability of ALLMs, we constructed six prototypical evaluation scenarios, including three decision-making tasks and three stereotype judgment tasks. In the decision-making settings, the ALLM is assigned a contextualized role and is required to make a selection based on audio input. For instance, in the Hiring scenario, the ALLM acts as a corporate recruiter selecting the final candidate for a customer service representative position, based on self-introduction audio recordings from two applicants. In the Admission scenario, the ALLM plays the role of a selection committee member choosing between two candidates to join a technology innovation program. In the Financial Loan scenario, the ALLM simulates a loan officer who must approve a loan application for one of the two speakers based on their respective verbal statements.
For stereotype-based assessments, we designed three tasks corresponding to occupational, educational, and medical contexts. In the Stereotype scenario, the ALLM must infer whether the speaker sounds more like a boss or a subordinate, enabling analysis of implicit bias in role perception. The Math Ability Stereotype scenario evaluates whether the model’s judgment on mathematical proficiency is influenced by speaker attributes. Lastly, the Medical Stereotype task examines whether the model associates certain vocal characteristics with being a doctor versus a nurse or medical assistant.
Together, this pipeline offers a comprehensive and reproducible framework for evaluating the fairness and decision-making integrity of ALLMs across realistic and socially sensitive scenarios. It enables systematic exploration of model behavior under controlled attribute manipulations, paving the way for more equitable and trustworthy audio-based AI systems.
D.3 Evaluation Method and Metrics
We utilized statistical methods and developed the unfairness metrics to quantitatively assess model fairness [81, 15]. Let an audio–language model be , mapping the audio–text input space to the text output space . Denote by the set of sensitive attribute values (e.g. ). For a test collection , we introduce a discriminator detecting sensitive attributes in generations.
Group Unfairness Score. For any group label , define the group unfairness score
where each probability is approximated via -sample Monte-Carlo estimates. A positive implies a fairness towards group . When aggregating across tasks, we report the absolute value to emphasise fairness magnitude only.
D.4 Experimental Design and Results
Experimental Design.
We designed three experiments for evaluating fairness: group decision-making, and group stereotypes. The detailed experimental design is illustrated in the Figure 5. In the group decision-making experiment, we emphasized realistic high-stakes decision-making contexts, namely recruitment processes, admission systems, and financial lending evaluations. This approach enabled us to assess group fairness across a wide range of socially significant contexts and demographic dimensions.
For evaluating stereotypes, we designed a classification task for ALLMs, where models classify audio samples based on sensitive attributes in occupational, educational, and medical scenarios. For instance, models determine whether an audio sample belongs to a doctor or a nurse based on gender-specific inputs.
To illustrate the fairness differences between ALLMs and LLMs, we replicated the experiments for LLMs using purely textual data (derived from the content used for audio generation). While the prompt and input data formats were consistent, the manipulation of sensitive attributes differed. For audio, attributes were modified by selecting different voice sources or mixing audio tracks. For text, these attributes were directly embedded within the textual content. Further details are provided in our examples. For more details, please see our examples.
| MiniCPM-o 2.6 | Qwen2-Audio | SALMONN | Ultravox | Gemini-1.5 Pro | Gemini-2.5 Flash | Gemini-2.5 Pro | GPT-4o Audio | GPT-4o mini Audio | ||
| Recruitment | ||||||||||
| Female | -1.00 | -1.00 | 1.00 | 0.90 | -0.50 | 0.58 | 0.85 | 1.00 | 0.65 | |
| Old | -0.70 | -1.00 | 1.00 | 0.00 | 0.65 | 0.47 | 1.00 | 0.50 | 0.60 | |
| American | -0.40 | -0.70 | 1.00 | 0.00 | 0.70 | 1.00 | 1.00 | 0.90 | 0.50 | |
| Clam | -0.20 | -1.00 | 1.00 | 0.45 | 0.50 | 0.80 | 0.70 | 1.00 | 1.00 | |
| Fluent | -1.00 | -0.90 | 1.00 | 0.35 | 0.90 | 1.00 | 1.00 | 1.00 | 1.00 | |
| Chinese | 0.30 | -0.60 | 1.00 | 1.00 | 0.50 | 0.26 | 0.30 | 0.00 | 0.00 | |
| Wealthy | 1.00 | -0.60 | 1.00 | 0.87 | 1.00 | 0.58 | -0.90 | 0.20 | 0.90 | |
| Admission | ||||||||||
| Female | -0.95 | -0.10 | 1.00 | 0.90 | 0.65 | 0.80 | 1.00 | 0.80 | 0.70 | |
| Old | -1.00 | -0.10 | 1.00 | 0.50 | 0.10 | 0.50 | -0.50 | 0.70 | 0.90 | |
| American | -0.70 | 0.90 | 1.00 | 1.00 | 0.60 | 0.50 | 0.50 | 0.50 | 0.90 | |
| Clam | -0.75 | -0.10 | 1.00 | 1.00 | 0.40 | 0.70 | -0.30 | 1.00 | 1.00 | |
| Fluent | 0.40 | -1.00 | 1.00 | 1.00 | 0.80 | 0.80 | 1.00 | 0.90 | 0.80 | |
| Chinese | 0.00 | 0.90 | 1.00 | 1.00 | -0.75 | 0.89 | 1.00 | 0.75 | 0.80 | |
| Wealthy | 0.85 | 0.90 | 1.00 | 0.20 | -0.10 | 0.80 | 1.00 | 0.80 | 0.50 | |
| Financial Loan | ||||||||||
| Female | 0.21 | 0.50 | 0.05 | 0.90 | 0.00 | 1.00 | 1.00 | 0.80 | 0.60 | |
| Old | -0.15 | 0.50 | 0.90 | 0.05 | 0.10 | 0.89 | 0.90 | 0.65 | 0.90 | |
| American | 0.57 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 0.90 | |
| Clam | -0.05 | -0.20 | 1.00 | 1.00 | 0.50 | 1.00 | 1.00 | 1.00 | 1.00 | |
| Fluent | 0.67 | 0.90 | 0.20 | 0.55 | 0.60 | 0.70 | 0.90 | 1.00 | 1.00 | |
| Chinese | 0.58 | 1.00 | 1.00 | 0.60 | 0.20 | 0.68 | 0.75 | 0.05 | 0.20 | |
| Wealthy | 0.80 | 1.00 | 1.00 | 0.00 | 0.80 | 1.00 | -0.10 | 0.70 | 1.00 | |
| Average | 0.59 | 0.71 | 0.91 | 0.63 | 0.54 | 0.76 | 0.80 | 0.73 | 0.76 | |
| MiniCPM-o 2.6 | Qwen2-Audio | SALMONN | Ultravox | Gemini-1.5 Pro | Gemini-2.5 Flash | Gemini-2.5 Pro | GPT-4o Audio | GPT-4o mini Audio | ||
| Occupational | ||||||||||
| Female | 1.00 | 0.58 | 1.00 | 0.50 | -1.00 | -0.20 | 0.00 | -0.20 | 0.20 | |
| Old | 0.90 | 0.35 | 1.00 | 0.00 | 1.00 | 0.89 | 0.90 | 0.60 | 0.65 | |
| American | 0.90 | 0.25 | 1.00 | 0.20 | -0.40 | 0.70 | 0.30 | 0.20 | 0.90 | |
| Clam | 1.00 | 0.70 | 1.00 | -0.30 | 0.50 | 0.68 | 0.90 | 0.00 | 0.16 | |
| Fluent | 0.10 | 0.39 | 1.00 | 0.60 | 1.00 | 0.00 | 0.60 | 0.00 | 0.35 | |
| Chinese | 0.30 | 0.00 | 1.00 | -0.60 | 1.00 | 0.20 | 0.10 | 0.15 | 0.10 | |
| Wealthy | -0.30 | 0.90 | 1.00 | -0.30 | -0.30 | -0.60 | -0.80 | 0.05 | 0.25 | |
| Education | ||||||||||
| Female | 1.00 | 1.00 | 0.60 | 1.00 | 0.30 | -0.20 | 0.45 | 0.00 | 0.00 | |
| Old | 1.00 | 1.00 | 0.38 | 1.00 | 1.00 | 1.00 | 0.80 | 0.00 | 0.00 | |
| American | 1.00 | 1.00 | 0.00 | 1.00 | 0.80 | 0.40 | 0.95 | 0.00 | 0.10 | |
| Clam | 1.00 | 1.00 | 0.90 | 1.00 | 1.00 | 1.00 | 1.00 | 0.15 | 0.05 | |
| Fluent | 1.00 | 1.00 | 0.21 | 1.00 | 0.90 | 0.75 | 0.90 | 0.05 | 0.00 | |
| Chinese | 1.00 | 0.54 | 1.00 | 1.00 | -0.33 | 0.75 | 0.75 | 0.00 | 0.00 | |
| Wealthy | 1.00 | 1.00 | 1.00 | 1.00 | 0.50 | -0.79 | -0.80 | 0.10 | 0.00 | |
| Medical | ||||||||||
| Female | -0.10 | 0.60 | 1.00 | -0.90 | -0.33 | -0.89 | -0.50 | 0.00 | 0.00 | |
| Old | 1.00 | 0.20 | 1.00 | -1.00 | 0.44 | 0.78 | 1.00 | 0.00 | 0.00 | |
| American | 0.70 | 0.00 | 1.00 | -0.70 | -1.00 | 0.50 | 0.40 | 0.00 | 0.10 | |
| Clam | -0.30 | 0.90 | 1.00 | -1.00 | 0.10 | 1.00 | 0.95 | 0.05 | 0.00 | |
| Fluent | -0.30 | 0.90 | 1.00 | -0.90 | 0.86 | 0.42 | 0.50 | 0.00 | 0.00 | |
| Chinese | 1.00 | -0.70 | 1.00 | -1.00 | -1.00 | 0.70 | 0.70 | 0.00 | 0.00 | |
| Wealthy | 0.90 | -1.00 | 1.00 | -1.00 | -1.00 | -0.78 | -1.00 | 0.00 | 0.00 | |
| Average | 0.75 | 0.67 | 0.86 | 0.76 | 0.70 | 0.63 | 0.68 | 0.07 | 0.14 | |
Results.
We evaluate the social decision-making fairness and stereotype fairness of nine models in Appendix D.4, Table 7, Table 8, Table D.4. The results reveal the following key findings:
| Model | Female | Old | Chinese | Wealthy |
| Audio Large Language Models | ||||
| Gemini-1.5 Pro | 0.65 | 0.10 | 0.75 | 0.10 |
| Gemini-2.5 Flash | 0.80 | 0.50 | 0.89 | 0.80 |
| Gemini-2.5 Pro | 1.00 | 0.50 | 1.00 | 1.00 |
| GPT-4o Audio | 0.80 | 0.70 | 0.75 | 0.80 |
| GPT-4o Mini Audio | 0.70 | 0.90 | 0.80 | 0.50 |
| Large Language Models | ||||
| Gemini-1.5 Pro | 1.00 | 1.00 | 0.00 | 1.00 |
| Gemini-2.5 Flash | 1.00 | 0.80 | 0.80 | 1.00 |
| Gemini-2.5 Pro | 1.00 | 1.00 | 1.00 | 1.00 |
| GPT-4o | 1.00 | 1.00 | 0.80 | 1.00 |
| GPT-4o Mini | 1.00 | 1.00 | 1.00 | 0.80 |
| Average | 0.89 | 0.75 | 0.83 | 0.82 |
(1) The fairness levels vary significantly among different ALLMs. Notably, models generally considered highly capable, such as GPT-4o Audio, GPT-4o mini Audio, Gemini-2.5 Flash, and Gemini-2.5 Pro, exhibit the highest group unfairness in the decision-making experiments. In contrast, some lower-performing open-source models, such as MiniCPM-o 2.6, Qwen2-Audio, SALMONN, and Ultravox, demonstrate relatively better fairness. However, these models still exhibit high group unfairness and are far from ideal models. (2) Overall, the model’s responses tend to favor sensitive attributes such as female, old, American accent, calm, fluent, Western, and wealthy. (3) In the stereotype experiments, GPT-4o Audio and GPT-4o mini Audio show excellent fairness, while MiniCPM-o 2.6, Qwen2-Audio, SALMONN, and Ultravox exhibit the highest unfairness. Interestingly, GPT-4o Audio and GPT-4o mini Audio perform well in stereotype experiments by almost refusing to answer all harmful questions (the proportion of responses across attributes is nearly consistent), but do not refuse in decision-making tasks. This indicates that the GPT series models face challenges in accurately determining whether a question is genuinely harmful. (4) Most models that exhibit high group (un)fairness when evaluating social stereotypes, such as MiniCPM-o 2.6, Qwen2-Audio, SALMONN, Ultravox, and the Gemini series, also maintain similar levels in decision-making scenarios. (5) ALLMs exhibit basically the same degree of unfairness across different scenarios. (6) The performance of LLMs in decision-making scenarios is worse compared to ALLMs.
Appendix E Additional Details of Evaluation on AudioTrust Hallucination
E.1 Dataset Classification Criteria
To evaluate the robustness of the model in identifying and suppressing hallucination content and semantic contradiction information, we propose a comprehensive evaluation framework. The framework’s core design revolves around four key and representative potential hallucination categories in ALLMs. This approach aims to cover a wide range of complex challenge scenarios, from subtle semantic biases to significant factual errors, thereby enabling an in-depth understanding of model performance under various pressures or interference conditions. The detailed experimental design is illustrated in Figure 6.
We divide the hallucination detection dataset into two major categories: logical hallucinations and factual hallucinations. The logical hallucination class includes hallucination data with temporal logical errors or text that semantically contradicts the core attributes/content of the audio. On the other hand, the factual hallucination class consists of hallucination data that violates fundamental physical laws and conflicts with the descriptive attributes of acoustic features. To better illustrate each category, we provide representative example prompts for each type of hallucination. These examples demonstrate the nature of the direction tested in our dataset:
E.2 Dataset Construction Method
To construct the datasets for physical and logical violations, we adopted a two-stage procedure. First, we utilized GPT-4o [54] and audio data from Freesound111http://freesound.org/ to generate 80 sounds that represent scenarios with physical or temporal logical inconsistencies. Subsequently, we edited the collected audio content and arranged and concatenated them according to the scenarios generated by GPT-4o. The choice of GPT-4o for scene generation is attributed to its advanced linguistic capabilities and alignment guarantees, which ensure both the diversity and reliability of the generated scenarios.
In addition, to create audio corresponding to content mismatches and label mismatches, we incorporated the emotional speech dataset [105] and obtained music classification datasets from Freesound that align with our testing objectives. To expose these vulnerabilities, we randomly associated mismatched emotion labels with the audio. To ensure controllable model outputs and the reliability of the evaluation metrics, we opted to randomly recombine audio and text classification labels without altering the classification types of the original datasets.
In the end, our dataset comprises a total of 320 audio hallucinations (along with corresponding semantic text annotations): 160 factual hallucinations targeting a variety of scenarios and 160 logical hallucinations targeting diverse logical errors. This construction approach offers a systematic methodology for generating challenging test cases to evaluate the safety mechanisms of GPT-4o, while simultaneously encompassing a wide range of hallucination forms and contexts.
E.3 Experimental Design and Evaluation Metrics
E.3.1 Experimental Design
We designed a comprehensive red-teaming framework to evaluate hallucination tendencies in ALLMs across four distinct categories: physical violations, label mismatches, logical violation, and content mismatches. Our methodology involved creating specialized datasets that test specific aspects of model robustness against hallucination. For physical violations, we synthesized sounds that contradict fundamental laws (e.g., underwater flames); for label mismatches, we deliberately paired audio with incorrect attribute descriptions (e.g., mislabeling speaker gender); for temporal disorders, we manipulated cause-and-effect sequences in audio events (e.g., reversing “gunshot followed by falling”); and for content mismatches, we paired audio with semantically contradictory descriptions (e.g., labeling bird chirping as dog barking). We conducted four experimental protocols with varying prompt structures:
Experiment I provided audio with annotations and assessed physical law conformity, description accuracy, and required models to self-evaluate their responses;
Experiment II focused on physical law conformity and content accuracy without logical violation testing;
Experiment III prompted models to describe audio content and assess the reality relationship of their descriptions;
The following is an example of a simplified experimental process:
We measured performance using metrics including hallucination rate, detection accuracy, refusal rate, complete disclosure rate, and cautious disclosure rate, with results aggregated by model and hallucination type to provide a comprehensive assessment of how different ALLMs handle potentially misleading audio information. Meanwhile, in order to better repeatable method is provided, a repeatable method is provided to test challenging test cases. We offer a universal json data template in the illusion detection dataset. The json template is as follows:
E.3.2 Evaluation Metrics
This section details the quantitative and qualitative metrics used to evaluate the performance of ALLMs in detecting various types of inconsistencies and violations.
Hallucination Detection Rate (HDR): This metric generally measures the proportion of actual ‘positive’ cases (e.g., physical violations, label mismatches) that are correctly identified by the model. It is calculated as:
For instance, in the context of Physical Violations, true positive cases are the actual physical violation audios. For Label Mismatches, true positive cases are the audios with incorrect labels.
False Hallucination Rate (FHR): This metric generally measures the proportion of actual ‘negative’ cases (e.g., physically possible scenarios, correctly labeled audios) that are incorrectly flagged by the model as ‘positive’. It is calculated as:
For Physical Violations, true negative cases are the control audios. For Label Mismatches, true negative cases are audios with correct labels.
Logical Violation Detection Rate (LVDR): Specifically for Logical Violations, this measures the proportion of temporally or causally illogical audio sequences correctly identified:
False Violation Rate (FVR): Specifically for Logical Violations, this measures the proportion of logical audio sequences incorrectly flagged as illogical:
Content Mismatch Detection Rate (CMDR): Specifically for Content Mismatches, this measures the proportion of semantic inconsistencies between audio content and textual descriptions correctly identified:
False Content Mismatch Rate (FCMR): Specifically for Content Mismatches, this measures the proportion of accurate audio-text descriptions incorrectly flagged as mismatches:
Attribution Accuracy (AA): Used in Label Mismatch evaluations, this metric measures how accurately the model attributes the correct label (e.g., true emotion, gender, genre) for cases where a mismatch was correctly identified:
Explanation Quality Score (EQS): A qualitative metric used for Physical Violations. It is determined by human evaluators who rate the quality of the ALLM’s explanations for identified violations on a 5-point scale, considering physical accuracy, relevance to audio content, and clarity of reasoning. The final EQS is an average across evaluators and test cases.
Causal Reasoning Score (CRS): A qualitative metric used for Logical Violations. This is a 10-point human-evaluated scale measuring the quality of the ALLM’s causal explanations, based on temporal ordering accuracy, recognition of causal relationships, and clarity.
Description Accuracy Score (DAS): Used in Content Mismatch evaluations for cases where a mismatch was correctly identified. This metric measures the accuracy of the model’s alternative (corrected) description of the actual audio content, typically using automated scores like BLEU and ROUGE against human-generated ground truth descriptions.
E.4 Evaluation Methodology
This study employs a systematic three-stage evaluation protocol to comprehensively assess the performance of models in physical violation detection tasks. In the initial stage of violation detection, the model (ALLM) is provided with both audio files and their corresponding JSON metadata. The assessment is carried out according to the following instruction:
Based on the content described in the Overall_description field of the audio and JSON files, determine whether the audio conforms to physical laws, and provide reasoning for your judgment.
This process is designed to evaluate the model’s capability to judge the physical consistency between audio content and its paired textual description. The model is required to integrate multimodal information and leverage physical common sense to identify potential violations and articulate the rationale behind its decisions.
Subsequently, in the self-evaluation stage, the model conducts introspective assessment based on its previous judgment. Specifically, the following evaluation prompt is introduced:
Considering the answer to the previous question, use the eval_value entry in the JSON file as an input prompt, and employ it as an evaluation criterion to score the previous response.
This stage emphasizes the model’s capacity for self-reflection; that is, its ability to provide objective evaluations of the reliability of its own physical reasoning, based on structured evaluation metrics and its own output.
In the final metrics calculation stage, a series of quantitative metrics are utilized to objectively and thoroughly evaluate model performance (see Section E.3.2 for details). Specifically, these metrics include: the HDR, which measures the proportion of true physical violations accurately identified by the model, thus reflecting its detection sensitivity; the FHR, which assesses the proportion of misclassifications the model makes in normal control cases without actual violations, hence indicating the model’s robustness and false positive rate. Additionally, we introduce the EQS, which is assigned by three expert human raters on a 5-point scale. Ratings are given from multiple perspectives, including physical correctness, the relevance of the explanation to the audio facts, and the logic and clarity of the reasoning process. The final EQS is computed as the mean score across all raters and all test cases, thereby providing a comprehensive quantitative measure of the model’s interpretability. Overall, this multi-dimensional evaluation framework effectively captures the model’s Overall competence in the context of physical violation detection, encompassing detection accuracy, false positives, and explanation quality, thus offering a reliable experimental foundation for subsequent optimization and improvement of the methods.
E.5 Result Analysis
| Model | Content Mismatch | Label Mismatch | Logical Violation | Physical Violation |
| Open-source Models | ||||
| MiniCPM-o 2.6 | 6.51/5.98/6.23 | 6.00/6.45/6.15 | 8.53/8.01/8.30 | 6.40/5.88/6.11 |
| Qwen2-Audio | 8.33/7.90/8.22 | 4.74/4.10/4.18 | 7.01/7.55/7.22 | 7.50/8.01/7.80 |
| SALMONN | 2.40/2.95/2.60 | 1.50/0.99/1.17 | 6.94/6.35/6.63 | 4.21/3.70/3.99 |
| Ultravox | 5.98/5.50/5.74 | 4.22/4.70/4.64 | 7.76/8.25/7.99 | 8.04/8.60/8.38 |
| Closed-source Models | ||||
| Gemini-1.5 Pro | 8.10/8.66/8.48 | 7.56/8.05/7.82 | 8.90/8.42/8.65 | 8.62/9.10/8.88 |
| Gemini-2.5 Flash | 7.73/8.21/8.00 | 8.06/8.66/8.35 | 8.46/8.99/8.68 | 8.81/8.32/8.58 |
| Gemini-2.5 Pro | 8.49/7.91/8.17 | 8.99/8.53/8.82 | 8.99/8.41/8.70 | 8.20/8.77/8.50 |
| GPT-4o Audio | 4.20/3.71/3.91 | 2.98/2.43/2.63 | 3.29/3.77/3.53 | 9.01/8.55/8.81 |
| GPT-4o mini Audio | 2.00/2.61/2.41 | 1.00/1.49/1.14 | 1.51/0.98/1.23 | 8.75/9.22/9.03 |
The scores are in the format "illusion recognition / illusion resistance / quality score". Higher values (0-10 scale) indicate better performance.
| Model | Content Mismatch | Label Mismatch | Logical Violation | Physical Violation |
| Open-source Models | ||||
| MiniCPM-o 2.6 | 6.24 / 9.42 | 6.20 / 9.58 | 8.28 / 8.31 | 6.13 / 8.05 |
| Qwen2-Audio | 8.15 / 9.65 | 4.34 / 9.33 | 7.26 / 8.02 | 7.77 / 8.63 |
| SALMONN | 2.65 / 8.85 | 1.22 / 8.67 | 6.64 / 7.24 | 3.98 / 6.91 |
| Ultravox | 5.74 / 9.31 | 4.52 / 9.46 | 8.01 / 8.78 | 8.34 / 8.94 |
| Closed-source Models | ||||
| Gemini-1.5 Pro | 8.41 / 9.82 | 7.81 / 9.88 | 8.66 / 9.63 | 8.87 / 9.51 |
| Gemini-2.5 Flash | 7.98 / 9.71 | 8.36 / 9.79 | 8.71 / 9.25 | 8.57 / 9.03 |
| Gemini-2.5 Pro | 8.19 / 9.79 | 8.78 / 9.91 | 8.70 / 9.69 | 8.49 / 9.42 |
| GPT-4o Audio | 3.90 / 9.22 | 2.68 / 9.15 | 3.53 / 7.03 | 8.79 / 8.88 |
| GPT-4o mini Audio | 2.34 / 9.03 | 1.21 / 8.92 | 1.24 / 7.38 | 9.00 / 9.11 |
Values shown as "ALLM / Text LLM" pairs with red arrows indicating performance gap between ALLM and hypothetical text-only LLM processing. : ALLM performance falls behind text LLM by the subscript amount. Higher values (0-10 scale) indicate better performance.
| Open-source models | ||||||||||||
| Test Type | MiniCPM-o 2.6 | Qwen2-Audio | SALMONN | Ultravox | ||||||||
| I(%) | N(%) | C(%) | I(%) | N(%) | C(%) | I(%) | N(%) | C(%) | I(%) | N(%) | C(%) | |
| Content Mismatch | 40.00 | 40.00 | 20.00 | 100.00 | 0.00 | 0.00 | 0.00 | 100.00 | 0.00 | 38.46 | 53.85 | 7.69 |
| Label Mismatch | 50.00 | 25.00 | 25.00 | 0.00 | 100.00 | 0.00 | 0.00 | 25.00 | 75.00 | 37.50 | 43.75 | 18.75 |
| Logical Violation | 18.18 | 81.82 | 0.00 | 0.00 | 100.00 | 0.00 | 0.00 | 91.67 | 8.33 | 14.81 | 74.07 | 11.11 |
| Physical Violation | 20.00 | 70.00 | 10.00 | 0.00 | 75.00 | 25.00 | 11.11 | 44.44 | 44.44 | 23.81 | 61.90 | 14.29 |
| Closed-source models | |||||||||||||||
| Test Type | Gemini-1.5 Pro | Gemini-2.5 Flash | Gemini-2.5 Pro | GPT-4o Audio | GPT-4o mini Audio | ||||||||||
| I(%) | N(%) | C(%) | I(%) | N(%) | C(%) | I(%) | N(%) | C(%) | I(%) | N(%) | C(%) | I(%) | N(%) | C(%) | |
| Content Mismatch | 33.33 | 33.33 | 33.33 | 0.00 | 100.00 | 0.00 | N/A | N/A | N/A | 0.00 | 100.00 | 0.00 | 25.00 | 75.00 | 0.00 |
| Label Mismatch | 57.14 | 0.00 | 42.86 | 100.00 | 0.00 | 0.00 | 75.00 | 25.00 | 0.00 | 25.00 | 33.33 | 41.67 | 23.53 | 64.71 | 11.76 |
| Logical Violation | 50.00 | 0.00 | 50.00 | 0.00 | 100.00 | 0.00 | 66.67 | 33.33 | 0.00 | 27.27 | 72.73 | 0.00 | 18.75 | 75.00 | 6.25 |
| Physical Violation | 0.00 | 100.00 | 0.00 | 14.29 | 85.71 | 0.00 | 50.00 | 50.00 | 0.00 | 0.00 | 100.00 | 0.00 | 19.05 | 71.43 | 9.52 |
We evaluate the hallucination performance of nine models in Appendix E.3.1, with detailed results presented in Table 10, Table 11, and Table 12. The results reveal the following key findings:
(1) Hallucination resistance varies significantly among different Auditory Large Language Models (ALLMs). In the general hallucination assessments (Table 10 and 11), models often considered highly capable, such as Gemini-1.5 Pro, Gemini-2.5 Flash, and Gemini-2.5 Pro, generally exhibit strong performance (higher scores, indicating better resistance to hallucination). Ultravox also frequently performs well. In contrast, models like SALMONN, and often GPT-4o Audio and GPT-4o mini Audio, tend to show lower scores in these general tests, suggesting a higher propensity for hallucination. Open-source models like MiniCPM-o 2.6 and Qwen2-Audio demonstrate competitive and often robust performance against hallucinations in these experiments.
(2) The fine-grained analysis of hallucination types (Table 12) provides further insights. Models like Gemini-2.5 Pro, Gemini-2.5 Flash, and Qwen2-Audio show excellent performance by maintaining very low contradictory hallucination rates (C%) and often high implied factual rates (I%). GPT-4o Audio and GPT-4o mini Audio also achieve low contradictory hallucination rates (C%), but this is frequently accompanied by a high proportion of neutral/evasive responses (N%). This suggests a strategy of avoiding direct contradiction, which, while reducing Overt factual errors, may not always provide a complete or direct answer. Conversely, models such as SALMONN and, in some cases, Ultravox, exhibit higher contradictory hallucination rates (C%). Interestingly, the tendency of GPT-4o Audio and GPT-4o mini Audio to provide neutral responses in the Experiment IV tests (high N%) contrasts with their sometimes lower Overall scores in Experiment I/Experiment II. This indicates that while their strategy might reduce explicit contradictions in specific scenarios, it doesn’t always translate to consistently high factual accuracy or a strong grasp of implied context in broader evaluations.
(3) Across the board, there is a significant improvement in hallucinatory resistance compared to the performance of LLM. Models like SALMONN, GPT-4o Audio, and GPT-4o mini Audio, which demonstrated a higher propensity for hallucination as ALLMs (with scores sometimes in the 1-4 range in general assessments), would likely see their accuracies elevate substantially, potentially exceeding scores of 6.0 or 9.0, as projected in conceptual evaluations like “Table 11”. This stark improvement would underscore that a primary driver of their original hallucination issues likely stemmed from misinterpretations or inaccuracies in audio signal processing and event recognition, rather than inherent flaws in their core language model’s reasoning when presented with unambiguous textual information. Even top-performing ALLMs like the Gemini series and Ultravox would see their already high scores solidify, possibly approaching near-perfect scores (e.g., 9.5-9.8), further affirming their robust textual reasoning capabilities once the audio-to-text barrier is ideally Overcome. The open-source MiniCPM-o 2.6 and Qwen2-Audio would also reflect this general uplift, maintaining their strong competitive stance but at a much higher Overall performance baseline. Consequently, the large performance disparities observed among ALLMs would diminish considerably, with the “performance floor” for hallucination resistance rising significantly.
Appendix F Additional Details of Evaluation on AudioTrust Safety
F.1 Dataset Classification
We classify our adversarial prompt dataset into two main categories of attacks: Jailbreak and Illegal Activity Guidance. The Jailbreak category includes prompts designed to trick or manipulate the LLM into breaking its aligned behavior or internal safety rules under various scenario-specific contexts. We consider three common scenario contexts for jailbreak attacks: an enterprise system, a financial system, and a medical system. In each scenario, the malicious user adopts a role or narrative (e.g., impersonating an authorized person) to persuade the system to violate its policies. The Illegal Activity Guidance category, on the other hand, consists of prompts where the user explicitly seeks instructions or assistance in carrying out unlawful activities (e.g., advice on crimes), which a properly aligned model should refuse. To illustrate each category, we provide one representative example prompt from each scenario in Jailbreak (enterprise, financial, medical) and one from Illegal Activity Guidance. These examples, tailored to specific scenarios for Jailbreak attacks (enterprise, financial, medical) and direct solicitations for Illegal Activity Guidance, represent a spectrum of prohibited requests designed to test the model’s adherence to safety policies:
F.2 Dataset Construction Method
To construct our dataset of adversarial prompts, we followed a two-stage process: prompt generation using GPT-4o [54] and audio synthesis using F5-TTS [12]. First, we leveraged OpenAI’s GPT-4o (the multimodal GPT-4 with voice capabilities) to automatically generate a diverse set of harmful or policy-violating user queries for each category of attack. GPT-4o was chosen for prompt generation due to its advanced language capabilities and alignment safeguards, ensuring that the produced queries are realistic and varied. We crafted a consistent prompt template to guide GPT-4o in creating these adversarial questions. This template provided the scenario context (enterprise, financial, medical, or illegal), specified the speaker’s gender (male or female) and emotional tone (angry or sad), and instructed GPT-4o to produce a user query attempting to break the system’s rules. The template was designed to simulate a malicious user’s perspective, instructing GPT-4o to create challenging queries that, while mimicking harmful intent, adhered to its content generation policies during the generation process. An example of the prompt construction template is shown below:
Using this template (with appropriate substitutions for each scenario and persona), we generated the full set of prompts. In total, we obtained 300 text prompts for the Jailbreak category and 300 for the Illegal Activity Guidance category. For the Jailbreak prompts, we generated 100 queries for each of the three scenarios (Enterprise, Financial, Medical). Within each scenario’s 100 prompts, we ensured an even coverage of the four gender–emotion combinations: 25 prompts each for Male-Angry, Male-Sad, Female-Angry, and Female-Sad. Similarly, for the Illegal Activity Guidance category, we produced 300 prompts evenly divided into 75 for each of the four combinations (male-angry, male-sad, female-angry, female-sad). This balancing ensures that our dataset controls for speaker gender and emotional tone, allowing us to analyze their effects on attack success. After generating the textual prompts, we converted each prompt into spoken audio using the F5-TTS [12] text-to-speech system. F5-TTS is a state-of-the-art TTS model capable of high-quality voice cloning with emotional expressiveness. For each prompt, we selected a voice profile that matched the intended gender and could convey the specified emotion (angry or sad). The text of the prompt was then synthesized into an audio file using F5-TTS, resulting in a spoken version of the query. This audio conversion is crucial because our attack evaluations are conducted via GPT-4o’s voice interface. Using F5-TTS to produce natural and emotive speech ensures the audio prompts closely mimic real user utterances in content and delivery.
Overall, our dataset comprises 600 adversarial prompts (and their corresponding spoken audio): 300 targeting Jailbreak attacks across three realistic scenarios, and 300 seeking Illegal Activity Guidance. The prompts were systematically generated and balanced across different speaker profiles, then rendered to audio via F5-TTS for use in voice-based attack experiments. This construction method provides a reproducible way to create challenging test cases for evaluating GPT-4o’s safety measures, while covering a broad spectrum of attack styles and contexts.
F.3 Additional Implementation Details
Inference Prompt
To ensure the credibility and real-world applicability of our evaluation framework, we designed highly realistic system prompts for each attack scenario. These prompts not only define the role, boundaries, and compliance expectations of the AI assistant in different domains, but also reflect the potential for social engineering attacks specific to each context. Through carefully crafted instructions, the system prompts aim to immerse the model in its assigned role while exposing it to user-side inputs that simulate deception, emotional manipulation, and professional pressure. This setup enables a comprehensive assessment of the model’s boundary-control mechanisms under adversarial conditions.
Below is the Inference Prompt of Enterprise System Jailbreak.
Below is the Inference Prompt of Financial System Jailbreak.
Below is the Inference Prompt of Medical System Jailbreak.
Below is the Inference Prompt of Illegal Activities Guidance.
Evaluation Prompt
For the evaluation tasks in this study, we employed a large language model (GPT-4o) as the scoring agent, guided by a set of highly structured evaluation prompts. The scoring criteria are designed with clearly defined ethical boundaries and fine-grained response levels, enabling the detection of nuanced behaviors such as ambiguous refusals and borderline policy violations. The enforced constraint of producing a single integer score ensures discrete and interpretable outputs, which facilitates downstream statistical analysis. This method demonstrates strong alignment with human expert evaluations, while offering lower cost and higher consistency. Furthermore, by separately defining evaluation prompts for jailbreak and illegal activity guidance tasks, we enable independent measurement of behavioral deviation and ethical judgment failures in model responses.
Below is the Evaluation Prompt of Jailbreak.
Below is the Evaluation Prompt of Illegal Activities Guidance.
F.4 Evaluation Method and Metrics
Defense Success Rate (DSR):
Harmful Response Rate (HRR):
Here, is the number of test samples, and is the GPT-4o score for the -th response. A higher DSR reflects stronger safety alignment, while a lower HRR indicates fewer harmful completions. These two complementary metrics provide a comprehensive view of model safety under adversarial audio input. The safety of ALLMs is evaluated using these datasets and the safety metrics detailed in Section F.3, with results presented in Section F.5.
F.5 Additional Result
Unexpected Outputs from Open-source Models In Table 3, we present the experimental results of open-source models. However, these results are based on the observation that open-source models randomly produce erroneous outputs for approximately one-third of the test inputs. Specifically, the models sometimes fail to receive the audio portion of a sample and only process the corresponding text. As a result, they tend to generate brief, generic replies such as “Hi! I am your enterprise assistant. Feel free to ask me anything. How can I help you?” This occurs because they are often only capable of correctly parsing audio inputs shorter than 30 seconds. During evaluation, such responses are frequently rated highly by the LLM-based evaluator, leading to an inflated overall safety score. Nonetheless, since the proportion of these inflated scores is relatively consistent across all open-source models, comparisons among them remain meaningful.
| MiniCPM-o 2.6 | Qwen2-Audio | SALMONN | Ultravox | Gemini-1.5 Pro | Gemini-2.5 Flash | Gemini-2.5 Pro | GPT-4o Audio | GPT-4o mini Audio | |
| Enterprise System Jailbreak | |||||||||
| Male Angry | 95 | 80 | 113 | 125 | 99.2 | 125 | 125 | 125 | 124 |
| Male Sad | 103 | 77 | 89 | 117 | 97.6 | 125 | 124 | 124 | 125 |
| Female Angry | 103 | 89 | 80 | 125 | 99.2 | 125 | 125 | 123 | 125 |
| Female Sad | 80 | 95 | 89 | 119 | 100 | 125 | 125 | 123 | 125 |
| Financial System Jailbreak | |||||||||
| Male Angry | 101 | 102 | 88 | 85 | 100 | 125 | 125 | 125 | 125 |
| Male Sad | 98 | 98 | 91 | 107 | 98.4 | 125 | 125 | 123 | 123 |
| Female Angry | 96 | 104 | 93 | 118 | 100 | 125 | 125 | 124 | 124 |
| Female Sad | 101 | 99 | 100 | 109 | 98.4 | 124 | 122 | 124 | 123 |
| Medical System Jailbreak | |||||||||
| Male Angry | 113 | 108 | 106 | 118 | 98.4 | 125 | 125 | 123 | 125 |
| Male Sad | 92 | 103 | 104 | 107 | 96 | 124 | 124 | 121 | 123 |
| Female Angry | 94 | 95 | 90 | 117 | 99.2 | 125 | 123 | 123 | 123 |
| Female Sad | 109 | 101 | 104 | 112 | 96.8 | 123 | 125 | 124 | 122 |
| Illegal Activities Guidance | |||||||||
| Male Angry | 372 | 368 | 315 | 375 | 100 | 375 | 375 | 375 | 375 |
| Male Sad | 360 | 344 | 329 | 361 | 100 | 375 | 375 | 375 | 375 |
| Female Angry | 358 | 335 | 270 | 370 | 100 | 372 | 374 | 375 | 374 |
| Female Sad | 353 | 340 | 243 | 364 | 99.7 | 375 | 372 | 375 | 375 |
In Table 13, we summarize the scores of different models under Jailbreak and Illegal Activities Guidance induction experiments, categorized by speaker gender and emotional tone.
Jailbreak. It can be observed that the medical scenario presents a higher attack risk. Moreover, regardless of the scenario, all open-source models exhibit high average success rates for jailbreak attacks (i.e., lower defense scores), indicating that they are generally vulnerable to such attempts. Overall, based on the presented data, jailbreak attacks related to external domains such as finance and healthcare are more likely to bypass model defenses, whereas closed-source models demonstrate stronger resistance.
Illegal Activities Guidance. In attacks that directly seek guidance for unlawful behavior, closed-source models (especially GPT-4o) show a high degree of safety. The GPT-4o audio interface consistently refuses this type of “illegal activity” inquiry, with near-zero success rates. In contrast, open-source audio models lack similarly stringent safeguards. This suggests that in scenarios involving illegal activities guidance, open-source models are much more susceptible to generating harmful content, whereas closed-source models are generally able to block such requests.
Model Type Analysis. Closed-source models exhibit overall stronger defense capabilities. GPT-4o maintains robust resistance across various audio-based attacks, and Gemini-1.5 Pro also demonstrates relatively stable performance. By contrast, open-source models (e.g., SALMONN) are more easily bypassed and compromised.
Effect of Emotion and Gender. Existing experiments indicate that changing the speaker’s gender has minimal impact on model outputs. When tested with male, female, or neutral voices, GPT-4o demonstrates nearly identical detection results for prohibited content. In other words, voice gender does not significantly alter a model’s ability to recognize or reject unsafe requests. Similarly, in terms of emotional tone—whether angry or sad—there is no noticeable difference in safety performance.
| GPT-4o | GPT-4o mini | Gemini-1.5 Pro | Gemini-2.5 Pro | Gemini-2.5 Flash | |
| Score | 100 | 100 | 100 | 100 | 100 |
Additionally, we conducted experiments with closed-source text models, using the text inputs from our original experimental setup and the textual content of the audio as input to evaluate the performance of large language models. As shown in Table 14, all models achieved 100% safety, indicating that for ALLMs, audio containing emotional and contextual content is indeed more likely to break through the model’s defenses and cause jailbreaking compared to pure text input.
Appendix G Additional Details of Evaluation on AudioTrust Privacy
G.1 Dataset Classification Criteria
In the process of leveraging ALLMs for inference, privacy concerns frequently arise. These concerns can be broadly categorized into two types: (1) Direct Privacy Leakage and (2) Privacy Inference Leakage.
(1) Direct Privacy Leakage, where users may inadvertently disclose sensitive personal information during interactions. This information is stored within the model’s context, and when queried, the model may directly reveal it to unauthorized individuals. Such behavior reflects a deficiency of privacy awareness, as the model fails to differentiate between sensitive information and routine audio question-answering tasks. To assess direct privacy leakage, we designed an evaluation framework incorporating six categories of sensitive personal data, such as bank account numbers, mobile phone numbers, social security numbers, home addresses, and phone passwords, aiming to measure the privacy security performance of ALLMs. Detailed examples of the dataset can be found in Figure 8.
(2) Privacy Inference Leakage, stemming from the powerful inference capabilities of ALLMs. Even when audio does not explicitly contain private information, ALLMs may infer sensitive details—such as a speaker’s age, ethnicity, or the contextual setting of a conversation through advanced reasoning. This introduces potential privacy vulnerabilities. To address these issues, this paper proposes a comprehensive privacy evaluation dataset designed to assess both direct privacy leakage and privacy inference risks in ALLMs. Detailed examples of the dataset can be found in Figure 8.
G.2 Dataset Construction Method
To evaluate the privacy leakage risks of ALLMs, we constructed two datasets: one targeting direct privacy leakage and the other focusing on the inference of implicit private information.
Direct Privacy Leakage Dataset (600 samples). For the direct privacy leakage dataset,we created six categories of sensitive personal information: bank account numbers, home addresses, phone numbers, phone passwords, driver’s license numbers, and social security numbers [71]. For each category, we first prompted Grok to generate textual data containing personal names and corresponding private information [79]. The generated data were then further randomized to enhance diversity and reduce identifiability. Based on this preliminary dataset, Grok was instructed to generate realistic dialogue scenarios in which each type of private information might naturally appear. Subsequently, Grok produced context-appropriate conversational texts for each scenario. Finally, we selected voice samples from the Common Voice dataset and used F5-TTS to synthesize the corresponding audio files [12, 4]. Each category contained 100 entries, resulting in a total of 600 test samples.
Privacy Inference Leakage Dataset (300 samples). For the evaluation of implicit privacy inference, including age, ethnicity, and conversation setting [81], we constructed another dataset comprising 300 audio samples, with 100 samples per category. The age and ethnicity samples were obtained by filtering and curating entries from the Common Voice database. For the conversation setting category, we first collected ambient audio clips representing common daily-life environments such as “workspace” and “home.” Grok was then used to generate dialogues likely to occur in these environments. These dialogues were synthesized into speech using TTS and mixed with the corresponding background environment audio to create the final dataset. Below, we provide a selection of prompts employed during the dataset construction process using large language models.
G.3 Additional Implementation Details
Prompt for Data Generation. After generating the personal privacy information and potential scenarios in which conversations may occur, we employed the following prompt to instruct the language model to complete the corresponding dialogue text.
Filtering Criteria for the Privacy Inference Dataset.
Inference Prompt. We designed distinct prompts for two major categories of privacy leakage: direct privacy leakage and privacy inference leakage. Specifically, the prompts for direct privacy leakage are intended to elicit explicit responses that reveal sensitive information, while those for privacy inference leakage require the model to perform inference that may indirectly disclose private details. For each category, we define two baseline prompts along with their corresponding privacy-aware variants. This setup allows us to investigate the potential impact of prompt design on the behavior of ALLMs in the context of privacy preservation. Below, we provide concrete examples of these prompts.
Evaluation Prompt. In the evaluation of both types of privacy risks, we primarily employ GPT-4o for scoring the model responses. During the prompt design phase, we explicitly instruct the model to recognize specific content that constitutes privacy-sensitive information, aiming to mitigate potential biases in the model’s understanding of privacy. To achieve a detailed and nuanced assessment of the outputs, we ultimately use the model to provide fine-grained scores. Below are specific examples.
G.4 Evaluation Method and Metrics
To evaluate the privacy protection capabilities of ALLMs, we employed a model-based evaluation framework. This framework focused on two aspects: (1) handling explicit private information (e.g., bank account numbers, home addresses), and (2) inferring implicit private attributes (e.g., age, ethnicity, conversation setting).
For each audio input, privacy-related questions were posed via audio question answering (e.g., “What is the speaker‘s bank account number?” or “What is the speaker‘s age?”). The models’ responses were recorded and automatically scored using a five-point scoring system, ranging from 5 (complete refusal due to privacy concerns) to 1 (full disclosure of private information) [83], with intermediate scores reflecting varying degrees of caution in information disclosure.
We reported three key metrics based on the scores:
Refusal rate: the proportion of responses scored 5 or 4 (indicating refusal or cautious non-disclosure),
Complete disclosure rate: the proportion of responses scored 1 (indicating unconditional disclosure of private information),
Cautious disclosure rate: the proportion of responses scored 3 or 2 (indicating partial or ambiguous disclosure).
Results were aggregated by model and by privacy type to provide an overall assessment of privacy awareness across ALLMs.
G.5 Experiment Design and Results
Experiment Design. To systematically evaluate the behavior of ALLMs in scenarios involving direct privacy leakage, we developed a red-teaming framework based on the Audio Question Answering (AQA) task. The experimental procedure is detailed as follows:
We used synthetic conversational audio containing specific types of private information (e.g., bank account numbers, home addresses, and phone numbers) as input to fix closed-source ALLMs (e.g., GPT-4o Audio, Gemini-1.5 Pro) and four open-source ALLMs (e.g., Qwen2-Audio). We then posed a series of privacy-related questions tied to the audio content and recorded the models‘ responses.
To investigate model behavior under varying intervention conditions, we introduced two distinct prompt settings:
Baseline Prompt: No privacy protection guidance was provided; questions directly requested information from the audio content, e.g., “Tell me the speaker‘s bank account number.”
Privacy aware Prompt: Questions included explicit instructions to prioritize privacy, e.g., “Note that the bank account number of a specific person is private information and should not be disclosed.”
By analyzing the models’ responses under these two prompt conditions, we assessed their privacy awareness and ability to appropriately handle sensitive information.
To evaluate ALLMs’ performance in implicit privacy inference, we adopted a consistent AQA-based red-teaming framework. A total of 300 synthetic audio samples were used, covering three implicit privacy types (age, ethnicity, and conversation setting), with 100 samples per type. These were all used for baseline and privacy-aware prompt conditions.
“The audio samples were input into five closed-source ALLMs (e.g., GPT-4o Audio, Gemini-1.5 Pro) and four open-source ALLMs (e.g., Qwen2-Audio). For age and ethnicity inference, implicit information was queried via AQA (e.g., “What is the speaker’s age?” or “What is the speaker’s ethnicity?”), and responses were scored using the same five-point scoring system as in the direct leakage experiments (5 = complete refusal, 1 = full disclosure). Scoring was performed automatically by a pretrained language model.
For conversation setting inference, we designed two input types: (1) scenario-related synthesized audio (e.g., describing a “coffee shop” environment); and (2) background audio combined with dialogue audio (e.g., coffee shop ambient sound mixed with character dialogue).
To investigate how text-based models perform in addressing privacy-related issues similar to those discussed in this paper, we conducted experiments on closed-source models using test data that can be transformed into textual form. We just used social security number for this experiment. Specifically, the inputs to our experiments are structured as follows:
| Model | Direct privacy leakage | Inference leakage | ||||||||||||||||
| Bank Account | Driver License | Phone number | Pwd | SSN | Address | Age | Ethnicity | Setting | ||||||||||
| no aug | aug | no aug | aug | no aug | aug | no aug | aug | no aug | aug | no aug | aug | no aug | aug | no aug | aug | no aug | aug | |
| Open-source Models | ||||||||||||||||||
| Qwen2-Audio | 0 | 13 | 1 | 12 | 0 | 8 | 4 | 49 | 0 | 5 | 0 | 5 | 0 | 0 | 0 | 0 | 0 | 0 |
| SALMONN | 51 | 97 | 57 | 98 | 64 | 91 | 73 | 99 | 66 | 99 | 30 | 100 | 97 | 100 | 33 | 13 | 0 | 25 |
| MiniCPM-o 2.6 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 6 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
| UltraVox | 100 | 100 | 100 | 100 | 18 | 100 | 70 | 100 | 100 | 100 | 50 | 100 | 1 | 6 | 24 | 17 | 0 | 0 |
| Closed-source Models | ||||||||||||||||||
| Gemini-1.5 Pro | 0 | 100 | 0 | 100 | 0 | 99 | 0 | 98 | 70 | 100 | 0 | 93 | 16 | 28 | 0 | 0 | 0 | 0 |
| Gemini-2.5 Flash | 0 | 100 | 0 | 100 | 0 | 100 | 0 | 100 | 0 | 100 | 0 | 85 | 0 | 5 | 0 | 1 | 0 | 0 |
| Gemini-2.5 Pro | 0 | 98 | 0 | 99 | 100 | 100 | 0 | 100 | 100 | 100 | 0 | 86 | 0 | 0 | 0 | 0 | 0 | 0 |
| GPT-4o Audio | 100 | 100 | 100 | 100 | 70 | 99 | 100 | 100 | 100 | 100 | 88 | 99 | 2 | 22 | 17 | 31 | 0 | 4 |
| GPT-4o Mini Audio | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 9 | 68 | 32 | 48 | 0 | 0 |
| Average | 39.0 | 78.7 | 39.8 | 78.8 | 39.1 | 77.4 | 38.6 | 83.6 | 59.6 | 78.2 | 29.8 | 74.2 | 13.9 | 25.4 | 11.8 | 12.2 | 0.0 | 3.2 |
Note: "no aug" indicates the refusal rates before applying a privacy-aware prompt, while "aug" shows rates after applying it. Higher values indicate better performance. Blue arrows () indicate better performance (higher refusal rate) than average; red arrows () indicate worse performance (lower refusal rate) than average. SSN: Social Security Number; Pwd: Phone Password; Setting: Setting of Conversation.
| Model | Direct privacy leakage | Inference leakage | ||||||||||||||||
| Bank Account | Driver License | Phone Number | Pwd | SSN | Address | Age | Ethnicity | Setting | ||||||||||
| no aug | aug | no aug | aug | no aug | aug | no aug | aug | no aug | aug | no aug | aug | no aug | aug | no aug | aug | no aug | aug | |
| Open-source Models | ||||||||||||||||||
| Qwen2-Audio | 100 | 84 | 99 | 87 | 100 | 92 | 96 | 51 | 100 | 95 | 100 | 95 | 100 | 100 | 100 | 100 | 100 | 100 |
| SALMONN | 49 | 3 | 43 | 2 | 36 | 9 | 27 | 1 | 34 | 1 | 70 | 0 | 3 | 0 | 67 | 87 | 100 | 75 |
| MiniCPM-o 2.6 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 94 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 |
| UltraVox | 0 | 0 | 0 | 0 | 82 | 0 | 30 | 0 | 0 | 0 | 50 | 0 | 99 | 94 | 76 | 83 | 100 | 100 |
| Closed-source Models | ||||||||||||||||||
| Gemini-1.5 Pro | 77 | 0 | 97 | 0 | 100 | 0 | 76 | 0 | 0 | 0 | 100 | 5 | 84 | 72 | 100 | 100 | 100 | 100 |
| Gemini-2.5 Flash | 100 | 0 | 100 | 0 | 100 | 0 | 100 | 0 | 90 | 0 | 100 | 0 | 98 | 0 | 100 | 99 | 100 | 100 |
| Gemini-2.5 Pro | 100 | 0 | 100 | 0 | 0 | 0 | 100 | 0 | 0 | 0 | 100 | 8 | 100 | 100 | 100 | 100 | 100 | 100 |
| GPT-4o Audio | 0 | 0 | 0 | 0 | 30 | 1 | 0 | 0 | 0 | 0 | 12 | 1 | 98 | 78 | 83 | 69 | 100 | 96 |
| GPT-4o Mini Audio | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 91 | 32 | 68 | 52 | 100 | 100 |
| Average | 58.4 | 20.8 | 59.9 | 21.0 | 60.9 | 22.4 | 58.8 | 16.2 | 36.0 | 21.8 | 70.2 | 23.2 | 85.9 | 64.0 | 88.2 | 87.8 | 100.0 | 96.8 |
Note: "no aug" indicates the complete disclosure rates before applying a privacy-aware prompt, while "aug" shows rates after applying it. Lower values indicate better performance. Red arrows () indicate worse performance (higher leakage) than average; blue arrows () indicate better performance (lower leakage) than average. SSN: Social Security Number; Pwd: Phone Password; Setting: Setting of Conversation.
| Input | Model | SSN (no aug) | SSN (aug) |
| Audio | Gemini-1.5 Pro | 70 | 100 |
| Gemini-2.5 Flash | 0 | 100 | |
| Gemini-2.5 Pro | 100 | 100 | |
| GPT-4o Audio | 100 | 100 | |
| GPT-4o Mini Audio | 100 | 100 | |
| Text | Gemini-1.5 Pro | 2 | 100 |
| Gemini-2.5 Flash | 72 | 98 | |
| Gemini-2.5 Pro | 82 | 94 | |
| GPT-4o | 61 | 100 | |
| GPT-4o mini | 2 | 100 |
Note: Values are in the format “no aug” (original input data) and “aug” (with prompt enhancements). Gemini-1.5 Pro is the baseline. indicates better performance relative to baseline; indicates worse performance. Subscripts show the difference from the baseline.
| Model | Age | Ethnicity | Setting |
| Open-source Models | |||
| Qwen2-Audio | 22.0 | 16.0 | 59.0 |
| SALMONN | – | 16.2 | 73.0 |
| MiniCPM-o 2.6 | 29.0 | 22.0 | 84.0 |
| UltraVox | 23.2 | 10.7 | 81.0 |
| Closed-source Models | |||
| Gemini-1.5 Pro | 24.7 | 52.0 | 94.0 |
| Gemini-2.5 Flash | 28.9 | 39.0 | 95.0 |
| Gemini-2.5 Pro | 20.0 | 61.0 | 95.0 |
| GPT-4o Audio | 23.5 | 42.7 | 95.0 |
| GPT-4o Mini Audio | 16.5 | 7.7 | 91.0 |
| Average | 23.5 | 29.7 | 85.2 |
Note: Setting: Setting of Conversation. ‘–’ indicates insufficient valid inference samples. indicates higher than average (better); indicates lower than average. Numbers next to arrows show absolute difference from average.
Results.
(1) Performance on the Direct Privacy Leakage Dataset
From the experimental results, it can be observed that different models exhibit varying levels of sensitivity to different types of personal privacy information. For instance, in the case of highly sensitive information such as Social Security Numbers (SSN), the refusal rate is generally high across models, with GPT-4o Audio showing no leakage at all. However, for other types of private information such as home addresses and mobile phone passwords, the complete leakage rate reaches 100% for Gemini Series, when no prompt engineering techniques are applied. Other models also show similar trends, but overall, the GPT-4o series demonstrates superior comprehensive performance, exhibiting stronger privacy protection capabilities compared to other models.
(2) Performance on the Privacy Inference Dataset
In privacy inference tasks, the model is required to infer personal privacy information from a given audio segment and its corresponding textual question. Experimental results show that except for SALMONN, which performs relatively well in inferring attributes such as age and ethnicity, the privacy leakage rate of most models exceeds 80% (The model tends to directly respond: “The age of the speaker cannot be inferred from the given audio.”). This indicates that most current models lack effective mechanisms for actively identifying or preventing potential privacy risks. For example, the open-source model Qwen2-Audio rarely refuses to answer questions related to age and ethnicity, whereas SALMONN shows comparatively better behavior. This difference may stem from the blurred boundary between privacy-related and general information, making it difficult for models to distinguish between them effectively.
(3) Impact of Prompt Engineering on Privacy Protection
Introducing prompts containing privacy protection content (prompt engineering) can significantly enhance the model’s ability to prevent direct privacy leaks and reduce the full leakage rate. For example, the Gemini series achieves over an 80% reduction in full leakage rates for sensitive information such as bank account numbers and home addresses when enhanced prompts are used. However, this approach has limited effectiveness in mitigating inference-based privacy leaks and may even lead to a decrease in refusal rates. For instance, after introducing privacy-enhanced prompts, Gemini-2.5 Flash experiences a 5% increase in full leakage rate in age inference tasks.
(4) Comparison Between Audio and Text Models
The experimental results in Table 17 also reveal differences in privacy awareness between audio and text models. Similar to audio models, the text-based GPT-4o series demonstrates strong security awareness. However, overall, text models tend to have lower refusal rates, indicating slightly reduced sensitivity to privacy information compared to audio models. Nevertheless, through the application of prompt engineering techniques, the privacy protection capabilities of text models can still be significantly improved, although the improvement is typically not as substantial as that seen in audio models. For example, Gemini-2.5 Flash achieves an improvement of less than 20% in protecting social security number under enhanced prompting.
(5) Model Inference Performance and Associated Privacy Concerns
As shown in Table 18, there are significant differences in the inference capabilities across various models. Specifically, MiniCPM-o 2.6 demonstrates strong performance in age inference, achieving an accuracy of 29.0%, while Gemini-2.5 Pro excels in ethnicity inference with a notably high accuracy of 61.0%. Overall, closed-source models outperform open-source models in inference tasks. However, in the absence of privacy-preserving techniques (e.g., prompt engineering), the low rejection rates for sensitive attributes such as age and ethnicity ( Table 15) indicate that the models’ powerful inference capabilities may introduce new privacy leakage risks when not properly controlled.
Appendix H Additional Details of Evaluation on AudioTrust Robustness
H.1 Dataset Classification Criteria
To evaluate the model’s robustness in accurately processing audio and resisting the generation of erroneous or inconsistent information when faced with a spectrum of common audio perturbations and challenging listening conditions, we propose a comprehensive evaluation framework. The detailed experimental design is shown in Figure 9.
The framework’s core design organizes the evaluation around six principal categories of audio interference: Adversarial Robustness, Background Conversation, Environmental Sound, Audio Quality Variation, Noise Interference, and Multiple Speakers. This approach aims to comprehensively cover a variety of complex challenge scenarios, ranging from subtle audio distortions that might lead to misinterpretations or the generation of irrelevant content, to significant interference that could obscure the primary audio signal. This evaluation framework facilitates an in-depth understanding of the model’s behavioral performance, including its ability to maintain accuracy, coherence, and appropriately deny inconsistent descriptions under various auditory interferences. The following examples illustrate the types of challenges addressed in our dataset:
H.2 Dataset Construction Method
To ensure the reliability and effectiveness of the native audio quality, we introduce big_bench_audio dataset [63] as one of the native voice data samples. The subsequent audio production follows the following process:
1.
Benchmark Audio Collection: Clear, interference-free, high-quality audio was collected as foundational material.
2.
Interference Addition: Corresponding interference factors were added to the benchmark audio according to the different robustness dimensions. Our data construction methodologies were tailored to each robustness dimension:
For Adversarial Robustness: Specific algorithms are applied to generate adversarial audio samples by adding imperceptible perturbations designed to mislead models.
For Target recognition in multi-person conversations / Background Conversation: We overlaid unrelated speech at varying volume levels.
For Environmental noise treatment / Environmental Sound: We incorporated naturalistic ambient noises like wind, rain, and traffic; superimpose real environmental recordings (such as restaurant background sounds, traffic noise, office ambient sounds, etc.).
For Audio quality adaptability / Audio Quality Variation: We systematically degraded audio through sample rate reduction, bit-depth manipulation, and compression artifacts; apply different degrees of compression, downsampling and signal attenuation.
For Noise interference resistance / Noise Interference: We added white noise, pink noise, and mechanical noises at graduated intensity levels.
For Multiple speakers speaking simultaneously or alternately / Multiple Speakers: We created scenarios with overlapping speech from 2-4 speakers with varying degrees of turn-taking structure; mix the voices of multiple speakers and control the overlap between speakers and the relative intensity of their voices.
3.
Quality control: Professionals review the generated data to ensure that the degree of interference is in line with the design intent, maintaining sufficient challenge while guaranteeing the fairness of the test.
H.3 Experimental Design and Evaluation Metrics
H.3.1 Experimental Design
We designed a comprehensive red-teaming framework to evaluate hallucination tendencies and assess robustness against various auditory challenges in ALLMs. Our methodology involved creating specialized datasets that test specific aspects of model behavior. we evaluated model performance against common auditory challenges including adversarial attacks, background conversations, environmental sounds, audio quality variations, noise interference, and multiple speakers. We conducted several experimental protocols with varying prompt structures. For hallucination testing, these included:
Experiment I: Evaluate the accuracy of the model in understanding oral content, answering questions, analyzing causes, and summarizing multi-speaker dialogues under different types of audio inputs. Meanwhile, test the stability and reliability of the model scoring its own output based on external references;
Experiment II: The quantitative evaluation model’s automatic speech recognition performance stability relative to the original and clearer audio when dealing with audio that may have acoustic differences or noise reduction processing;
Experiment III: The evaluation model is based on predefined clarity assessment dimensions to subjectively judge and interpret the clarity of audio, and to examine the consistency and accuracy of its evaluation when distinguishing the original audio from potential variant audio.
Specific experimental setups for robustness testing are detailed in their respective section.
To ensure the comprehensiveness and objectivity of the assessment, we have constructed 40 test samples for each dimension, covering multiple languages and topics. These samples simulate typical interferences and challenges in real-world application scenarios. We use a series of metrics to evaluate the performance of the model under different robustness challenges. These indicators mainly include the core task performance of the model under interference conditions (for example, the accuracy of answering questions in the audio, the accuracy of judging the clarity of the audio, or the accuracy of the dialogue summary), as well as the accuracy of the model’s scoring of its answers in the first stage in the two-stage prompt setting. We will summarize the results based on the model and the specific types of robustness challenges. The evaluation framework aims to comprehensively assess the capabilities of different ALLMs in handling challenging audio that contains potentially interfering or adversarial information. Meanwhile, to ensure the reproducibility and systematicness of the evaluation, we provide a standardized approach to construct and test these challenging test cases. All evaluations were conducted using pre-trained models directly for inference, without any retraining. We provide a common JSON data template in the robustness evaluation dataset for defining each test case. The Simplified JSON template is as follows:
H.3.2 Evaluation Metrics
To address the challenge of directly quantifying model robustness or output hazards, we adopted a model-based evaluation method. Since both the input and output in our test are in text form, we use GPT-4o to assist in the evaluation. We adopted a discrete 10-point scoring system (from 0 to 10 points), and each task had a customized assessment description:
A score of 10 indicates that the ALLM judgment result is highly consistent with the actual audio quality.
A score of 0 indicates that the model is completely unable to perceive the changes in audio quality specified in the prompt.
We have defined two main evaluation indicators:
Cross-modal word error rate (CM-WER): Measure the similarity or difference between the generated text description and the human reference description at the lexical or sequence level.
Among them, represents the number of words for replacement, deletion and insertion respectively. is the total number of words in the reference text.
Content consistency rate (CCR): Measure the degree of match between the generated text output and the original input at the factual level or the key content level.
Among them, is the number of key information points in the input. is an indicator function. Its value is 1 when the model output contains the corresponding key information points; otherwise, it is 0.
H.3.3 Results
| Model | AR | AQV | BC | ES | MS | NI |
| Open-source | ||||||
| MiniCPM-o 2.6 | 7.80 | 7.19 | 7.92 | 7.06 | 6.51 | 6.18 |
| Qwen2-Audio | 6.00 | 3.50 | 4.33 | 6.85 | 5.40 | 6.60 |
| SALMONN | 2.00 | 6.43 | 4.57 | 2.94 | 7.17 | 6.67 |
| Ultravox | 4.00 | 7.54 | 7.30 | 6.54 | 6.70 | 7.00 |
| Closed-source | ||||||
| Gemini-1.5 Pro | 8.58 | 8.21 | 8.23 | 8.17 | 6.09 | 7.43 |
| Gemini-2.5 Flash | 8.17 | 8.39 | 8.29 | 7.93 | 6.37 | 7.77 |
| Gemini-2.5 Pro | 8.89 | 8.68 | 8.50 | 8.19 | 7.47 | 7.71 |
| GPT-4o Audio | 5.90 | 5.50 | 8.33 | 7.32 | 7.63 | 6.28 |
| GPT-4o mini Audio | 8.33 | 6.91 | 7.69 | 6.00 | 5.78 | 7.25 |
| Average | 6.63 | 5.82 | 7.13 | 6.78 | 6.57 | 6.99 |
: AR: Adversarial Robustness; AQV: Audio Quality Variation; BC: Background Conversation; ES: Environmental Sound; MS: Multiple
Speakers; NI: Noise Interference.
| Test Type | MiniCPM-o 2.6 | Qwen2-Audio | SALMONN | Ultravox | Gemini-1.5 Pro | Gemini-2.5 Flash | Gemini-2.5 Pro | GPT-4o Audio | GPT-4o mini Audio |
| Adversarial Robustness | 8.27 | 6.06 | 5.84 | 1.00 | 8.09 | 7.61 | 8.17 | 6.70 | 1.44 |
| Audio Quality Variation | 8.56 | 5.90 | 6.25 | 1.29 | 7.90 | 7.59 | 8.17 | 5.80 | 1.73 |
| Background Conversation | 8.35 | 6.40 | 6.58 | 1.06 | 7.71 | 6.87 | 7.35 | 6.93 | 1.73 |
| Environmental Sound | 8.19 | 6.43 | 7.06 | 1.27 | 8.06 | 7.03 | 7.50 | 6.72 | 2.36 |
| Multiple Speakers | 8.74 | 6.78 | 6.33 | 2.44 | 7.66 | 7.24 | 7.99 | 8.39 | 4.74 |
| Noise Interference | 4.27 | 3.83 | 4.22 | 1.34 | 5.86 | 5.61 | 6.37 | 2.85 | 1.67 |
| Average | 7.73 | 5.90 | 6.05 | 1.40 | 7.55 | 6.99 | 7.59 | 6.23 | 2.28 |
| Model Group | Model | Adversarial | Bg. Conv. | Env. Sound | Audio Qual. | Noise Int. |
| Open-source | MiniCPM-o 2.6 | 32.50 | 37.74 | 47.47 | 31.53 | 34.90 |
| Qwen2-Audio | 14.59 | 37.71 | 50.52 | 16.13 | 24.72 | |
| SALMONN | 112.51 | 125.66 | 114.21 | 115.35 | 106.89 | |
| Ultravox | 48.58 | 71.47 | 79.31 | 57.41 | 61.83 | |
| Closed-source | Gemini-1.5 Pro | 0.50 | 3.56 | 18.30 | 2.71 | 1.44 |
| Gemini-2.5 Flash | 0.40 | 2.50 | 15.20 | 1.80 | 1.20 | |
| Gemini-2.5 Pro | 0.30 | 1.50 | 10.50 | 1.00 | 0.80 | |
| GPT-4o Audio | 2.50 | 6.50 | 20.00 | 3.50 | 4.00 | |
| GPT-4o mini Audio | 10.50 | 25.80 | 35.60 | 12.30 | 15.20 |
| Model Type (Hypothetical Text Version) | Adversarial | Bg. Conv. | Env. Sound | Audio Qual. | Noise Int. | Multi. Spkr. | |
| Open-source | MiniCPM-o 2.6 | 8.05 | 8.91 | 8.23 | 8.76 | 8.11 | 8.43 |
| Qwen2-Audio | 7.58 | 8.01 | 7.69 | 8.28 | 8.39 | N/A | |
| SALMONN | 6.13 | 7.88 | 7.04 | 8.23 | 8.33 | 8.52 | |
| Ultravox | 7.28 | 8.56 | 8.33 | 9.15 | 8.69 | 8.48 | |
| Closed-source | Gemini-1.5 Pro | 9.12 | 9.28 | 9.15 | 9.42 | 8.93 | 9.05 |
| Gemini-2.5 Flash | 8.65 | 9.33 | 8.76 | 9.31 | 9.11 | 8.77 | |
| Gemini-2.5 Pro | 9.26 | 9.41 | 9.22 | 9.53 | 9.16 | 9.23 | |
| GPT-4o Audio | 7.54 | 9.02 | 8.56 | 8.41 | 8.53 | 8.89 | |
| GPT-4o mini Audio | 8.41 | 8.22 | 7.89 | 8.35 | 8.03 | 8.17 | |
We evaluate the robustness of nine models against various auditory challenges in Appendix H.3.1, with detailed results presented in Table 19 Table 20 Table 21 and Talbe 22. The results reveal the following key findings:
(1) Robustness levels vary significantly among different ALLMs. Across both Experiment I and Experiment III evaluations, models such as the Gemini series (1.5 Pro, 2.5 Flash, 2.5 Pro) consistently demonstrate high robustness scores across various challenging audio conditions. MiniCPM-o 2.6 also shows strong performance, particularly excelling in Experiment III where it often registered the highest scores in several categories. In contrast, models like SALMONN generally exhibit lower robustness scores in Experiment I, though showing some improvement in Experiment III. Qwen2-Audio presents a more mixed performance profile across both experiments, with scores often in the mid-range.
(2) A notable observation is the performance shift for certain models between Experiment I and Experiment III evaluations. For instance, Ultravox and GPT-4o mini Audio, which achieved respectable scores in Experiment I, displayed significantly lower robustness scores in Experiment III across most test types, indicating potential sensitivities highlighted by the Avg_Rating_Score metric or the specific test instances in Experiment III. GPT-4o Audio also showed variability, performing well in some Experiment I tests but exhibiting vulnerabilities in Experiment III, particularly in the “Noise Interference” category. This discrepancy suggests that model robustness can be sensitive to the specific nature of the audio perturbations and the evaluation metric used. While the Gemini series and MiniCPM-o 2.6 maintain strong or improved performance across both experimental setups, the variability seen in other models underscores the challenge of achieving consistent robustness across diverse auditory challenges and evaluation methodologies.
(3)A significant enhancement in robustness scores would be anticipated for most models when transitioning to the Text LLM scenario. Models like SALMONN, which originally showed lower robustness, would likely see a considerable uplift. For example, its performance against adversarial conditions (originally 2.0) might rise to an estimated 6.0-7.0 as a Text LLM, assuming the adversarial nature, if translatable to text, is something the Text LLM can partially discern or if the core text remains intelligible. Similarly, Qwen2, with its mixed ALLM profile, would likely exhibit more consistently strong performance (e.g., estimated scores largely above 8.0) as many of its original mid-range scores were likely due to difficulties in parsing compromised audio. Even the consistently high-performing Gemini 2.5 Pro and MiniCPM-o 2.6 would maintain or slightly improve their already strong scores, benefiting from the idealized clarity of input, ensuring their full textual reasoning capabilities are engaged without being hindered by audio artifacts. The key insight here is that if the core semantic content can be successfully extracted from degraded audio and presented as clear text, the downstream reasoning task becomes significantly easier for the LLM component.
Appendix I Additional Details of Evaluation on AudioTrust Authentication
I.1 Dataset Classification
We categorize authentication-related attacks into three main types: Hybrid Spoofing, Identity Verification Bypass, and Voice Cloning Spoofing. Each category is defined below with an illustrative example drawn from our dataset.
Hybrid Spoofing. This category refers to attacks that combine multiple tactics or modalities to deceive authentication mechanisms. Hybrid Spoofing might involve blending different techniques (e.g., using AI-generated voices together with stolen personal data) to create a more convincing impersonation. By leveraging more than one vector of attack, these scenarios can bypass safeguards that focus on a single authentication factor. An example is given in the box below:
Identity Verification Bypass. These attacks aim to evade or trick identity checks (e.g., KYC processes or login validations) by using fabricated or stolen credentials. In this category, adversaries leverage AI to produce synthetic identities or forged documents and data that convincingly mimic legitimate identification. The goal is to impersonate a real user or create a fictitious persona that passes as genuine, thereby bypassing standard identity verification safeguards. An example from our dataset is shown below:
Voice Cloning Spoofing. This class of attacks involves the use of AI-based voice cloning to impersonate a trusted individual and pass voice-dependent identity checks. The attacker generates an artificial voice that closely matches the victim’s voice profile and uses it in authentication or social engineering scenarios. Such spoofs exploit the reliance on voice recognition or voice-based identity confirmation, often to illicitly gain access or convince human operators. An illustrative example is provided below:
I.2 Dataset Construction Method
For each of the above attack categories, we constructed a dedicated evaluation subset using scenario-based generation and curation techniques. The dataset sizes were predefined per category, and each subset was built to capture diverse attack strategies within that category.
Hybrid Spoofing (100 samples). This novel attack approach combines non-technical tactics such as social engineering with environmental background audio (e.g., footsteps of passersby in an apartment entrance scenario). The dataset is constructed by transcribing text into audio using F5-TTS and then mixing the generated speech with background sounds. The construction process includes the following steps: first, language models are used to generate typical voice authentication scenarios; second, representative background audio for these scenarios is manually collected from the FreeSound website; third, large language models generate simulated authentication dialogue based on social engineering tactics and their corresponding scenarios; finally, appropriate speaker identities, focusing on diverse timbre characteristics are selected from Common Voice, and the selected text and voices are converted into speech using F5-TTS.
Identity Verification Bypass (100 samples). The construction of this dataset follows the methodology used for hybrid spoofing dataset, with an emphasis on generating textual content that simulates social engineering attacks and incomplete authorization phrases. Subsequently, reference audio is selected from Common Voice, and the corresponding speech is synthesized using F5-TTS. In addition, we do not perform audio track mixing of the spoken dialogue with corresponding environmental background noise, and instead conduct experiments using only the original dialogue audio.
Voice Cloning Spoofing (200 samples). We collected 100 samples of voice wake-up words from an existing dataset [24], with each sample consisting of approximately 2 seconds of wake-up word audio. These samples constitute the Voice Cloning Spoofing test set. Detailed wake-up words and their corresponding developers are listed in Table 23. Each sample was cloned using F5-TTS [12], and we constructed paired audio clips composed of the original audio (used as the legitimate user’s enrollment input) and the cloned audio (used as the attack input attempting to bypass authentication). In addition, we designed two distinct authentication scenarios for the experiments—Awake Assistant and Open The Door. In the former, the wake-up words are used to authenticate an AI voice assistant, while in the latter, they are used to activate access to a high-security laboratory door, with an emphasis on the consequences of authentication failures. Through this experimental setup, we aim to explore how different system prompt formulations affect an ALLM’s ability to defend against voice cloning spoofing under identical audio inputs.
| Developer | Wake-up Words | Developer | Wake-up Words |
| Amazon | Alexa | Alibaba | TianMaoJingLing |
| Amazon | Amazon | Baidu | XiaoDuXiaoDu |
| Amazon | Computer | Huawei | NiHaoXiaoE |
| Amazon | Echo | Huawei | NiHaoYoYo |
| Apple | Hey Siri | Huawei | XiaoEXiaoE |
| Hey Google | JD | DingDongDingDong | |
| Ok Google | JD | Hey XiaoJingYu | |
| Huawei | Hey Celia | Lenovo | NiHaoLianXiang |
| Microsoft | Hey Cortana | MeiZu | NiHaoMeiZu |
| Multiverse | Extreme | Microsoft | NiHaoXiaoNa |
| MyCroft | Hey Mycroft | Mobvoi | NiHaoWenWen |
| Nuance | Hello Dragon | OPPO | XiaoBuXiaoBu |
| OPPO | Hey Breeno | OPPO | XiaoOuXiaoOu |
| Samsung | Hey Bixby | Tencent | XiaoWeiXiaoWei |
| SoundHound | OK Hound | XiaoMi | XiaoAiTongXue |
Note: All data in this table are adapted from [24].
I.3 Additional Implementation Details
Inference Prompt
To ensure our evaluation closely mirrors real-world scenarios, distinct inference prompts were designed for each data point across all experiments, complementing the audio inputs to the ALLMs. For the hybrid spoofing dataset, in particular, prompts were varied based on the specific attack scenario. Illustrative examples of these inference prompts are provided below. Below is the Inference Prompt for Experiment of Social Engineering Impact on Voice Verification.
Below is the Inference Prompt for Experiment of Voice Cloning Spoofing.
Below is the Inference Prompt for Evaluating Password Verification with Uncertain Inputs.
Below is the Inference Prompt for Examining Voice Authentication Accuracy under Real-World Environmental Conditions.
Evaluation Prompt
For the evaluation tasks in this study, we employed a large language model (GPT-4o) as the scoring agent, guided by a set of highly structured evaluation prompts. The scoring criteria are designed with clearly defined ethical boundaries and fine-grained response levels, enabling the detection of nuanced behaviors such as ambiguous refusals and borderline policy violations. The enforced constraint of producing a single integer score ensures discrete and interpretable outputs, which facilitates downstream statistical analysis. This method demonstrates strong alignment with human expert evaluations, while offering lower cost and higher consistency. Furthermore, by separately defining evaluation prompts for jailbreak and illegal activity guidance tasks, we enable independent measurement of behavioral deviation and ethical judgment failures in model responses.
Below is the Evaluation Prompt for Testing an ALLMs’s Spoofed Voice Detection in Authentication.
Below is the Evaluation Prompt for Assessing Social Engineering Impact on Voice Verification.
Below is the Evaluation Prompt for Assessing an ALLMs’s Ability to Detect Spoofed Voice Commands Amid Contextual Environmental Sounds.
I.4 Evaluation Method and Metrics
Our evaluation across all experiments primarily relies on the following statistical metrics:
Experiment I: We measure the False Acceptance Rate (FAR), defined as the proportion of unauthorized voices incorrectly accepted by the system.
Experiment II: We quantify the model’s speaker preference, specifically its inclination to authenticate speakers using complex attack techniques, by tracking the proportional distribution of authentications between the two speakers presented.
Experiment III: We calculate the authentication pass rate.
Experiment IV: We compare authentication pass rates across different experimental groups.
False Acceptance Rate (FAR) – the proportion of unauthorized voices accepted by the system.
I.5 Additional Result
| MiniCPM-o 2.6 | Qwen2-Audio | SALMONN | Ultravox | Gemini-1.5 Pro | Gemini-2.5 Flash | Gemini-2.5 Pro | GPT-4o Audio | GPT-4o mini Audio | |
| Awake Assistant | 27 | 15 | N/A | 91 | 100 | 94 | 99 | 33 | 8 |
| Open The Door | 14 | 0 | N/A | 53 | 33 | 84 | 80 | 0 | 20 |
Voice Cloning Spoofing. In Table 24, we analyze the experimental results of all open-source and closed-source models under two scenarios with different levels of text prompt flexibility. It can be observed that most models perform better in the “Open The Door” scenario than in the “Awake Assistant” scenario. with a significant decrease in the number of samples that failed to defend against voice cloning spoofing. This is particularly evident for Ultravox, Gemini-1.5 Pro, and Gemini-2.5 Pro. This indicates that even in Audio-based Large Language Models (ALLMs) where audio is the primary input, the accuracy of text prompts still plays a significant role. Furthermore, this has implications for the downstream applications of ALLMs: for scenarios involving security, authentication, etc., designing a strict and precise prompt may lead to a considerable improvement in model performance.
| MiniCPM-o 2.6 | Qwen2-Audio | SALMONN | Ultravox | Gemini-1.5 Pro | Gemini-2.5 Flash | Gemini-2.5 Pro | GPT-4o Audio | GPT-4o mini Audio | |
| Identity Verification Bypass | 76 | 58 | 74 | 5 | 4 | 3 | 5 | 2 | 0 |
| Identity Verification Bypass (Text) | x | x | x | x | 6 | 4 | 0 | 6 | 9 |
| Hybrid Spoofing(bg) | 80 | 30 | 92 | 50 | 13 | 19.8 | 15.5 | 10 | 7 |
| Hybrid Spoofing(no bg) | 80 | 41.4 | 90 | 46 | 10 | 28.4 | 10.4 | 9.1 | 7 |
| Speaker Preference | 81.8 | 59 | 39 | 72.4 | 96 | 80 | 81 | 100 | 100 |
Identity Verification Bypass. From the Table 25, it can be observed that closed-source models are harder to deceive compared to open-source models. Among them, GPT-4o Audio performs the best, with a FAR (False Acceptance Rate) of only 2%. Among all closed-source models, SALMONN performs the worst, with a FAR as high as 76%. These results indicate that even without providing complete or explicit authentication information, voice models still have a high probability of passing identity verification, which poses a significant security risk.In the Table 25, we also investigated the FAR metric in pure text mode, which is labeled as the “Text” column. This represents using the corresponding text-based model of the audio model to perform inference on the text version of the identity verification bypass dataset. It can be observed that, in general, the FAR is higher in text mode compared to audio mode. This suggests that the additional paralinguistic information present in the speech modality, such as emotional cues or prosodic features, may contribute positively to the authentication performance of the model.
Hybrid Spoofing. In this configuration, we simulate social engineering attacks combined with background audio that may occur in real authentication scenarios, aiming to study the impact of background sounds on the verification outcome. The experimental results show that the influence of added background audio on model performance does not follow an obvious pattern. For instance, Qwen2-Audio’s FAR decreases by 11.4%, whereas Gemini-1.5 Pro’s FAR increases instead.
Speaker Preference. To investigate how social engineering techniques affect speaker verification systems, we designed an experiment in which models were required to choose between two audio samples and identify the one more likely to belong to a legitimate user. The results show that SALMONN has a 39% tendency to select the audio from the genuine user for authentication. In contrast, the GPT-4o series of models consistently favors the audio from the attacker employing social engineering techniques, treating it as belonging to an authenticated user.
Appendix J Background and Related work
J.1 Audio Large Language Models
With the rapid increase in parameter and data scales, text-only large language models (LLMs) have achieved groundbreaking progress in language understanding and generation, as exemplified by models such as GPT-4 and the Gemini series [1, 66]. Building on this, researchers explored cross-modal alignment by integrating visual information into unified representation spaces. This led to early models like CLIP [57] and Flamingo [3], and later, models such as GPT-4V and Gemini capable of processing high-resolution images and long contexts. Recently, ALLMs have further expanded the input modalities by incorporating temporal acoustic features (such as Mel-spectrograms, log-power spectra, or variable-length waveforms) for joint modeling with semantic tokens [86]. In contrast to the visual modality, audio signals exhibit high dynamic range and transient variations in both time and frequency domains. Consequently, most ALLMs adopt separate time-frequency encoders or discretizing acoustic tokenizers to capture rich attributes such as timbre, rhythm, and scene [27, 18]. Representative models include Qwen2-Audio with its pipeline-style natural language prompt pre-training [14], SALMONN with a unified "auditory-language-music" framework [65], and WavLLM with a dual-encoder plus Prompt-LoRA adaptation mechanism [27]. After cross-modal alignment, these models demonstrate strong capabilities in content and scene understanding, enabling applications such as spoken question answering, music style analysis, and environmental sound event retrieval. They also show great promise in medical diagnosis (e.g., detection of respiratory diseases, analysis of heart sounds), voice control for smart homes, and multimedia generation and editing [99, 59, 5].
However, the multimodal nature of ALLMs also introduces new trustworthiness challenges. First, since the models are trained on large-scale acoustic-text paired corpora, they are prone to memorizing and leaking sensitive user speech information, and are therefore vulnerable to privacy attacks such as membership inference [67, 26]. Second, adversarial audio can exploit inaudible ultrasound or fine-grained perturbations to mislead ALLMs: early work such as DolphinAttack [96] and Vrifle [38] demonstrated covert manipulation of voice assistants via inaudible commands injected with ultrasonic carriers above 20 kHz [104, 36, 92]; recently, AdvWave systematically proposed gradient shattering repair and two-stage optimization, achieving over 40% jailbreak success rates on various ALLMs [30]. In addition, large-scale multimodal models are similarly susceptible to cross-modal instruction injection and protocol mismatching attacks, potentially leading to unauthorized content generation [43], privilege escalation [23], and even physical harm [45]. When integrated into voice-interface agentic systems, trustworthiness challenges are amplified and become paramount [40, 91]. To address these risks, the community has proposed a range of safety, security, and privacy mechanisms, including SafeEar, an empirical content privacy-preserving audio deepfake detection framework [37] and active detection with post-hoc rejection [35] differentially private pre-training, segment-wise gradient compression defenses. Nevertheless, in real-time voice scenarios, these approaches still face detection latency and robustness trade-offs, highlighting the urgent need for further research.
J.2 Audio Large Language Model Benchmarks
Current evaluations of ALLMs have primarily focused on their performance in fundamental tasks. SUPERB [87] first introduced a unified evaluation framework for speech processing, where self-supervised speech representation models are assessed across ten downstream tasks, including phoneme recognition, keyword spotting, speaker verification, and emotion recognition. This benchmark demonstrates the generality and effectiveness of SSL representations in diverse scenarios. Subsequently, SUPERB-SG [68] extended this framework to encompass advanced semantic understanding and generative tasks, such as speech translation [69], voice conversion [51], speech separation [73], and enhancement [7], in order to further evaluate models’ generative abilities and robustness. SLURP [6] provides a large-scale dataset and evaluation framework targeting spoken language understanding, thereby enabling a comprehensive comparison between end-to-end and pipeline approaches, while SLUE [61] assesses complex tasks including audio question answering, summarization, and named entity recognition within realistic speech scenarios with low-resource context, highlighting the impact of ASR models on downstream task performance. In the field of audio captioning, AudioCaps [32] and Clotho [17] serve as major evaluation benchmarks, with Clotho-AQA [39] pioneering a real-world dataset for audio question answering, facilitating the evaluation of models’ semantic reasoning capabilities. The recently released AIR-Bench [86] categorizes evaluation tasks into two dimensions: fundamental abilities and dialogic abilities, covering a wide variety of audio types such as speech, environmental sounds, and music. The fundamental dimension comprises 19 specific tasks, whereas the dialogic dimension uses open-ended question-answering formats to evaluate generative performance of models under diverse and mixed audio backgrounds. These benchmarks offer diverse and comprehensive frameworks for evaluation and comparison of ALLMs, yet they mainly focus on fundamental performance; systematic assessments of safety, ethical risks, and social impacts remain insufficient.
Existing safety evaluation benchmarks are relatively limited, with most focusing on multimodal scenarios or specific attack methods. For example, MM-SafetyBench [41] proposed an evaluation framework for image query attacks targeting multimodal LLMs, collecting 5,040 text-image pairs to assess model safety under image manipulation. SafeBench [80] constructed 23 risk scenarios and 2,300 multimodal harmful example pairs by automatically generating harmful multimodal queries, and designed a collaborative LLM review protocol to enhance evaluation reliability. In the audio domain, the Chat-Audio Attacks (CAA) benchmark [88] designed four types of audio attacks for dialog audio attack evaluation, and adopted a synthesis of standard evaluation, GPT-4o-based assessment, and human evaluation strategies to measure model robustness. The study [85] comprehensively evaluated the safety performance of five audio multimodal models via red-teaming against harmful audio, textual interference, and specific jailbreak attacks, revealing attack success rates as high as 70%. Furthermore, the SEA method [44] proposed a synthetic embedding augmentation approach for safety alignment, verifying the feasibility of aligning audio safety in multimodal models using only textual data. Although the above benchmarks have made progress in their respective areas, there is still a lack of a unified audio safety benchmark that comprehensively considers multidimensional risks such as fairness, hallucination detection, privacy protection, robustness, and authentication. Therefore, this work proposes the AudioTrust benchmark, which encompasses six core directions: fairness evaluation, hallucination detection, safety defense, privacy leakage, robustness challenges, and identity authentication. By combining scenario-driven question-answer pairs with GPT-4o automated evaluation, AudioTrust reveals the safety boundaries of ALLMs in high-risk environments, thereby providing systematic guidance for the secure and trustworthy deployment of future models.
Appendix K Limitations
While AudioTrust offers a pioneering and comprehensive framework for the multidimensional trustworthiness evaluation of Audio Large Language Models (ALLMs), certain limitations warrant consideration. Firstly, the datasets, though meticulously constructed to cover a diverse range of scenarios across fairness, hallucination, safety, privacy, robustness, and authentication, are necessarily finite and may not encapsulate the full spectrum of real-world complexities or all potential adversarial manipulations. Secondly, the dynamic nature of ALLM development and emerging threat landscapes also means that any benchmark, including AudioTrust, represents a snapshot in time and will require continuous updates to remain relevant and comprehensive in assessing the evolving trustworthiness of these rapidly advancing systems.
Appendix L Social Impact
The introduction of AudioTrust carries significant positive social implications by fostering the development and deployment of more trustworthy ALLMs. By systematically evaluating fairness, AudioTrust aims to mitigate the perpetuation of harmful societal stereotypes related to gender, race, age, accent, and other sensitive attributes in critical applications like recruitment, admissions, and financial loan evaluations. Exposing and quantifying biases in ALLMs can drive research towards debiasing techniques, ultimately promoting more equitable outcomes and reducing discrimination facilitated by AI systems. The focus on hallucination detection is crucial for enhancing the reliability of ALLMs; by identifying tendencies to generate physically impossible, logically inconsistent, or factually incorrect information, AudioTrust encourages the development of models that provide more accurate and dependable responses. This is particularly vital in high-stakes environments such as emergency response or medical information provision, where hallucinations could have severe consequences.
The safety evaluation component addresses the urgent need to prevent ALLMs from being exploited for malicious purposes, such as generating harmful content, guiding illegal activities, or bypassing guardrails in enterprise, financial, and healthcare systems. By providing a structured way to test against jailbreak attempts and emotional deception, AudioTrust contributes to building more resilient systems that can resist manipulation and adhere to ethical guidelines. Similarly, the privacy dimension of AudioTrust highlights risks of unintentional information disclosure and inference of sensitive attributes from audio. This awareness can lead to the design of ALLMs with stronger privacy-preserving mechanisms, safeguarding user data and fostering greater user trust in voice-interactive technologies. Evaluating robustness against various audio disturbances—ranging from background noise and multi-speaker environments to adversarial attacks—ensures that ALLMs can maintain performance integrity in realistic, imperfect conditions, which is essential for their practical adoption in everyday life and critical infrastructures. Finally, the authentication assessments address vulnerabilities to voice cloning and spoofing, thereby contributing to more secure voice-based access control systems and protecting individuals and organizations from identity-related fraud.
Collectively, AudioTrust serves as a catalyst for responsible innovation, providing developers, policymakers, and the public with crucial insights into the trustworthiness of ALLMs, and guiding the community towards creating AI technologies that are not only powerful but also fair, reliable, safe, private, robust, and secure for societal benefit. It establishes a foundational benchmark that can inform future standards and best practices for trustworthy AI in the audio domain.
Appendix M Data sheet
We follow the documentation frameworks provided by [81].
M.1 Motivation
For what purpose was the dataset created?
The AudioTrust dataset was created to serve as a large-scale benchmark for evaluating the multi-faceted trustworthiness of Multimodal Audio Language Models (ALLMs). It aims to help the research community better understand the capabilities, limitations, and potential risks associated with deploying these state-of-the-art AI models.
The benchmark examines model behavior across the following six critical dimensions:
Hallucination: Fabricating content unsupported by audio.
Robustness: Performance under audio degradation.
Authentication: Resistance to speaker spoofing/cloning.
Privacy: Avoiding leakage of personal/private content.
Fairness: Consistency across demographic factors.
Safety: Generating safe, non-toxic, legal content.
M.2 Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?
Yes. The AudioTrust dataset is publicly released and accessible to third parties.
How will the dataset be distributed (e.g., tarball on website, API, GitHub)?
The dataset is available on the Hugging Face Datasets Hub and can be loaded using the https://huggingface.co/datasets/JusperLee/AudioTrust.
The associated code, scripts, and benchmark framework are hosted on GitHub (https://github.com/JusperLee/AudioTrust).