### v4

# AudioTrust: Benchmarking the Multifaceted Trustworthiness of Audio Large Language Models

Abstract
The rapid development and widespread adoption of Audio Large Language Models (ALLMs) require a rigorous assessment of their trustworthiness. However, existing evaluation frameworks, primarily designed for text, are not equipped to handle the unique vulnerabilities introduced by audio’s acoustic properties. We find that significant trustworthiness risks in ALLMs arise from non-semantic acoustic cues, such as timbre, accent, and background noise, which can be used to manipulate model behavior. To address this gap, we propose AudioTrust, the first framework for large-scale and systematic evaluation of ALLM trustworthiness concerning these audio-specific risks. AudioTrust spans six key dimensions: fairness, hallucination, safety, privacy, robustness, and authenticition. It is implemented through 26 distinct sub-tasks and a curated dataset of over 4,420 audio samples collected from real-world scenarios (e.g., daily conversations, emergency calls, and voice assistant interactions), purposefully constructed to probe the trustworthiness of ALLMs across multiple dimensions. Our comprehensive evaluation includes 18 distinct experimental configurations and employs human-validated automated pipelines to objectively and scalably quantify model outputs. Experimental results reveal the boundaries and limitations of 14 state-of-the-art (SOTA) open-source and closed-source ALLMs when confronted with diverse high-risk audio scenarios, thereby offering critical insights into the secure and trustworthy deployment of future audio models. Our platform and benchmark are publicly available at this link.
1 Introduction
Despite rapid progress in ALLMs (Du et al., 2023; Tang et al., 2024; Chu et al., 2024; Zhang et al., 2023a; 2024a; Xu et al., 2025c; Yao et al., 2024), there is still no comprehensive benchmark evaluating their potential risks. Existing safety evaluation frameworks, such as SafeDialBench (Cao et al., 2025) and SafetyBench (Zhang et al., 2024b), mainly focus on the text modality while overlooking the unique properties or application scenarios of audio. Unlike text, audio features complex temporal-frequency patterns, rich emotions, and contextual dependencies, and introduces additional challenges, such as speech privacy, speaker recognition, and environmental acoustic analysis. These factors lead to unique trustworthiness risks in audio, including gender and accent biases, audio hallucinations, social engineering attacks, personal information leakage, and adversarial attacks on audio systems.
The integration of audio modalities into large models, while functionally powerful, introduces a new attack surface and exacerbates existing trustworthiness vulnerabilities. To systematically quantify these emergent risks, we introduce AudioTrust, the first comprehensive benchmark designed to evaluate the trustworthiness of ALLMs. AudioTrust establishes a rigorous evaluation framework across six critical dimensions where audio introduces unique safety concerns: (1) Fairness: Evaluating biases derived from vocal delivery, rather than the semantic content of speech; (2) Hallucinations: Testing for audio-grounded hallucinations, where model outputs violate the physical laws or temporal logic of an acoustic scene. (3) Safety: Assessing resilience to harmful queries that leverage persuasive or emotional vocal tones to bypass safety filters; (4) Privacy: Quantifying the leakage of sensitive information from spoken content and inference of personal attributes from acoustic cues; (5) Robustness: Assessing model performance in acoustically complex and imperfect environments, such as those with background noise. (6) Authentication: Evaluating the ability to distinguish authentic speakers from sophisticated impersonation attacks, including voice clones and audio-based social engineering. Underpinning our benchmark is a curated dataset of over 4,420 audio samples spanning 18 distinct evaluation tasks, from emergency communications to adversarial settings. We deploy a large-scale automated evaluation pipeline to ensure rigorous and reproducible assessment. The reliability of our automated metrics and results is verified by human experts (over 97% agreement rate). The initial findings on representative models are summarized in a public leaderboard (see Figure˜2(b)). Details of the benchmark are provided in Appendix˜C and Figure˜5.
Fairness. The introduction of audio inputs brings new fairness risks by introducing new biases linked to audio characteristics. To investigate these risks, we conducted a comprehensive evaluation of model fairness along two dimensions: decision-making experiments and stereotype-association experiments. Our main findings are as follows: (1) Audio-based attributes (e.g., accent, emotion) can introduce biases that are stronger than those from traditional sensitive attributes (e.g., age, gender), indicating that audio information is a key carrier of bias; (2) We observed that closed-source models exhibited stronger decision biases, while open-source models were more susceptible to stereotype associations; (3) The identified biases tend to disadvantage non–socially-dominant groups, such as older-sounding accents, perceived calmness, and markers of lower socioeconomic status (SES). Further details are provided in Section˜2.
Hallucination: The introduction of audio gives rise to new forms of hallucination, including the misinterpretation of paralinguistic features (e.g., emotion or accent) and failures to capture temporal causality within speech. We study these vulnerabilities on a carefully curated benchmark and identify several key weaknesses. (1) Closed-source models exhibit stronger robustness when confronted with acoustically implausible events. (2) Many models remain vulnerable to misleading meta-attributes, revealing insufficient alignment with domain knowledge. (3) We further observe pronounced fragility in tasks that require temporal reasoning. (4) Substantial variability exists across models in terms of cross-modal semantic consistency. Further details are provided in Section˜3.
Safety: Incorporating audio inputs substantially broadens the attack surface. Unlike text, speech carries emotional, contextual, and anthropomorphic cues that adversaries can exploit by modulating tone, injecting affective signals, or impersonating identities. We design a composite attack framework spanning emotion-driven deception and identity-verification evasion. Empirical analysis shows: (1) closed-source models exhibit stronger overall robustness but remain sensitive to highly emotional speech; (2) open-source models are disproportionately vulnerable to identity- and emotion-based manipulation, often yielding unsafe outputs in high-stakes settings such as healthcare; (3) for prompts concerning illicit guidance, closed-source systems largely resist, whereas some open-source models deliver risky recommendations in multi-turn audio dialogues. Further details are provided in Section˜4.
Privacy: ALLMs face two closely related but mechanistically distinct audio privacy risks. The first is content-level leakage, such as reading out and repeating bank account numbers, social security numbers, or addresses. The second is paralinguistic-level inference leakage, where attributes such as age, gender, race, geographic location, or socioeconomic status are inferred from voiceprints, timbre, intonation, accent, or background sounds. To conduct a systematic evaluation, we created targeted scenarios to assess both explicit information disclosure and implicit attribute inference. Our findings show: (1) ALLMs are relatively robust in preventing direct content leakage; (2) existing semantic-oriented defenses fall short in addressing paralinguistic attack surfaces unique to audio, underscoring the need to integrate acoustic and environmental cues into privacy-aware decision boundaries. See Section˜5 for details.
Robustness: Since ALLMs interact directly through audio, they are inevitably affected by noise and distortion. To systematically characterize their robustness, we evaluated the models against a comprehensive suite of real-world audio degradations, including environmental noise, speaker overlap, and signal perturbations. Our analysis reveals: (1) mainstream closed-source ALLMs achieve stronger task performance under overlapping speech, non-stationary noise, and reverberant conditions, while most open-source models exhibit substantial performance degradation; (2) existing ALLMs generally demonstrate an “over-textualization” tendency, where models continue reasoning based on partially correct transcripts while neglecting acoustic cues when transcription is correct but acoustic attribution is mistaken. See Section˜6 for details.
Authentication: In applications of ALLMs, speech-related authentication issues are particularly critical. To investigate these risks, we evaluated the models against several key attack vectors, including identity verification bypass and voice cloning deception. The results show: (1) certain closed-source models exhibit some resilience in identity verification scenarios, whereas open-source models are generally more vulnerable to sophisticated voice-based attacks; (2) adversaries may exploit social engineering or acoustic interference, such as background crowd noise, to compromise verification reliability; (3) employing more stringent speech-text prompting strategies can substantially improve the ability of ALLMs to withstand voice cloning attacks. See Section˜7 for details.
2 AudioTrust: Fairness
This section examines the fairness issues associated with ALLMs. Fairness risks in audio models are fundamentally different from those in text or vision systems. For instance, a text-based model might exhibit bias based on a name mentioned in a hiring application, but an ALLM can develop biases from the acoustic cues of an applicant’s voice alone. A hesitant speaking style could be misinterpreted as a lack of confidence, or a non-native accent could trigger stereotypes, regardless of the spoken content’s quality. Traditional fairness metrics focusing on textual protected attributes are insufficient to capture these audio-native biases. We investigate these new audio forms of bias that arise from how auditory characteristics are perceived.
Attack Strategies. To systematically probe these risks, we categorize fairness into two dimensions: traditional fairness and audio-based fairness. Traditional fairness assesses biases linked to demographic attributes like gender, race, and age (Cho et al., 2023; Zhang et al., 2017b; Wan et al., 2024) that can be inferred from a voice. We test, for example, if a model’s loan approval decisions are skewed by whether the applicant sounds male or female. Audio-based fairness isolates biases triggered by instrinsic acoustic properties. We divide these biases into four sub-categories: accent, linguistic characteristic (e.g., speech disfluencies may adversely impact the model’s fairness), economic status (e.g., noisy environments, often correlated with lower economic status, might yield more negative outputs), and personality traits (e.g., negative emotions may cause the model to produce negative outputs) (see Figure 6). For each sensitive attribute, we designed decision-making and stereotype-driven scenarios (Xu et al., 2025a; Wang et al., 2025c). Decision-making covers recruitment, admissions, and loan evaluations. Stereotype-driven contexts address beliefs such as men outperforming women in mathematics and gender roles in medicine and occupations. During the data construction phase, we utilized the multi-modal generation model GPT-4o (OpenAI et al., 2024) to produce textual content, which was then transformed into audio using the F5-TTS model (Chen et al., 2025). The entire data generation process begins with a text-based prompt, constructed entirely by GPT-4o to support natural language fluency and semantic consistency. We constructed a dedicated dataset consisting of 840 audio samples, each lasting approximately 20 seconds. The samples were annotated with seven key sensitive attributes: gender, age, and race for evaluating traditional fairness, as well as accent, linguistic characteristic, economic status, and personality traits for assessing audio-baseed fairness. Notably, due to the limitations of the audio modality, attributes such as appearance are not included (Nangia et al., 2020; Parrish et al., 2022). Detailed dataset construction procedures can be found in Sections˜D.1 and D.2.
Evaluation & Metrics. For fairness evaluation, we introduced a key metric: Group Fairness Score (Xu et al., 2025a; Deldjoo & Nazary, 2024). Group fairness examines the distributional equity of model outputs across different groups (e.g., male versus female), measured by the divergence or balance between group distributions. indicates perfect fairness, while indicates extreme unfairness. In computing the fairness score, we used human annotation and counting throughout to ensure data correctness and validity. For detailed formulas in the Section˜D.3.
| Model | ||
| Open-source Models | ||
| SALMONN | 0.139 | 0.089 |
| Ultravox | 0.238 | 0.392 |
| Qwen2-Audio | 0.333 | 0.290 |
| MiniCPM-o 2.6 | 0.260 | 0.415 |
| Step-Fun | 0.658 | 0.505 |
| Qwen2.5-Omni | 0.067 | 0.202 |
| Kimi-Audio | 0.036 | 0.086 |
| OpenS2S | 0.017 | 0.157 |
| Step-Audio2 | 0.074 | 0.098 |
| Closed-source Models | ||
| Gemini-1.5 Pro | 0.297 | 0.460 |
| GPT-4o Audio | 0.926 | 0.264 |
| GPT-4o mini Audio | 0.864 | 0.245 |
| Gemini-2.5 Flash | 0.370 | 0.246 |
| Gemini-2.5 Pro | 0.319 | 0.205 |
| Average | 0.328 | 0.261 |
Note: : higher than column average, : lower than column average, subscript is signed difference from mean.
Results. We evaluated the group fairness of 14 ALLMs in terms of social stereotypes and decision-making in Table 1. Complete results and examples are provided in Section˜D.4. The main findings are as follows: (1) Existing ALLMs exhibit severe unfairness across different sensitive attributes, falling far short of the ideal fairness (i.e., ). (2) The GPT-4o series shows a pronounced disparity between decision and stereotype . This is because we have designed extreme decision scenarios, and the GPT-4o series models sacrifice fairness to maintain accuracy in response. (3) Although leading closed-source models such as GPT-4o Audio exhibit stable fairness performance, open-source models vary widely. Notably, Step-Fun demonstrates strong fairness, with scores comparable to the best closed-source models. By contrast, models like OpenS2S and SALMONN display pronounced vulnerabilities, underscoring a substantial capability gap within the open-source ecosystem. (4) The two models in the Step series exhibit a stark disparity in fairness, suggesting substantial differences in their underlying fairness mechanisms.
3 AudioTrust: Hallucination
In this section, we examine the hallucination problem in ALLMs. Audio hallucinations extend beyond the factual errors seen in text (Ji et al., 2023; Sriramanan et al., 2024; Yao et al., 2023). An ALLM does not just process information; it interprets a simulated physical world. For instance, if an audio recording contains the sound of a gunshot followed by a body falling, but the model describes the fall happening first, it is not just a factual error, which is a violation of causality. Similarly, describing a fire burning underwater is a violation of physical laws. These audio-grounded errors are undetectable by text-based fact-checking and pose unique safety risks. Our work is the first to systematically define and evaluate these physically and logically grounded hallucinations.
Attack Strategies. In AudioTrust, we identify two main categories of audio hallucinations (see Appendix E.1): Physical Logic and Chronological Order. The former relates to violations of acoustic laws and environment properties, and the latter reflects failures in reasoning about temporal and causal relations. These distinctions directly connect to safety risks in real-world use. To evaluate them, we built a dataset of 320 samples from synthetic and real sources, According to the different logical or temporal requirements of the testing tasks, the clips and sequences are edited and spliced to varying degrees and in different orders (see Appendix E.2). For Acoustic-Physical Hallucinations, we focus on two specific manifestations: (1) Contravention of Physicochemical Constraints, generating impossible events (e.g., the flames are burning in the seawater) to test propagation understanding (Liu et al., 2024); (2) Source–Environment Mismatch, applying contradictory reverberation (e.g., casual speech with cathedral acoustics) to test disentanglement of source vs. environment (Zhou et al., 2022). For Temporo-Logical Hallucinations, we examine (1) Temporal-Causal Inversion: reversing causal chains (e.g., engine start before ignition) to probe event logic (Chen et al., 2023); (2) Cross-Modal Contradiction: pairing audio with conflicting text (e.g., fast footsteps described as peaceful rain) to test cross-modal reasoning (Deshmukh et al., 2023).
Evaluation & Metrics. We introduce a comprehensive evaluation framework to assess model capabilities across four key dimensions: hallucination detection, attribute verification, real-world consistency, and transcription accuracy. For hallucination detection, models are required to identify inaccuracies in audio-text pairs and provide justifications. Performance is quantified via a multi-dimensional, GPT-4o-based (OpenAI et al., 2024) and Qwen3-based (Yang et al., 2025a) evaluators.The scores mentioned above range from 0 to 10 and are used to evaluate the accuracy of the detection and the quality of the interpretation. These scores were then subject to further review by the human evaluators. We probe for attribute-level hallucinations related to physical properties, labels, and content using multi-stage prompting (Madaan et al., 2023). To measure alignment with factual information, we adopt the two-stage protocol from Li et al. (2023a) for real-world consistency assessment. Finally, we evaluate transcription robustness under hallucinatory interference using both the standard Word Error Rate (WER) (Deshmukh et al., 2023) and the cross-modal WER (CM-WER) (Tang et al., 2024). The complete experimental design and metric details are provided in Appendices E.3.1 and E.3.2.
| Model | CM | LM | LV | PV |
| Open-source Models | ||||
| MiniCPM-o 2.6 | 6.24 / 4.62 | 6.20 / 5.13 | 8.28 / 7.53 | 6.13 / 5.30 |
| Qwen2-Audio | 8.15 / 8.66 | 4.34 / 7.05 | 7.26 / 5.48 | 7.77 / 8.18 |
| SALMONN | 2.65 / 1.70 | 1.22 / 1.16 | 6.64 / 5.97 | 3.98 / 3.92 |
| Ultravox | 5.74 / 8.29 | 4.52 / 7.83 | 8.01 / 6.10 | 8.34 / 8.70 |
| Qwen2.5-Omni | 8.12 / 8.44 | 5.63 / 3.94 | 7.89 / 6.45 | 6.11 / 6.17 |
| Step-Fun | 3.96 / 3.93 | 4.84 / 5.04 | 5.80 / 5.83 | 8.72 / 9.25 |
| Kimi Audio | 1.86 / 1.88 | 5.77 / 5.67 | 5.82 / 7.22 | 8.54 / 8.62 |
| Step-Audio2 | 3.62 / 1.74 | 1.94 / 4.09 | 2.60 / 6.10 | 2.76 / 8.39 |
| OpenS2S | 1.92 / 1.30 | 5.03 / 4.11 | 5.97 / 5.75 | 7.89 / 8.27 |
| Closed-source Models | ||||
| Gemini-1.5 Pro | 8.41 / 8.05 | 7.81 / 7.17 | 8.66 / 8.35 | 8.87 / 9.78 |
| Gemini-2.5 Flash | 7.98 / 7.15 | 8.36 / 8.24 | 8.71 / 9.03 | 8.57 / 9.74 |
| Gemini-2.5 Pro | 8.19 / 7.02 | 8.78 / 5.08 | 8.70 / 9.00 | 8.49 / 9.70 |
| GPT-4o Audio | 3.94 / 1.65 | 2.68 / 1.68 | 3.53 / 3.41 | 8.79 / 9.43 |
| GPT-4o Mini Audio | 2.34 / 1.35 | 1.21 / 1.77 | 1.24 / 3.25 | 9.00 / 9.61 |
| Average | 5.22 / 4.70 | 4.88 / 4.85 | 6.36 / 6.39 | 7.43 / 8.22 |
Note: Values are shown as "GPT-4o / Qwen3 evaluators. Subscript is the absolute difference from the respective column average. Scenarios: Content Mismatch (CM), Label Mismatch (LM), Logical Violation (LV), and Physical Violation (PV).
Results. As shown in Table 2, our evaluation highlights both the progress and the critical limitations of current ALLMs in resisting hallucinations. Complete results and examples are provided in Section˜E.4 and Section˜E.5. We observe two main findings: (1) Although certain open source models, such as Gemini-2.5, demonstrate the ability to detect specific types of hallucinations, particularly those with explicit physical or temporal contradictions.(e.g., statements claiming an object exists in two distinct places at the same time or a water bottle made a sound by hitting the ground during its fall) Nonetheless, overarching vulnerabilities persist. Most models falter on subtler instances , including source-environment incongruities (e.g., an audio track describing a visual scene that contradicts the dialogue content) or cross-modal semantic discrepancies (e.g. The audio background information of a scene contradicts the content of the dialogue.), indicating that their perceptual understanding remains fragmented lacking an integrated cognitive architecture. (2) A striking observation is the negative correlation between the subjective complexity of tasks from a human perspective and the actual performance of models. While models attain comparatively high accuracy in identifying violations of physical laws, they underperform on content mismatches such as scenarios and independent tasks in different scenarios that humans intuitively discern with ease. This divergence highlights a core disparity in auditory perception and reasoning between humans and machines: models excel at low-level acoustic anomaly detection but struggle to emulate human-like commonsense reasoning.
4 AudioTrust: Safety
The safety landscape for ALLMs presents challenges distinct from text-based systems (Wei et al., 2023; Wang et al., 2023b). The tone of a voice (whether urgent, distressed, or authoritative) can serve as a powerful tool to bypass the model’s safety alignment. For example, a user might pretend to have a medical emergency with a panicked voice to request dangerous information. This emotional attack vector, which leverages the persuasive nature of the human voice, is a novel challenge that text-only safety protocols do not address. Due to the high-dimensional and complex nature of these acoustic features, they may introduce new vulnerabilities during the encoding process and are challenging to align effectively with the safety constraints established for the text modality.
Attack Strategies. To systematically evaluate the safety of ALLMs, AudioTrust is structured along two orthogonal dimensions: Domain-Specific Policy Jailbreak Attacks and General Illegal Activity Guidance. The first category, Domain-Specific Policy Jailbreak Attacks, encompasses adversarial attempts to circumvent a model’s intrinsic behavioral constraints through techniques like prompt injection and role-playing (Li et al., 2025; Mo et al., 2024; Zhou et al., 2024). To ensure real-world applicability, these attacks are stratified into three critical domains: enterprise, financial, and healthcare systems, each populated with 100 test instances. The second category, General Illegal Activity Guidance, assesses the model’s propensity to generate harmful content for illicit requests (e.g., fraud, weapon crafting), drawing upon established benchmarks (Cao et al., 2025; Zhang et al., 2024b) to assemble 300 diverse prompts. In AudioTrust, we specifically investigated an attack vector based on the audio modality: Emotional Deception, which involves using speech prompts with specific emotional cues (e.g., anger, sadness) to induce the model to bypass its safety filters. All textual prompts were augmented via GPT-4o (OpenAI et al., 2024) and subsequently synthesized into audio using F5-TTS (Chen et al., 2025). Further details on data classification and construction are available in Appendices F.1 and F.2, respectively.
| Model Name | Jailbreak | Illegal Activities Guidance | ||
| Enterprise | Financial | Medical | ||
| Open-source Models | ||||
| SALMONN | 74.2 / 83.4 | 74.4 / 91.2 | 80.8 / 92.2 | 77.1 / 89.4 |
| Ultravox | 97.2 / 97.0 | 83.8 / 97.0 | 90.8 / 99.2 | 98.0 / 96.9 |
| Qwen2-Audio | 68.2 / 95.2 | 80.6 / 91.8 | 81.4 / 90.2 | 92.5 / 93.3 |
| MiniCPM-o 2.6 | 76.2 / 83.0 | 79.2 / 86.0 | 81.6 / 94.8 | 96.2 / 99.3 |
| Step-Fun | 70.6 / 92.4 | 86.2 / 92.8 | 89.0 / 94.8 | 94.5 / 94.2 |
| Qwen2.5-omni | 97.2 / 99.2 | 94.8 / 99.4 | 94.2 / 97.8 | 99.1 / 99.6 |
| Kimi-Audio | 99.4 / 99.8 | 98.2 / 100.0 | 95.2 / 99.6 | 99.9 / 99.9 |
| OpenS2S | 51.4 / 47.6 | 67.8 / 87.4 | 75.2 / 83.0 | 71.8 / 72.9 |
| Step-Audio2 | 88.0 / 100.0 | 68.4 / 99.8 | 73.0 / 96.8 | 69.0 / 99.6 |
| Closed-source Models | ||||
| Gemini-1.5 Pro | 99.0 / 94.0 | 99.2 / 97.2 | 97.6 / 99.0 | 99.9 / 99.9 |
| GPT-4o Audio | 99.0 / 99.2 | 99.2 / 100.0 | 98.8 / 100.0 | 100.0 / 99.9 |
| GPT-4o mini Audio | 99.8 / 97.6 | 99.0 / 100.0 | 98.8 / 99.0 | 99.9 / 99.9 |
| Gemini-2.5 Flash | 100.0 / 100.0 | 99.8 / 93.2 | 99.4 / 96.0 | 99.8 / 93.6 |
| Gemini-2.5 Pro | 99.8 / 82.6 | 99.4 / 87.2 | 99.4 / 100.0 | 99.8 / 99.0 |
Note: Due to the common issue of random audio recognition failures in open-source models, these scores may be inflated. Values are shown as GPT-4o / Qwen3 evaluators.
Evaluation & Metrics. Quantifying the success of jailbreak attacks and the harmfulness of generated content via rule-based methods is notoriously challenging. We therefore employed a model-based evaluation, leveraging GPT-4o and Qwen3. This choice is motivated by its strong alignment with human preferences on safety-sensitive assessments (Tan et al., 2025). We first generated text-based malicious prompts and converted them into audio; specific emotional expressions were achieved by selecting reference audio clips with distinct emotional timbres during the TTS process. In particular, we selected TTS voices and configurations that realize the same sentence with different emotional tones and different genders. GPT-4o and Qwen3 rated each model-generated response on a 5-point Likert scale, where a score of 5 indicates complete refusal of a malicious request, and 1 denotes a fully compliant, harmful generation. Subsequently, these scores were further reviewed by human evaluators. Based on these scores, we computed the primary metric: the Defense Success Rate (DSR), which measures the model’s robustness against jailbreak attempts (higher is better). The detailed evaluation protocols, which are adapted for each task, are provided in Sections˜F.3 and F.4.
Results. We evaluated the safety performance of both open- and closed-source ALLMs across different scenarios, as shown in Table˜3. Our analysis reveals several observations: (1) While leading closed-source models such as Gemini-2.5 Flash maintain strong safety performance, open-source models exhibit substantial variation. Notably, Kimi-Audio demonstrates remarkable robustness, achieving scores comparable to the best closed-source counterparts. In contrast, models such as OpenS2S and SALMONN display considerable vulnerability, highlighting the large capability gap within the open-source ecosystem. (2) For closed-source models, the medical domain remains relatively more susceptible to jailbreak attacks, suggesting that domain-specific alignment in specialized areas is still an open challenge even for highly capable systems. In open-source models, no single domain consistently emerges as the weakest link, with vulnerabilities appearing to be model-dependent. (3) Most models, regardless of being open-source or closed-source, generally exhibit stronger defenses against General illegal activity guidance prompts compared to domain-specific jailbreak attempts. This indicates that broad safety training against overtly illegal content is largely effective, whereas nuanced, domain-targeted jailbreaks remain a more successful pathway for adversaries. Detailed results are provided in Section˜F.5.
5 AudioTrust: Privacy
This section examines privacy challenges specific to ALLMs. In text-based systems (Wang et al., 2023a; Huang et al., 2024), privacy risks typically involve the model memorizing and repeating sensitive information from its training data. ALLMs face this risk, but also a more subtle and pervasive one: information leakage from the acoustic signal itself. The sound of a voice can reveal a speaker’s approximate age; the background noise can betray their location (e.g., a quiet office or a busy cafe). This means ALLMs can infer private information even when it is never explicitly stated, creating a new class of privacy risks beyond simple content disclosure.
| Model | Direct leakage | Inference leakage | ||
| w/o | w/ | w/o | w/ | |
| Open-source Models | ||||
| SALMONN | 57.50 / 57.00 | 96.83 / 96.83 | 48.33 / 47.00 | 49.00 / 49.00 |
| UltraVox | 73.46 / 72.67 | 99.67 / 99.33 | 8.67 / 8.67 | 9.67 / 7.33 |
| Qwen2-Audio | 0.83 / 0.83 | 23.67 / 23.17 | 1.33 / 0.00 | 1.00 / 0.00 |
| MiniCPM-o 2.6 | 0.00 / 0.00 | 0.67 / 0.67 | 1.33 / 0.00 | 1.00 / 0.33 |
| Step Fun | 41.50 / 41.67 | 98.33 / 98.17 | 15.33 / 14.67 | 19.67 / 17.00 |
| Qwen2.5-Omni | 0.00 / 0.00 | 1.17 / 1.00 | 0.33 / 0.00 | 0.00 / 0.00 |
| Kimi Audio | 0.17 / 0.17 | 1.00 / 1.00 | 15.00 / 15.33 | 8.67 / 8.67 |
| OpenS2S | 7.68 / 7.17 | 43.83 / 42.83 | 8.67 / 6.87 | 7.33 / 7.83 |
| Step Audio2 | 0.00 / 0.00 | 38.83 / 38.83 | 0.00 / 0.00 | 0.00 / 0.00 |
| Closed-source Models | ||||
| Gemini-1.5 Pro | 11.85 / 12.12 | 98.50 / 98.67 | 5.33 / 6.33 | 9.33 / 8.67 |
| GPT-4o Audio | 92.00 / 91.83 | 99.67 / 99.50 | 6.67 / 6.33 | 20.00 / 17.33 |
| GPT-4o mini Audio | 100.00 / 100.00 | 100.00 / 100.00 | 14.00 / 13.33 | 40.00 / 40.00 |
| Gemini-2.5 Flash | 1.36 / 0.00 | 96.44 / 96.82 | 1.00 / 1.00 | 3.33 / 0.67 |
| Gemini-2.5 Pro | 33.50 / 33.33 | 94.17 / 94.33 | 0.33 / 0.00 | 0.67 / 0.00 |
| Average | 29.99 / 29.77 | 63.77 / 63.65 | 9.02 / 8.54 | 12.12 / 11.20 |
Note: Values are shown as GPT-4o / Qwen3 evaluators. Scores are refusal rates (higher is better). “w/o” vs. “w/” compares standard prompting without and with privacy-aware prompt engineering.
Attack Strategies. In AudioTrust, we categorize the privacy risks associated with ALLMs into two distinct groups: (1) Direct Privacy Leakage, which pertains to sensitive information explicitly stated within the conversational content. In this category, the ALLMs might reveal data such as a bank account number mentioned during a conversation. The formulation of this risk is informed by similar challenges in traditional large language models. (Wang et al., 2023a; Huang et al., 2024). (2) Privacy Inference Leakage, where private attributes are inferred from paralinguistic cues rather than the explicit content. Such cues include a speaker’s tone of voice, speech rate, accent, and vocal quality. This risk, which is unique to ALLMs, involves the model deducing personal attributes from the audio itself, independent of the semantic content. Consistent with recent studies (Xu et al., 2025a; Feng et al., 2023; Maltezou-Papastylianou et al., 2025), we treat the systematic inference of demographic attributes (e.g., age, ethnicity) from voice as a critical privacy leakage. While distinct from unique identity recognition, such profiling exploits paralinguistic cues to reveal sensitive personal data, which can be aggregated for user profiling at scale. Therefore, we evaluate demographic inference as a primary dimension of privacy risk. To evaluate these risks, we constructed two distinct datasets. The first, a direct privacy leakage set, contains 600 synthetic dialogues. This set was created by first synthesizing privacy information, then generating the conversations using LLMs, and finally transcribing them into audio. The second, a privacy inference set, comprises 300 samples. For this set, we used speech from Common Voice (Ardila et al., 2020) and mixed it with background audio sourced from freesound (Freesound, ). For a detailed taxonomy and construction methods, see Appendixes G.1 and G.2.
Evaluation & Metrics. We evaluated our experiment within an Audio Question Answering (AQA) framework, employing two distinct settings: one utilizing privacy-enhancing prompts and the other using standard prompts. GPT-4o and Qwen3 serve as the evaluators, assigning scores from 1 to 5 to each response and then the results were checked by human. A score of 1 indicates complete disclosure of private information, while a score of 5 signifies a refusal of the request due to privacy safety concerns. Our main evaluation metrics are refusal rate and accuracy. A higher refusal rate indicates stronger privacy protection. Accuracy is calculated only when privacy information is leaked. Further details on evaluation prompts and evaluation metrics can be found in Appendixes G.3 and G.4.
Results. We evaluated the privacy protection performance of both open-source and closed-source ALLMs. The results are shown in Table˜4. Our analysis revealed several key observations: (1) Direct Privacy Leakage: In experiments using the direct privacy leakage dataset, closed-source ALLMs generally achieved superior results. The GPT-4o series demonstrated the best refusal rate, reaching 100%. Furthermore, the implementation of privacy-enhancing prompts significantly improved performance across almost every model, showing approximately a 25% improvement. This indicates that traditional prompt engineering can effectively enhance privacy protection when sensitive information is explicitly present in conversation content. (2) Privacy Inference Leakage: Both open-source and closed-source ALLMs performed poorly in addressing privacy inference leakage. They rarely refused requests for certain types of privacy information, such as age and ethnicity, with the refusal rate of only 9.02%. Unlike direct privacy leakage, privacy-enhancing prompts had a minimal impact, yielding only about a 3% improvement. Across all our results, we observed that ALLMs struggle to process privacy information that is not directly stated in a conversation but rather inferred from paralinguistic cues. This suggests that ALLMs might not identify such inferred information as private or requiring protection. This could be due to the training process, where paralinguistic cues may have been considered less important than conversational content, or because of insufficient data for this specific type of information. Detailed results accuracy analysis are available in Appendix G.5.
6 AudioTrust: Robustness
This section investigates the robustness of ALLMs in maintaining performance against real-world audio distortions. Unlike in text (Wang et al., 2024b) and vision (Hendrycks & Dietterich, ), audio robustness presents unique challenges due to its physical nature. Audio signals are temporal encodings of acoustic phenomena, with inherent properties like source, medium, and reverberation shaping complex auditory scenes. A truly robust ALLM must be able to disentangle the primary speech signal from this acoustic clutter and maintain its performance. This is not simply a matter of better speech recognition, but a test of the model’s fundamental ability to function in imperfect acoustic environments.
| Model | AR | AQV | BC | ES | MS | NI |
| Open-source Models | ||||||
| MiniCPM-o 2.6 | 7.80 / 4.85 | 7.18 / 5.35 | 7.92 / 7.57 | 7.06 / 7.98 | 6.50 / 7.90 | 6.17 / 5.50 |
| Qwen2-Audio | 6.00 / 2.95 | 3.50 / 3.97 | 4.33 / 6.29 | 6.84 / 6.23 | 5.40 / 8.07 | 6.60 / 4.97 |
| SALMONN | 2.00 / 2.58 | 6.42 / 5.01 | 4.57 / 5.59 | 2.94 / 4.45 | 7.16 / 5.58 | 6.66 / 5.63 |
| Ultravox | 4.00 / 4.76 | 7.53 / 6.29 | 7.30 / 7.94 | 6.53 / 8.05 | 6.70 / 8.30 | 7.00 / 5.93 |
| Qwen2.5-Omni | 8.14 / 5.63 | 7.10 / 6.59 | 7.50 / 7.93 | 7.93 / 7.94 | 7.12 / 7.28 | 7.17 / 6.03 |
| Step-Fun | 5.00 / 4.78 | 7.48 / 6.50 | 8.20 / 8.36 | 7.42 / 7.46 | 5.89 / 4.97 | 7.08 / 6.86 |
| Kimi Audio | 5.67 / 4.00 | 6.83 / 5.63 | 6.00 / 5.70 | 6.83 / 5.00 | 7.08 / 6.41 | 6.94 / 5.98 |
| OpenS2S | 8.25 / 3.83 | 6.46 / 4.10 | 5.17 / 6.11 | 6.39 / 5.15 | 2.33 / 8.92 | 6.25 / 4.85 |
| Step-Audio2 | 6.18 / 4.93 | 6.58 / 6.45 | 7.92 / 7.91 | 6.82 / 7.61 | 0.00 / 0.12 | 6.78 / 7.25 |
| Closed-source Models | ||||||
| Gemini-1.5 Pro | 8.57 / 6.63 | 8.21 / 7.89 | 8.23 / 8.01 | 8.16 / 8.08 | 6.09 / 8.39 | 7.43 / 7.02 |
| Gemini-2.5 Flash | 8.16 / 6.98 | 8.38 / 7.89 | 8.28 / 8.14 | 7.93 / 7.74 | 6.36 / 8.12 | 7.76 / 7.54 |
| Gemini-2.5 Pro | 8.88 / 7.74 | 8.68 / 8.19 | 8.50 / 8.36 | 8.18 / 7.87 | 7.46 / 7.87 | 7.71 / 7.60 |
| GPT-4o Audio | 5.90 / 4.66 | 5.50 / 5.33 | 8.33 / 8.31 | 7.31 / 7.70 | 7.62 / 9.88 | 6.27 / 5.91 |
| GPT-4o mini Audio | 8.33 / 5.45 | 6.90 / 7.09 | 7.69 / 8.05 | 6.00 / 7.73 | 5.77 / 8.53 | 7.25 / 6.30 |
| Average | 6.63 / 4.98 | 6.91 / 6.16 | 7.14 / 7.45 | 6.88 / 7.07 | 5.82 / 7.17 | 6.93 / 6.24 |
Note: Values are shown as GPT-4o / Qwen3 evaluators. Subscript is the absolute difference from the respective column average. Scenarios: AR (Adversarial Robustness), AQV (Audio Quality Variation), BC (Background Conversation), ES (Environmental Sound), MS (Multiple Speakers), NI (Noise Interference).
Attack Strategies. We categorize robustness challenges for ALLMs into two primary types: intentional adversarial attacks (Madry et al., 2021) and naturally occurring phenomenon of performance degradation (Radford et al., 2023). Adversarial attacks employ carefully crafted, imperceptible perturbations to induce model failure (Carlini & Wagner, 2018) Details of the creation of additional interference data can be found in Section˜H.2). In contrast, non-adversarial challenges encompass common real-world interferences. We evaluate model robustness across several key dimensions: (1) adversarial resilience, including three categories: natural noise, speaker identification, and voice overlap situations (Bredin et al., 2020); (2) robustness to environmental noise (Zhang et al., 2023b) and variations in audio quality. To this end, we constructed dedicated datasets simulating these interferences. In order to ensure the fluency of natural language and the logical validity of the original speech, we created the dataset by adding the aforementioned simulated interference to the introduced big bench audio (Suzgun et al., 2023). Each dimension contains 40 multilingual and multi-topic samples to ensure a comprehensive assessment. Further dataset details are available in Sections˜H.1 and H.2.
Evaluation & Metrics. Given the challenges in directly measuring robustness or output risk, we adopt a model-based evaluation using GPT-4o (OpenAI et al., 2024) and Qwen3-based (Yang et al., 2025a), following recent evidence (Zheng et al., 2023a). Each test output is rated on a discrete 10-point scale, with scoring rubrics tailored per prompt and task: 10 indicates strong consistency with audio quality, while 0 means perceptual failure or inability to recognize the specified variation. These scores were then subject to further review by the human evaluators. Prompt templates are detailed in Section˜H.3.1. For comprehensive evaluation, we also report two quantitative metrics (Section˜H.3.2): CM-WER, measuring dissimilarity between generated and human-annotated transcriptions (Radford et al., 2023); and Content Consistency Rate (CCR),(e.g., transcribe the voice with added interfering information through multiple rounds of model dialogue prompts to understand its semantic context, and then score the transcription against the original voice text content.) assessing factual alignment between ALLMs outputs and ground-truth audio content (Min et al., 2023).
Results. Our robustness evaluation (Table 5) reveals a significant performance gap between closed- and open-source ALLMs. Detailed results are in Section˜H.3.3. (1) Superior robustness in closed source models. Such as the closed-source models like Gemini-2.5 Pro,a cross nearly all tested conditions including background noise, multi-speaker conversations, and audio quality variation leading closed source systems consistently outperform their open source counterparts. Notably, this advantage is most apparent under severe acoustic distortion, suggesting that proprietary models benefit from more mature front-end signal processing and advanced noise suppression architectures. (2) In contrast, many open-source systems experience a steep decline in transcription accuracy and semantic coherence when exposed to moderate noise or compression. Their audio encoders often fail to disentangle source speech from channel artifacts, leading to semantic hallucinations in which non-speech noise is incorrectly interpreted as meaningful content.
7 AudioTrust: Authentication
In this section, we investigate the reliability of ALLMs for authentication. Text-based authentication relies on semantic secrets like passwords. Audio authentication is more complex because a voice signal contains both a semantic component (the passphrase) and an acoustic one (the speaker’s unique voiceprint). This dual nature creates a unique attack surface. For example, an attacker could use a perfect AI-generated voice clone to speak a correct passphrase, defeating systems that rely on either modality alone. We evaluate how well ALLMs can defend against different impersonation attacks.
| Model Name | IVB | HS | VCS |
| Open-source Models | |||
| SALMONN | 26 / 26 | 7 / 8 | N/A |
| Ultravox | 95 / 96 | 57 / 57 | 28 |
| Qwen2-Audio | 42 / 39 | 71 / 71 | 92.5 |
| MiniCPM-o 2.6 | 24 / 20 | 43 / 39 | 79.5 |
| Step-Fun | 79 / 79 | 97 / 97 | 22 |
| Qwen2.5-omni | 19 / 19 | 64 / 64 | 12.5 |
| Kimi-Audio | 79 / 74 | 76 / 76 | 24.5 |
| OpenS2S | 97 / 93 | 66 / 65 | 50 |
| Step-Audio2 | 37 / 37 | 15 / 15 | 51 |
| Open-source Avg. | 55.3 / 53.7 | 55.1 / 54.7 | 45.0 |
| Closed-source Models | |||
| Gemini-1.5 pro | 96 / 96 | 95 / 95 | 33.5 |
| GPT-4o Audio | 98 / 98 | 100 / 100 | 83.5 |
| GPT-4o mini Audio | 100 / 100 | 100 / 100 | 86 |
| Gemini-2.5 Flash | 97 / 97 | 93 / 93 | 11 |
| Gemini-2.5-Pro | 95 / 95 | 97 / 97 | 10.5 |
| Closed-source Avg. | 97.2 / 97.2 | 97.0 / 97.0 | 44.9 |
Note: Values in IVB and HS columns are shown as GPT-4o evaluator / Qwen3 evaluator. SALMONN consistently disregarded prompt instructions by outputting audio descriptions, which prevented obtaining valid results for voice cloning spoofing. For authentication metrics, higher values indicate better security (more successful defenses against attacks). The VCS component was not assessed by large language model evaluators.
Attack Strategies. We devise a taxonomy of authentication attacks that exploit the multi-dimensional attributes of audio signals, categorizing them into three primary classes. (1) Identity Verification Bypass (IVB), which injects strong emotional cues (e.g., feigned urgency or distress) to exploit social engineering principles and induce the model to lower its security thresholds, and (2) Hybrid Spoofing (HS), this attack convolves cloned or synthesized speech with background noise and reverberation characteristic of specific scenes (e.g., a bustling office). Such attacks challenge the model’s robustness in source separation and noise suppression, and more critically, its ability to jointly reason about the logical consistency between the purported speaker identity and their acoustic context. (3) Voice Cloning Spoofing (VCS), which employs advanced speech synthesis (e.g., F5-TTS (Chen et al., 2025)) to generate highly realistic cloned voices saying a correct passphrase, thereby testing the model’s sensitivity to subtle physical traits in the vocal fingerprint. This taxonomy moves beyond traditional spoofing distinctions to directly probe ALLMs’ intrinsic vulnerabilities in processing complex auditory signals. Based on this framework, we constructed a 400-sample evaluation dataset (details in Sections˜I.1 and I.2).
Evaluation and Metrics. Our evaluation framework is designed to comprehensively assess the robustness of models under different attack scenarios. Imposter Rejection Rate (IRR) is defined as the proportion of impostor voices that are correctly rejected by the system, which will serve as the privacy metric for all attack scenarios. GPT-4o and Qwen3 will be our main evaluator, used to verify if ALLMs deny attacker requests. Human evaluators will review GPT-4o’s outputs before computing the final IRR. In the HS scenario, where the model must handle both identity spoofing and environmental interference, the IRR serves as the primary performance indicator. A success is recorded only when the model correctly rejects a forged input,such as refusing to acknowledge an identity confirmation request or denying access. Further details on the evaluation are provided in Section˜I.3 and I.4.
Results. In Table˜6, We evaluate both open- and closed-source ALLMs on our benchmark. We report IRR for the authentication task. Several key observations emerge: (1) Our findings in the IVB and HS settings reveal a critical attack vector absent in text-only systems: non-semantic features. We observed that the average IRR for open-source models was 55.3% in IVB settings, but slightly higher at 55.1% in HS settings. These settings differ considerably in paralinguistic features. The HS setting, for example, includes background audio and other speakers’ voices, both common elements in social engineering attacks conveyed through audio. This suggests that paralinguistic cues have limited influence on successful authentication by attackers. (2) In the VCS tests, we observe substantial performance discrepancies across models, with the Gemini family exhibiting comparatively weaker defenses. Interestingly, we find that simply increasing the strictness of system prompts consistently improves resilience against spoofing attacks across all systems. This suggests that in downstream ALLM applications, carefully crafted system prompts provide an efficient means of strengthening authentication security. Further details can be found in Section˜I.5.
8 Conclusions
This paper introduces AudioTrust, the first comprehensive benchmark framework for reliability assessment specifically designed for ALLMs. Unlike prior evaluations targeting text-based LLMs, AudioTrust places particular emphasis on the unique characteristics of the audio modality and the novel security challenges it entails. The framework systematically spans six key dimensions: fairness, hallucination, safety, privacy, robustness, and authentication, and also includes audio-specific risks into the design space and threat modeling. To ensure broad coverage, AudioTrust constructs a large-scale audio dataset that reflects a wide range of complex conditions. Also, we develop dedicated metrics to assess these risks, integrated with an automated pipeline powered by GPT-4o, enabling scalable evaluation. Our experimental results demonstrate that both open-source and closed-source ALLMs exhibit pronounced limitations when faced with high-risk challenges unique to the audio domain. Beyond these empirical findings, AudioTrust offers actionable insights for researchers. It defines the reliability boundaries of current ALLMs in real-world audio scenarios and lays a foundation for future work on trustworthy model design. We have publicly released our framework and evaluation platform to foster broader community-driven research in this critical area.
Acknowledgments
This research was supported by the National Research Foundation, Singapore, and the Cyber Security Agency of Singapore under the National Cybersecurity R&D Programme and the CyberSG R&D Programme Office (Award CRPO-GC3-NTU-001), NTU-NAP startup grant (024584-00001), NTU startup funding (025559-00001), and the Singapore Ministry of Education Tier 1 Grant (RG19/25). The authors from Tsinghua University are partially supported by the National Natural Science Foundation of China (No. U2341228).
Any opinions, findings, conclusions, or recommendations expressed in these materials are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore, the Cyber Security Agency of Singapore, or the CyberSG R&D Programme Office.
Reproducibility Statement
To ensure the reproducibility of our work, we release all related code, dataset metadata, and detailed experimental configurations. Our evaluation framework AudioTrust, the automated evaluation scripts, and the leaderboard generation code are hosted in an GitHub repository: https://github.com/JusperLee/AudioTrust. Appendix C describes the decoupled two-stage architecture (inference and evaluation) of our platform, which provides guidance for reproducing our evaluation pipeline. Throughout the main paper (Sections 2–7), we explicitly indicate the corresponding appendix sections. Specifically, Appendices D through I present detailed treatments of six evaluation dimensions: fairness, hallucination, safety, privacy, robustness, and verification, which cover attack strategies, dataset construction, evaluation protocols, and the metrics employed. For instance, Appendices D.1 and D.2 describe the dataset construction process for fairness evaluation, while Appendices E.3.1 and E.3.2 explain the details of the metrics for hallucination evaluation. We believe that the public release of these resources will facilitate both reproducibility and extension of our study by the community.
Ethical Statement
Our work adheres to the ICLR Code of Ethics. The primary objective of this research is to foster the development of more trustworthy, secure, and equitable Audio Large Language Models (ALLMs) by providing a comprehensive evaluation framework, AudioTrust. We aim to identify and understand the potential risks within these models, thereby guiding the community toward building safer and more reliable AI systems. Ethical considerations have been carefully addressed at every stage of our research, from dataset creation to the potential impact of our findings.
Dataset Curation and Human Participants
All data used in the AudioTrust benchmark was curated with strict ethical considerations. The majority of samples are synthetically generated (e.g., via text-to-speech systems and large language models) or drawn from publicly available datasets with permissive research licenses (e.g., Common Voice and freesound). No real personal or private information was included; sensitive content such as account numbers was generated synthetically to evaluate information-leakage risks. Annotations regarding attributes such as gender, accent, and age were taken from public datasets or inferred only when necessary, and were used solely to study potential biases. Human annotators were involved exclusively for quality control tasks (e.g., checking transcription accuracy), without collection of personal data or interventions, and thus this study does not constitute research involving human subjects.
Potential Misuse and Dual Use Concerns
Our research, by its nature, reveals vulnerabilities in current ALLMs concerning safety, privacy, robustness, and authentication. We recognize that while the methods and scenarios within AudioTrust are designed for evaluation, they could theoretically be adapted by malicious actors to develop more effective attacks. We systematically explore strategies for jailbreaking, audio based social engineering, and spoofing authentication systems. Disclosing these vulnerabilities is a double edged sword. We believe that the benefit of transparently presenting these risks to the research community through a benchmark far outweighs the risk of misuse. By establishing a public benchmark, we enable developers, both open source and proprietary, to test, fortify, and improve their models against these specific threats. Our objective is to catalyze defensive research and promote the adoption of robust safety alignment. The purpose of our public release is to accelerate this positive feedback loop, leading to safer ALLMs for everyone.
Appendix
Appendix A Introduction to Audio Large Language Models
The emergence of ALLMs signifies a pivotal paradigm shift in the domain of multimodal artificial intelligence systems (Wu et al., 2024; Peng et al., 2024). These models fundamentally extend the capabilities of traditional LLMs (Zhao et al., 2023; Hadi et al., 2023), which have demonstrated remarkable proficiency in processing and generating textual information. They achieve this by enabling the comprehension and synthesis of auditory signals. This advancement substantially surpasses conventional Automatic Speech Recognition (ASR) systems (Malik et al., 2021), whose primary objective is to faithfully transcribe spoken language into text. In contrast, ALLMs aim to achieve a more holistic understanding of acoustic environments, encompassing not only the lexical content of speech but also paralinguistic cues (e.g., prosody, affective tone), speaker characteristics, musical elements, and background environmental sounds (Tang et al., 2024). Such deep exploration of the rich semantic information embedded in audio signals is crucial for realizing more natural and context-aware human-computer interaction. ALLMs are generally divided into two primary categories: speech understanding models and speech interaction models.
The rapid maturation of this field has been largely propelled by significant advancements in self-supervised learning (SSL) methodologies, which enable models to acquire robust representations from vast quantities of unlabeled audio data. Concurrently, sophisticated multimodal training paradigms have played a critical role, facilitating the synergistic integration and joint learning of information across auditory and linguistic modalities (Girdhar et al., 2023; Rubenstein et al., 2023; Zeng et al., 2024; Wang et al., 2024a). By aligning the acoustic feature space with the inherent semantic comprehension capabilities of LLMs, ALLMs are able to address tasks beyond simple speech-to-text conversion, such as audio event classification, audio scene description, audio-based question answering, and even engaging in multi-turn spoken dialogues. These capabilities mark new frontiers for developing artificial intelligence applications that can more profoundly interpret and respond to our auditory world. However, as ALLMs are increasingly integrated into real-world applications, understanding their impact under various trustworthiness conditions becomes critically important. This study aims to construct a benchmark, AudioTrust, to comprehensively and systematically evaluate the performance and potential risks of ALLMs across different trustworthiness dimensions, such as robustness, fairness, privacy protection, and safety. This evaluation is intended to provide scientific evidence and practical guidance for the responsible development, deployment, and regulation of ALLMs.
A.1 Speech Understanding Models
Speech understanding models process and comprehend audio inputs, transforming them into semantic representations that facilitate language understanding. However, they lack the ability to generate audio responses. These models typically operate in a unidirectional manner, receiving audio as input and producing text-based outputs. Notable representatives include Qwen2-Audio (Chu et al., 2024), which integrates audio understanding capabilities into the Qwen2 (Yang et al., 2024a) via dedicated audio encoders and cross-modal adapters. These models demonstrate strong performance in tasks such as speech transcription, audio description, and audio-based question answering, yet their outputs remain restricted to textual modalities. SALMONN (Tang et al., 2024) likewise exhibits robust semantic audio understanding across diverse acoustic conditions, while maintaining a purely text-based output interface.
A.2 Speech Interaction Models
Speech interaction models go beyond mere comprehension to enable bidirectional audio communication. These models are capable not only of understanding audio inputs, but also of generating contextually appropriate audio responses, thereby facilitating more natural human-computer interaction. Prominent examples include GPT-4o (OpenAI et al., 2024), which represents a significant advance in multimodal interactive capability by processing and generating audio in near real-time conversational scenarios. MiniCPM-o 2.6 (Yao et al., 2024) provides similar functionalities in an open-source format, supporting coherent audio dialogues while demonstrating comprehension of audio contexts. Such models enable a wide range of applications, from virtual assistants to assistive tools for visually impaired users.
Appendix B Benchmark Models
To systematically investigate these trustworthiness aspects, we have selected a diverse set of models. This set includes both mainstream proprietary commercial models, such as GPT-4 (OpenAI et al., 2024) and Gemini (Team et al., 2023), as well as representative and robust open-source ALLMs, including Qwen2-Audio (Chu et al., 2024) and MiniCPM-o 2.6 (Yao et al., 2024). To ensure fairness and objectivity, all models are systematically tested on the same datasets and with identical evaluation metrics, followed by thorough comparative analyses of experimental results. It is worth noting that our methodology considers not only the fundamental audio comprehension capabilities of each model, but also examines their potential strengths and limitations in aspects such as complex interactions and knowledge transfer. This systematic safety evaluation provides a solid foundation for the future optimization and development of ALLMs. To vividly visualize these comparative strengths and weaknesses, we present a six-dimensional trustworthiness profile for each model (see Figure 3), covering fairness (F), hallucination (H), safety (S), privacy (P), robustness (R), and authentication (A). In the following descriptions, we analyze the specific radar chart performance of each model to highlight their distinct trade-offs and alignment characteristics.
B.1 Open-Source Models
In conducting trustworthiness evaluations of unified ALLMs, we selected nine representative open-source audio and multimodal models: SALMONN, Ultravox, Qwen2-Audio, MiniCPM-o 2.6, Step-Fun, Qwen2.5-omni, Kimi-Audio, OpenS2S, Step-Audio2.
1. SALMONN (Tang et al., 2024) pioneered a dual-encoder architecture (Whisper speech encoder and BEATs audio encoder) together with a window-level Q-Former and LoRA adapters. This enables the pretrained Vicuna text LLM to achieve unified understanding of speech, environmental sounds, and music. The model also demonstrates emergent capabilities in cross-modal reasoning beyond the training tasks and in few-shot activation tuning. The radar chart reveals a notable contradiction in the design of SALMONN: although it achieves surprisingly high scores in privacy protection as an open-source model, this comes at a significant cost to fairness and reliability. We hypothesize that SALMONN’s relatively weak capability in environmental sound perception limits its performance in privacy inference tasks, which in turn leads to artificially elevated privacy scores.
2. Ultravox (AI, 2024) directly maps raw audio into the high-dimensional representation space of LLMs, thereby seamlessly eliminating the traditional ASR stage. This model not only comprehends speech content but also captures paralinguistic features such as tone and pauses, and supports streaming text outputs. Ultravox demonstrates excellent safety and decent hallucination resistance, but its robustness is critically low. This indicates that while the model is well-aligned semantically, its end-to-end architecture is extremely fragile to acoustic perturbations.
3. Qwen2-Audio (Chu et al., 2024) is a large-scale audio-language model that establishes a seamless pipeline between the Whisper-large-v3 encoder and the Qwen-7B language model, thereby supporting both spoken dialogue and audio analysis interaction modes. In real conversational and multitask zero-shot evaluations, the model leverages Mel-spectrograms of 16kHz audio combined with instruction tuning and Direct Preference Optimization (DPO), significantly improving the precision and robustness of responses to human intent. The cross-dimensional analysis highlights a significant discrepancy in Qwen2-Audio: although it performs strongly in terms of safety and authentication, its privacy score remains alarmingly low. This indicates that the model can effectively identify speakers but fails to safeguard their sensitive attributes. A potential reason may lie in the lack of effective safety-alignment training after integrating the new modality, or an excessive reliance on ASR outputs, which leads to insufficient protection of sensitive audio attributes.
4. MiniCPM-o 2.6 (Yao et al., 2024) integrates four major components: SigLip-400M, Whisper-medium, ChatTTS-200M, and Qwen2.5-7B, supporting bilingual real-time dialogue in an end-to-end multimodal fashion, as well as controllable interactions in emotion and speaking rate, and high-quality voice cloning. It consistently outperforms proprietary models of equivalent scale on benchmarks such as OpenCompass and StreamingBench. The model exhibits a clear trade-off: it prioritizes robustness and safety while largely neglecting privacy and performing poorly in controlling hallucinations. This characteristic suggests that the system demonstrates strong resistance to noise and harmful prompts yet is prone to private information leakage and the generation of unsupported content.
5. Step-Fun (Huang et al., 2025) is a production-ready open-source real-time speech–text multimodal system that tackles data-collection cost, weak dynamic control, and limited intelligence via four pillars: a 130B unified understanding–generation model, a generative speech data engine enabling affordable voice cloning and distilling the lightweight Step-Audio-TTS-3B, an instruction-driven fine-control mechanism spanning dialects, emotions, singing, and rap, and an enhanced cognitive layer with tool calling and role-playing for complex tasks. Step-Fun stands out as the most balanced open-source model, particularly in terms of fairness and robustness. Unlike comparable models that often sacrifice fairness for safety, Step-Fun maintains moderate performance across all dimensions, although its authentication results indicate remaining room for improvement in identity verification.
6. Qwen2.5-Omni (Xu et al., 2025b) builds upon Qwen2.5-VL/Audio by introducing the Thinker-Talker architecture and TMRoPE (Time-aligned Multimodal RoPE) temporal alignment embedding. This allows the model to stream and process text, image, audio, and video inputs concurrently within a single framework, with the ability to produce both textual and natural speech outputs in synchronization. Qwen2.5-Omni enhances safety but exacerbates cross-dimensional imbalance: fairness decreases while privacy becomes nearly negligible. This reflects a broader trend wherein cross-modal alignment training may inadvertently suppress the fine-grained reasoning capabilities necessary for achieving both fairness and privacy protection.
7. Kimi-Audio (Ding et al., 2025) is an open-source audio foundation model for understanding, generation, and conversation; it adopts a 12.5Hz audio tokenizer and an LLM-based architecture that ingests continuous features and emits discrete tokens, alongside a chunk-wise streaming detokenizer via flow matching for low-latency inference. Kimi-Audio exhibits the most extreme alignment trade-off in our benchmark. While it achieves near-perfect safety, such over-cautious behavior likely results in the lowest fairness score and very limited privacy protection. This may occur because the model indiscriminately rejects legitimate queries from marginalized groups or fails to distinguish between safe and private contexts.
8. OpenS2S (Wang et al., 2025a) built on the BLSP-Emo empathetic speech-to-text backbone, it introduces a streaming interleaved decoding architecture for low-latency speech generation while capturing rich paralinguistic cues for expressive responses. Interestingly, OpenS2S exhibits characteristics opposite to most models: it performs strongly in hallucination resistance and authentication but shows weaker robustness and fairness. This suggests that the model achieves high precision under clean conditions yet lacks the generalization ability to handle diverse speakers or noisy environments.
9. Step-Audio 2 (Wu et al., 2025) is an end-to-end multimodal LLM for industry-grade audio understanding and speech conversation, combining a latent audio encoder with reasoning-centric RL to boost ASR and audio comprehension; by folding discrete audio-token generation into language modeling, it becomes highly responsive to paralinguistic cues in real time. The model performs reasonably well in robustness and safety but exhibits systematic weaknesses in fairness and privacy. This indicates that its reasoning-centered reinforcement learning optimization primarily focuses on utility and ASR performance rather than ethical alignment.
B.2 Closed-Source Models
Among closed-source ALLMs, Google’s Gemini series (Team et al., 2023) and OpenAI’s GPT-4o series (OpenAI et al., 2024) represent the industry’s state-of-the-art in audio understanding and interaction technologies. In our evaluation of various safety concerns, we employ both the Gemini and GPT-4o model series.
10. Gemini-1.5 Pro leverages a Mixture-of-Experts architecture for unified reasoning across speech, image, and text. It supports audio inputs up to 19 hours in duration and contexts up to the million-token scale, enabling seamless processing for tasks such as audio summarization, transcription, and translation. Although Gemini-1.5 Pro demonstrates the best safety performance and excels in robustness and authentication, its relatively low hallucination score reveals a critical vulnerability. The model exhibits strong resistance to adversarial attacks but tends to fabricate content, highlighting a notable gap between safety alignment and factual grounding.
11. GPT-4o Audio is the first developer-oriented interactive audio model that supports both understanding and generation of speech. It is capable of speech transcription, summarization, sentiment analysis, and conversational dialogue. GPT-4o Audio demonstrates the most comprehensive performance, leading in authentication and achieving high scores in safety, privacy, and fairness. However, similar to Gemini, it remains affected by hallucinations, suggesting that scaling up the model can enhance nearly all dimensions of trustworthiness, while hallucination issues may stem from inaccuracies in the training data.
12. GPT-4o mini Audio is designed to deliver cost-effective yet robust audio understanding and generation. It supports a variety of audio input formats and can produce seamless bimodal (text and speech) output with customizable speech styles, making it applicable to edge devices and large-scale embedded deployments. The "mini" variant illustrates a clear compression trade-off: it retains the superior authentication and safety of the larger model (even improves privacy), but its robustness collapses to 0.228. This suggests that lightweight models can be safe and private but lack the parameter redundancy to handle adversarial noise.
13. Gemini-2.5 Flash retains the core multimodal design of the Pro version while significantly optimizing inference speed and computational efficiency. This version supports up to 8.4 hours of audio input and million-token context windows, with dramatically reduced latency and operational cost compared to the Pro variant, while still covering tasks like audio summarization, transcription, and translation. Gemini-2.5 Flash prioritizes safety and robustness to serve as a reliable production model. However, cross-dimensional analysis shows it lags in hallucination and fairness, suggesting that optimization for speed and safety may have compromised its ability to handle nuanced equity tasks and factual verification.
14. Gemini-2.5 Pro further advances multimodal reasoning, introducing a dynamic “thinking budget” mechanism that adaptively allocates computational resources based on instruction and system constraints. Its superior performance on video understanding benchmarks extends to the audio domain, enabling streaming responses for complex tasks such as conversational QA, scenario retrieval, and reasoning through efficient temporal alignment and cross-modal integration. Compared to its predecessor, Gemini-2.5 Pro has improved hallucination resistance while maintaining elite safety and robustness. Nevertheless, fairness remains a persistent challenge, reinforcing the observation that advanced reasoning capabilities do not automatically translate to equitable decision-making without explicit intervention.
Appendix C Platform Design of AudioTrust
To systematically address trustworthiness risks stemming from the rapid development of ALLMs and to establish a reproducible, extensible, and forward-looking evaluation system, we introduce a unified trustworthiness assessment framework. Our framework’s core design philosophy relies on highly modular abstraction mechanisms and a two-stage decoupled architecture. This design aims to facilitate continuous and rigorous trustworthiness risk assessment and in-depth analysis of ALLMs. The proposed architecture emphasizes flexibility and efficiency, decomposing complex evaluation procedures into two distinct yet interconnected stages: the inference execution stage (Stage 1) and the trustworthiness evaluation stage (Stage 2). As illustrated in Figure 4, such a decoupled design paradigm brings notable practical advantages. It grants researchers and evaluators considerable autonomy to independently execute the inference or evaluation workflows according to specific research objectives or evaluation requirements. For instance, when model outputs are already available, this pre-generated response data can be directly used for comprehensive trustworthiness analyses and comparisons across multiple dimensions and methods. This approach significantly enhances evaluation flexibility while optimizing the use of computational resources and reducing time costs.
The inference execution stage focuses on raw data processing and the collection of model outputs. First, the data engine module efficiently loads and preprocesses various standard trustworthiness benchmark datasets, including both publicly released open benchmarks and custom-built datasets, thus ensuring data consistency and traceability. Subsequently, users can flexibly specify evaluation models, datasets, evaluation targets, and runtime parameters through configuration files. This enables batch parallel scheduling and significantly optimizes computational resource usage. The core inference module supports mainstream ALLMs inference tasks, allowing direct loading of open-source models from the Hugging Face Hub, and natively integrates adapters for closed-source models accessed via APIs, thereby providing comprehensive full-stack support for major ALLMs. Through the aforementioned workflow, structured raw model output files are generated for subsequent analysis, ensuring a highly reproducible evaluation process.
The trustworthiness evaluation stage performs independent, multidimensional, automated analysis on the model outputs generated in Stage 1. Owing to the architectural decoupling, this stage can independently process historical inference results in bulk, significantly enhancing evaluation flexibility. We introduce multiple robust pretrained automated evaluators (evaluator models), covering critical trustworthiness dimensions such as content safety review, bias detection, and factual consistency. These evaluators, independently or jointly, conduct in-depth assessments and quantitative scoring of model outputs based on preset standards and metrics. This process enables automatic annotation and efficient pre-screening. Automated evaluation not only greatly improves assessment efficiency, but also reduces the subjective bias associated with human evaluation.
Platform Validation.
Beyond automated evaluators, we include a human-aided validation protocol: 10% of the data are randomly sampled and cross-checked by 20 trained annotators, with each QA pair independently reviewed by three annotators. Final labels are decided by majority voting, yielding a 97–98% agreement with the platform’s automated assessments, thereby validating the reliability of the framework.
Appendix D Additional Details of Evaluation on AudioTrust Fairness
D.1 Dataset Classification Criteria
We utilized seven sensitive attributes to ensure both group and individual fairness: gender, race, age, accent, economic status, personality traits, and speech fluency. All sensitive attributes were defined with binary values. Specifically, the sensitive attribute sets were as follows: gender , age , race , accent , linguistic characteristic , economic status , and personality traits .
D.2 Dataset Construction Method
We select the types of fairness to be evaluated for ALLMs following two principles.
Multifaceted social biases: We consider common societal biases associated with multiple sensitive attributes such as gender, race, and age (Cho et al., 2023; Zhang et al., 2017b; Wan et al., 2024; Hirota et al., 2023). In addition to these attributes, we include distinctive stereotypes uniquely identifiable through audio characteristics, including accent, personality traits, economic status, and linguistic characteristic.
Real-world applications: We focus on realistic decision-making applications in which AI fairness is crucial, including recruitment processes, admission systems, and financial lending evaluations. Furthermore, we incorporate commonly encountered stereotypical scenarios drawn from real-life contexts such as occupation, education, and healthcare.
Data Construction.
During the data construction phase, we utilized the multi-modal generation model GPT-4o (OpenAI et al., 2024) to produce textual content, which was then transformed into audio using the F5-TTS model (Chen et al., 2025). The entire data generation process begins with a text-based prompt, constructed entirely by GPT-4o to support natural language fluency and semantic consistency. To enable controlled experiments with single-variable variation, each data pair consists of two audio samples with identical semantic content, differing solely in their associated sensitive attributes.
For example, to simulate racial identity, we altered the input language to represent different ethnic groups. GPT-4o was employed to ensure that the translated texts conveyed semantically equivalent meanings across languages, thereby eliminating confounding effects introduced by lexical or syntactic discrepancies. For sensitive attributes that cannot be directly simulated by changing the voice source, such as linguistic style or economic status, we implemented more sophisticated data augmentation strategies.
To simulate economic status, we enriched the base audio with representative background sounds to enhance the perceptual cues of social identity. Examples include upscale restaurant ambience for high-income individuals and street market noise for lower-income individuals. individuals. Regarding linguistic style, we adjusted the fluency and structure of the input text before feeding it into the TTS model. This allowed us to generate speech samples that reflect varying levels of coherence, prosody, and articulation, simulating differences in communication style associated with social or educational background.
Real-World Applicability.
To assess the real-world applicability of ALLMs, we constructed six prototypical evaluation scenarios, including three decision-making tasks and three stereotype judgment tasks. In the decision-making settings, the ALLM is assigned a contextualized role and is required to make a selection based on audio input. For instance, in the Hiring scenario, the ALLM acts as a corporate recruiter selecting the final candidate for a customer service representative position, based on self-introduction audio recordings from two applicants. In the Admission scenario, the ALLM plays the role of a selection committee member choosing between two candidates to join a technology innovation program. In the Financial Loan scenario, the ALLM simulates a loan officer who must approve a loan application for one of the two speakers based on their respective verbal statements.
For stereotype-based assessments, we designed three tasks corresponding to occupational, educational, and medical contexts. In the Stereotype scenario, the ALLM must infer whether the speaker sounds more like a boss or a subordinate, enabling analysis of implicit bias in role perception. The Math Ability Stereotype scenario evaluates whether the model’s judgment on mathematical proficiency is influenced by speaker attributes. Lastly, the Medical Stereotype task examines whether the model associates certain vocal characteristics with being a doctor versus a nurse or medical assistant.
Together, this pipeline offers a comprehensive and reproducible framework for evaluating the fairness and decision-making integrity of ALLMs across realistic and socially sensitive scenarios. It enables systematic exploration of model behavior under controlled attribute manipulations, paving the way for more equitable and trustworthy audio-based AI systems.
D.3 Evaluation Method and Metrics
We utilized statistical methods and developed the fairness metrics to quantitatively assess model fairness (Xu et al., 2025a; Deldjoo & Nazary, 2024). Let an audio large language model be , mapping the audio–text input space to the text output space . Denote by the set of sensitive attribute values (e.g. ). For a test collection , we introduce a discriminator detecting sensitive attributes in generations.
Group Fairness Score. For any group label , define the group fairness score
where each probability is approximated via ‑sample Monte‑Carlo estimates. A positive implies a fairness towards group . When aggregating across tasks, we report the absolute value to emphasise fairness magnitude only.
D.4 Experimental Design and Results
Experimental Design.
We designed three experiments for evaluating fairness: group decision-making, and group stereotypes. The detailed experimental design is illustrated in the Figure 6. In the group decision-making experiment, we emphasized realistic high-stakes decision-making contexts, namely recruitment processes, admission systems, and financial lending evaluations. This approach enabled us to assess group fairness across a wide range of socially significant contexts and demographic dimensions.
For evaluating stereotypes, we designed a classification task for ALLMs, where models classify audio samples based on sensitive attributes in occupational, educational, and medical scenarios. For instance, models determine whether an audio sample belongs to a doctor or a nurse based on gender-specific inputs.
To illustrate the fairness differences between ALLMs and LLMs, we replicated the experiments for LLMs using purely textual data (derived from the content used for audio generation). While the prompt and input data formats were consistent, the manipulation of sensitive attributes differed. For audio, attributes were modified by selecting different voice sources or mixing audio tracks. For text, these attributes were directly embedded within the textual content. Further details are provided in our examples. For more details, please see our examples.
| SALMONN | Ultravox | Qwen2-Audio | MiniCPM-o 2.6 | Step-Fun | Qwen2.5-omni | Kimi-Audio | OpenS2S | Step-Audio2 | ||
| Recruitment | ||||||||||
| Female | ||||||||||
| Old | ||||||||||
| American | ||||||||||
| Clam | ||||||||||
| Fluent | ||||||||||
| Chinese | ||||||||||
| Wealthy | ||||||||||
| Admission | ||||||||||
| Female | ||||||||||
| Old | ||||||||||
| American | ||||||||||
| Clam | ||||||||||
| Fluent | ||||||||||
| Chinese | ||||||||||
| Wealthy | ||||||||||
| Financial Loan | ||||||||||
| Female | ||||||||||
| Old | ||||||||||
| American | ||||||||||
| Clam | ||||||||||
| Fluent | ||||||||||
| Chinese | ||||||||||
| Wealthy | ||||||||||
| Average | 0.088 | 0.368 | 0.290 | 0.415 | 0.505 | 0.202 | 0.086 | 0.157 | 0.098 | |
| Gemini-1.5 Pro | GPT-4o Audio | GPT-4o mini Audio | Gemini-2.5 Flash | Gemini-2.5 Pro | ||
| Recruitment | ||||||
| Female | ||||||
| Old | ||||||
| American | ||||||
| Clam | ||||||
| Fluent | ||||||
| Chinese | ||||||
| Wealthy | ||||||
| Admission | ||||||
| Female | ||||||
| Old | ||||||
| American | ||||||
| Clam | ||||||
| Fluent | ||||||
| Chinese | ||||||
| Wealthy | ||||||
| Financial Loan | ||||||
| Female | ||||||
| Old | ||||||
| American | ||||||
| Clam | ||||||
| Fluent | ||||||
| Chinese | ||||||
| Wealthy | ||||||
| Average | 0.460 | 0.274 | 0.245 | 0.240 | 0.205 | |
| SALMONN | Ultravox | Qwen2-Audio | MiniCPM-o 2.6 | Step-Fun | Qwen2.5-omni | Kimi-Audio | OpenS2S | Step-Audio2 | ||
| Occupational | ||||||||||
| Female | ||||||||||
| Old | ||||||||||
| American | ||||||||||
| Clam | ||||||||||
| Fluent | ||||||||||
| Chinese | ||||||||||
| Wealthy | ||||||||||
| Education | ||||||||||
| Female | ||||||||||
| Old | ||||||||||
| American | ||||||||||
| Clam | ||||||||||
| Fluent | ||||||||||
| Chinese | ||||||||||
| Wealthy | ||||||||||
| Medical | ||||||||||
| Female | ||||||||||
| Old | ||||||||||
| American | ||||||||||
| Clam | ||||||||||
| Fluent | ||||||||||
| Chinese | ||||||||||
| Wealthy | ||||||||||
| Average | 0.139 | 0.238 | 0.333 | 0.248 | 0.657 | 0.067 | 0.036 | 0.017 | 0.074 | |
| Gemini-1.5 Pro | GPT-4o Audio | GPT-4o mini Audio | Gemini-2.5 Flash | Gemini-2.5 Pro | ||
| Occupational | ||||||
| Female | ||||||
| Old | ||||||
| American | ||||||
| Clam | ||||||
| Fluent | ||||||
| Chinese | ||||||
| Wealthy | ||||||
| Education | ||||||
| Female | ||||||
| Old | ||||||
| American | ||||||
| Clam | ||||||
| Fluent | ||||||
| Chinese | ||||||
| Wealthy | ||||||
| Medical | ||||||
| Female | ||||||
| Old | ||||||
| American | ||||||
| Clam | ||||||
| Fluent | ||||||
| Chinese | ||||||
| Wealthy | ||||||
| Average | 0.297 | 0.930 | 0.860 | 0.370 | 0.320 | |
Results.
We evaluate the social decision-making fairness and stereotype fairness of nine models in Appendix D.4, Table 7, Table 8, Table 9, Table 10, Table D.4. The results reveal the following key findings:
| Model | Female | Old | Chinese | Wealthy |
| Audio Large Language Models | ||||
| Gemini-1.5 Pro | 0.35 | 0.90 | 0.25 | 0.90 |
| Gemini-2.5 Flash | 0.20 | 0.50 | 0.11 | 0.20 |
| Gemini-2.5 Pro | 0.00 | 0.50 | 0.00 | 0.00 |
| GPT-4o Audio | 0.20 | 0.30 | 0.25 | 0.20 |
| GPT-4o Mini Audio | 0.30 | 0.10 | 0.20 | 0.50 |
| Large Language Models | ||||
| Gemini-1.5 Pro | 0.00 | 0.00 | 1.00 | 0.00 |
| Gemini-2.5 Flash | 0.00 | 0.20 | 0.20 | 0.00 |
| Gemini-2.5 Pro | 0.00 | 0.00 | 0.00 | 0.00 |
| GPT-4o | 0.00 | 0.00 | 0.20 | 0.00 |
| GPT-4o Mini | 0.00 | 0.00 | 0.00 | 0.20 |
| Average | 0.11 | 0.25 | 0.17 | 0.18 |
(1) The fairness levels vary significantly among different ALLMs. Notably, models generally considered highly capable, such as GPT-4o Audio, GPT-4o mini Audio, Gemini-2.5 Flash, and Gemini-2.5 Pro, exhibit the highest group unfairness in the decision-making experiments. In contrast, some lower-performing open-source models, such as MiniCPM-o 2.6, Qwen2-Audio, SALMONN, and Ultravox, and Step-Fun, demonstrate relatively better fairness. However, these models still exhibit high group unfairness and are far from ideal models. (2) Overall, the model’s responses tend to favor sensitive attributes such as female, old, American accent, calm, fluent, Western, and wealthy. (3) In the stereotype experiments, GPT-4o Audio and GPT-4o mini Audio show excellent fairness, while MiniCPM-o 2.6, Qwen2-Audio, SALMONN, and Ultravox exhibit the highest unfairness. Interestingly, GPT-4o Audio and GPT-4o mini Audio perform well in stereotype experiments by almost refusing to answer all harmful questions (the proportion of responses across attributes is nearly consistent), but do not refuse in decision-making tasks. This indicates that the GPT series models face challenges in accurately determining whether a question is genuinely harmful. (4) Most models that exhibit high group (un)fairness when evaluating social stereotypes, such as MiniCPM-o 2.6, Qwen2-Audio, SALMONN, Ultravox, and the Gemini series, also maintain similar levels in decision-making scenarios. (5) ALLMs exhibit basically the same degree of unfairness across different scenarios. (6) The performance of LLMs in decision-making scenarios is worse compared to ALLMs.
Appendix E Additional Details of Evaluation on AudioTrust Hallucination
E.1 Dataset Classification Criteria
To evaluate the robustness of the model in identifying and suppressing hallucination content and semantic contradiction information, we propose a comprehensive evaluation framework. The framework’s core design revolves around four key and representative potential hallucination categories in ALLMs. This approach aims to cover a wide range of complex challenge scenarios, from subtle semantic biases to significant factual errors, thereby enabling an in-depth understanding of model performance under various pressures or interference conditions. The detailed experimental design is illustrated in Figure 7.
We divide the hallucination detection dataset into two major categories: logical hallucinations and factual hallucinations. The logical hallucination class includes hallucination data with temporal logical errors or text that semantically contradicts the core attributes/content of the audio. On the other hand, the factual hallucination class consists of hallucination data that violates fundamental physical laws and conflicts with the descriptive attributes of acoustic features. To better illustrate each category, we provide representative example prompts for each type of hallucination. These examples demonstrate the nature of the direction tested in our dataset:
E.2 Dataset Construction Method
To construct the datasets for physical and logical violations, we adopted a two-stage procedure. First, we utilized GPT-4o (OpenAI et al., 2024) and audio data from Freesound111http://freesound.org/ to generate 80 sounds that represent scenarios with physical or temporal logical inconsistencies. Subsequently, we edited the collected audio content and arranged and concatenated them according to the scenarios generated by GPT-4o. The choice of GPT-4o for scene generation is attributed to its advanced linguistic capabilities and alignment guarantees, which ensure both the diversity and reliability of the generated scenarios.
In addition, to create audio corresponding to content mismatches and label mismatches, we incorporated the emotional speech dataset (Zhou et al., 2022) and obtained music classification datasets from Freesound that align with our testing objectives. To expose these vulnerabilities, we randomly associated mismatched emotion labels with the audio. To ensure controllable model outputs and the reliability of the evaluation metrics, we opted to randomly recombine audio and text classification labels without altering the classification types of the original datasets.
In the end, our dataset comprises a total of 320 audio hallucinations (along with corresponding semantic text annotations): 160 factual hallucinations targeting a variety of scenarios and 160 logical hallucinations targeting diverse logical errors. This construction approach offers a systematic methodology for generating challenging test cases to evaluate the safety mechanisms of GPT-4o, while simultaneously encompassing a wide range of hallucination forms and contexts.
E.3 Experimental Design and Evaluation Metrics
E.3.1 Experimental Design
We designed a comprehensive red-teaming framework to evaluate hallucination tendencies in ALLMs across four distinct categories: physical violations, label mismatches, logical violation, and content mismatches. Our methodology involved creating specialized datasets that test specific aspects of model robustness against hallucination. For physical violations, we synthesized sounds that contradict fundamental laws (e.g., underwater flames); for label mismatches, we deliberately paired audio with incorrect attribute descriptions (e.g., mislabeling speaker gender); for temporal disorders, we manipulated cause-and-effect sequences in audio events (e.g., reversing “gunshot followed by falling”); and for content mismatches, we paired audio with semantically contradictory descriptions (e.g., labeling bird chirping as dog barking). At the same time, we found that if the model’s response relies on overly forced explanations (for example, "The object landed precisely on a perfect sound-absorbing material"), to avoid ambiguity and excessive rationalization, this evaluation does not rely on the open-ended responses of the model. We use GPT-4o as the automatic evaluator, which follows pre-defined common sense rules. This evaluator is explicitly instructed to ignore special environments or recording settings and only focus on whether the described event follows the laws of everyday physics. Models that unquestioningly accept extreme or temporary set scenarios will be regarded as having produced hallucinations and will receive a lower score. We conducted four experimental protocols with varying prompt structures:
Experiment I provided audio with annotations and assessed physical law conformity, description accuracy, and required models to self-evaluate their responses;
Experiment II focused on physical law conformity and content accuracy without logical violation testing;
Experiment III prompted models to describe audio content and assess the reality relationship of their descriptions;
The following is an example of a simplified experimental process:
We measured performance using metrics including hallucination rate, detection accuracy, refusal rate, complete disclosure rate, and cautious disclosure rate, with results aggregated by model and hallucination type to provide a comprehensive assessment of how different ALLMs handle potentially misleading audio information. Meanwhile, in order to better repeatable method is provided, a repeatable method is provided to test challenging test cases. We offer a universal json data template in the illusion detection dataset. The json template is as follows:
E.3.2 Evaluation Metrics
This section details the quantitative and qualitative metrics used to evaluate the performance of ALLMs in detecting various types of inconsistencies and violations.
Hallucination Detection Rate (HDR): This metric generally measures the proportion of actual ‘positive’ cases (e.g., physical violations, label mismatches) that are correctly identified by the model. It is calculated as:
For instance, in the context of Physical Violations, true positive cases are the actual physical violation audios. For Label Mismatches, true positive cases are the audios with incorrect labels.
False Hallucination Rate (FHR): This metric generally measures the proportion of actual ‘negative’ cases (e.g., physically possible scenarios, correctly labeled audios) that are incorrectly flagged by the model as ‘positive’. It is calculated as:
For Physical Violations, true negative cases are the control audios. For Label Mismatches, true negative cases are audios with correct labels.
Logical Violation Detection Rate (LVDR): Specifically for Logical Violations, this measures the proportion of temporally or causally illogical audio sequences correctly identified:
False Violation Rate (FVR): Specifically for Logical Violations, this measures the proportion of logical audio sequences incorrectly flagged as illogical:
Content Mismatch Detection Rate (CMDR): Specifically for Content Mismatches, this measures the proportion of semantic inconsistencies between audio content and textual descriptions correctly identified:
False Content Mismatch Rate (FCMR): Specifically for Content Mismatches, this measures the proportion of accurate audio-text descriptions incorrectly flagged as mismatches:
Attribution Accuracy (AA): Used in Label Mismatch evaluations, this metric measures how accurately the model attributes the correct label (e.g., true emotion, gender, genre) for cases where a mismatch was correctly identified:
Explanation Quality Score (EQS): A qualitative metric used for Physical Violations. It is determined by human evaluators who rate the quality of the ALLM’s explanations for identified violations on a 5-point scale, considering physical accuracy, relevance to audio content, and clarity of reasoning. The final EQS is an average across evaluators and test cases.
Causal Reasoning Score (CRS): A qualitative metric used for Logical Violations. This is a 10-point human-evaluated scale measuring the quality of the ALLM’s causal explanations, based on temporal ordering accuracy, recognition of causal relationships, and clarity.
Description Accuracy Score (DAS): Used in Content Mismatch evaluations for cases where a mismatch was correctly identified. This metric measures the accuracy of the model’s alternative (corrected) description of the actual audio content, typically using automated scores like BLEU and ROUGE against human-generated ground truth descriptions.
E.4 Evaluation Methodology
This study employs a systematic three-stage evaluation protocol to comprehensively assess the performance of models in physical violation detection tasks. In the initial stage of violation detection, the model (ALLM) is provided with both audio files and their corresponding JSON metadata. The assessment is carried out according to the following instruction:
Based on the content described in the Overall_description field of the audio and JSON files, determine whether the audio conforms to physical laws, and provide reasoning for your judgment.
This process is designed to evaluate the model’s capability to judge the physical consistency between audio content and its paired textual description. The model is required to integrate multimodal information and leverage physical common sense to identify potential violations and articulate the rationale behind its decisions.
Subsequently, in the self-evaluation stage, the model conducts introspective assessment based on its previous judgment. Specifically, the following evaluation prompt is introduced:
Considering the answer to the previous question, use the eval_value entry in the JSON file as an input prompt, and employ it as an evaluation criterion to score the previous response.
This stage emphasizes the model’s capacity for self-reflection; that is, its ability to provide objective evaluations of the reliability of its own physical reasoning, based on structured evaluation metrics and its own output.
In the metrics calculation stage, we utilize a series of quantitative metrics to evaluate model performance (see Section˜E.3.2 for complete formulations). Our evaluation framework is twofold, comprising both automated metrics and human-based judgments.First, we programmatically compute two metrics by parsing the model’s textual output and comparing it against ground-truth labels, requiring no human judgment. These are: 1) The HDR (Hallucination Detection Rate), which measures detection sensitivity by calculating the proportion of true physical violations (our positive class) that the model correctly identifies. 2) The FHR (False Hallucination Rate), which assesses the false positive rate by calculating the proportion of normal control cases (our negative class) that the model incorrectly flags as violations.Second, to complement these automated metrics, we introduce the EQS (Explanation Quality Score) as a human-based measure of interpretability. This score is assigned by three expert human raters on a 5-point scale. Ratings evaluate multiple perspectives, including physical correctness, the relevance of the explanation to the audio facts, and the logic and clarity of the reasoning process. The final EQS is computed as the mean score across all raters and test cases.
Overall, this multi-dimensional framework, combining automated detection accuracy (HDR/FHR) with human-rated explanation quality (EQS), effectively captures the model’s competence and provides a reliable experimental foundation.
E.5 Result Analysis
| Model | Content Mismatch | Label Mismatch | Logical Violation | Physical Violation |
| Open-source Models | ||||
| MiniCPM-o 2.6 | 6.51/5.98/6.23 | 6.00/6.45/6.15 | 8.53/8.01/8.30 | 6.40/5.88/6.11 |
| Qwen2-Audio | 8.33/7.90/8.22 | 4.74/4.10/4.18 | 7.01/7.55/7.22 | 7.50/8.01/7.80 |
| SALMONN | 2.40/2.95/2.60 | 1.50/0.99/1.17 | 6.94/6.35/6.63 | 4.21/3.70/3.99 |
| Ultravox | 5.98/5.50/5.74 | 4.22/4.70/4.64 | 7.76/8.25/7.99 | 8.04/8.60/8.38 |
| Step-Fun | 3.97/3.83/4.09 | 6.17/6.33/5.78 | 5.88/5.75/6.30 | 8.89/8.50/8.96 |
| OpenS2S | 2.01/1.79/1.89 | 2.75/2.75/4.75 | 8.00/7.00/5.97 | 8.89/8.89/8.31 |
| Kimi-Audio | 1.38/1.39/1.38 | 2.42/3.15/2.75 | 5.00/6.00/6.09 | 4.50/4.00/8.58 |
| Qwen2.5-Omni | 7.96/8.02/7.99 | 3.80/3.00/5.57 | 7.67/7.67/8.12 | 5.20/5.00/6.36 |
| Step-Audio2 | 3.51/3.73/3.61 | 0.00/0.00/5.82 | 0.00/0.00/7.80 | 0.00/0.00/8.28 |
| Closed-source Models | ||||
| Gemini-1.5 Pro | 8.10/8.66/8.48 | 7.56/8.05/7.82 | 8.90/8.42/8.65 | 8.62/9.10/8.88 |
| Gemini-2.5 Flash | 7.73/8.21/8.00 | 8.06/8.66/8.35 | 8.46/8.99/8.68 | 8.81/8.32/8.58 |
| Gemini-2.5 Pro | 8.49/7.91/8.17 | 8.99/8.53/8.82 | 8.99/8.41/8.70 | 8.20/8.77/8.50 |
| GPT-4o Audio | 4.20/3.71/3.91 | 2.98/2.43/2.63 | 3.29/3.77/3.53 | 9.01/8.55/8.81 |
| GPT-4o mini Audio | 2.00/2.61/2.41 | 1.00/1.49/1.14 | 1.51/0.98/1.23 | 8.75/9.22/9.03 |
Scores follow the format “DIM 1 / DIM 2 / DIM 3”. Higher values indicate better performance.
| Model | Content Mismatch | Label Mismatch | Logical Violation | Physical Violation |
| Open-source Models | ||||
| MiniCPM-o 2.6 | 6.24 / 9.42 | 6.20 / 9.58 | 8.28 / 8.31 | 6.13 / 8.05 |
| Qwen2-Audio | 8.15 / 9.65 | 4.34 / 9.33 | 7.26 / 8.02 | 7.77 / 8.63 |
| SALMONN | 2.65 / 8.85 | 1.22 / 8.67 | 6.64 / 7.24 | 3.98 / 6.91 |
| Ultravox | 5.74 / 9.31 | 4.52 / 9.46 | 8.01 / 8.78 | 8.34 / 8.94 |
| Closed-source Models | ||||
| Gemini-1.5 Pro | 8.41 / 9.82 | 7.81 / 9.88 | 8.66 / 9.63 | 8.87 / 9.51 |
| Gemini-2.5 Flash | 7.98 / 9.71 | 8.36 / 9.79 | 8.71 / 9.25 | 8.57 / 9.03 |
| Gemini-2.5 Pro | 8.19 / 9.79 | 8.78 / 9.91 | 8.70 / 9.69 | 8.49 / 9.42 |
| GPT-4o Audio | 3.90 / 9.22 | 2.68 / 9.15 | 3.53 / 7.03 | 8.79 / 8.88 |
| GPT-4o mini Audio | 2.34 / 9.03 | 1.21 / 8.92 | 1.24 / 7.38 | 9.00 / 9.11 |
Values shown as "ALLM / Text LLM" pairs with red arrows indicating performance gap between ALLM and hypothetical text-only LLM processing. : ALLM performance falls behind text LLM by the subscript amount. Higher values (0-10 scale) indicate better performance.
| Open-source models | ||||||||||||
| Test Type | MiniCPM-o 2.6 | Qwen2-Audio | SALMONN | Ultravox | ||||||||
| I(%) | N(%) | C(%) | I(%) | N(%) | C(%) | I(%) | N(%) | C(%) | I(%) | N(%) | C(%) | |
| Content Mismatch | 40.00 | 40.00 | 20.00 | 100.00 | 0.00 | 0.00 | 0.00 | 100.00 | 0.00 | 38.46 | 53.85 | 7.69 |
| Label Mismatch | 50.00 | 25.00 | 25.00 | 0.00 | 100.00 | 0.00 | 0.00 | 25.00 | 75.00 | 37.50 | 43.75 | 18.75 |
| Logical Violation | 18.18 | 81.82 | 0.00 | 0.00 | 100.00 | 0.00 | 0.00 | 91.67 | 8.33 | 14.81 | 74.07 | 11.11 |
| Physical Violation | 20.00 | 70.00 | 10.00 | 0.00 | 75.00 | 25.00 | 11.11 | 44.44 | 44.44 | 23.81 | 61.90 | 14.29 |
| Test Type | Step-Fun | OpenS2S | Kimi-Audio | Qwen2.5-Omni | ||||||||
| I(%) | N(%) | C(%) | I(%) | N(%) | C(%) | I(%) | N(%) | C(%) | I(%) | N(%) | C(%) | |
| Content Mismatch | 54.6 | 16.2 | 29.2 | 81.7 | 2.5 | 15.8 | 67.5 | 29.2 | 3.3 | 26.2 | 0.0 | 73.8 |
| Label Mismatch | 50.0 | 20.4 | 29.6 | 57.5 | 4.6 | 37.9 | 28.7 | 33.3 | 37.9 | 77.1 | 1.2 | 21.7 |
| Logical Violation | 34.4 | 14.4 | 51.2 | 54.4 | 3.1 | 42.5 | 33.1 | 55.0 | 11.9 | 50.6 | 31.2 | 18.1 |
| Physical Violation | 11.7 | 11.7 | 76.7 | 41.6 | 15.1 | 43.3 | 23.8 | 17.5 | 58.8 | 52.5 | 5.8 | 41.7 |
| Closed-source models | |||||||||||||||
| Test Type | Gemini-1.5 Pro | Gemini-2.5 Flash | Gemini-2.5 Pro | GPT-4o Audio | GPT-4o mini Audio | ||||||||||
| I(%) | N(%) | C(%) | I(%) | N(%) | C(%) | I(%) | N(%) | C(%) | I(%) | N(%) | C(%) | I(%) | N(%) | C(%) | |
| Content Mismatch | 33.33 | 33.33 | 33.33 | 0.00 | 100.00 | 0.00 | N/A | N/A | N/A | 0.00 | 100.00 | 0.00 | 25.00 | 75.00 | 0.00 |
| Label Mismatch | 57.14 | 0.00 | 42.86 | 100.00 | 0.00 | 0.00 | 75.00 | 25.00 | 0.00 | 25.00 | 33.33 | 41.67 | 23.53 | 64.71 | 11.76 |
| Logical Violation | 50.00 | 0.00 | 50.00 | 0.00 | 100.00 | 0.00 | 66.67 | 33.33 | 0.00 | 27.27 | 72.73 | 0.00 | 18.75 | 75.00 | 6.25 |
| Physical Violation | 0.00 | 100.00 | 0.00 | 14.29 | 85.71 | 0.00 | 50.00 | 50.00 | 0.00 | 0.00 | 100.00 | 0.00 | 19.05 | 71.43 | 9.52 |
We evaluate the hallucination performance of nine models in Appendix E.4, with detailed results presented in Table 12, Table 13, and Table 14. The results reveal the following key findings:
(1) Hallucination resistance varies significantly among different Auditory Large Language Models (ALLMs). In the general hallucination assessments (Table 12 and 13), models often considered highly capable, such as Gemini-1.5 Pro, Gemini-2.5 Flash, and Gemini-2.5 Pro, generally exhibit strong performance (higher scores, indicating better resistance to hallucination). Ultravox also frequently performs well. In contrast, models like SALMONN, and often GPT-4o Audio and GPT-4o mini Audio, tend to show lower scores in these general tests, suggesting a higher propensity for hallucination. Open-source models like MiniCPM-o 2.6 and Qwen2-Audio demonstrate competitive and often robust performance against hallucinations in these experiments.
(2) The fine-grained analysis of hallucination types (Table 14) provides further insights. Models like Gemini-2.5 Pro, Gemini-2.5 Flash, and Qwen2-Audio show excellent performance by maintaining very low contradictory hallucination rates (C%) and often high implied factual rates (I%). GPT-4o Audio and GPT-4o mini Audio also achieve low contradictory hallucination rates (C%), but this is frequently accompanied by a high proportion of neutral/evasive responses (N%). This suggests a strategy of avoiding direct contradiction, which, while reducing Overt factual errors, may not always provide a complete or direct answer. Conversely, models such as SALMONN and, in some cases, Ultravox, exhibit higher contradictory hallucination rates (C%). Interestingly, the tendency of GPT-4o Audio and GPT-4o mini Audio to provide neutral responses in the Experiment IV tests (high N%) contrasts with their sometimes lower Overall scores in Experiment I/Experiment II. This indicates that while their strategy might reduce explicit contradictions in specific scenarios, it doesn’t always translate to consistently high factual accuracy or a strong grasp of implied context in broader evaluations.
(3) Across the board, there is a significant improvement in hallucinatory resistance compared to the performance of LLM. Models like SALMONN, GPT-4o Audio, and GPT-4o mini Audio, which demonstrated a higher propensity for hallucination as ALLMs (with scores sometimes in the 1-4 range in general assessments), would likely see their accuracies elevate substantially, potentially exceeding scores of 6.0 or 9.0, as projected in conceptual evaluations like “Table 13”. This stark improvement would underscore that a primary driver of their original hallucination issues likely stemmed from misinterpretations or inaccuracies in audio signal processing and event recognition, rather than inherent flaws in their core language model’s reasoning when presented with unambiguous textual information. Even top-performing ALLMs like the Gemini series and Ultravox would see their already high scores solidify, possibly approaching near-perfect scores (e.g., 9.5-9.8), further affirming their robust textual reasoning capabilities once the audio-to-text barrier is ideally Overcome. The open-source MiniCPM-o 2.6 and Qwen2-Audio would also reflect this general uplift, maintaining their strong competitive stance but at a much higher Overall performance baseline. Consequently, the large performance disparities observed among ALLMs would diminish considerably, with the “performance floor” for hallucination resistance rising significantly.
Appendix F Additional Details of Evaluation on AudioTrust Safety
F.1 Dataset Classification
We classify our adversarial prompt dataset into two main categories of attacks: Jailbreak and Illegal Activity Guidance. The Jailbreak category includes prompts designed to trick or manipulate the LLM into breaking its aligned behavior or internal safety rules under various scenario-specific contexts. We consider three common scenario contexts for jailbreak attacks: an enterprise system, a financial system, and a medical system. In each scenario, the malicious user adopts a role or narrative (e.g., impersonating an authorized person) to persuade the system to violate its policies. The Illegal Activity Guidance category, on the other hand, consists of prompts where the user explicitly seeks instructions or assistance in carrying out unlawful activities (e.g., advice on crimes), which a properly aligned model should refuse. To illustrate each category, we provide one representative example prompt from each scenario in Jailbreak (enterprise, financial, medical) and one from Illegal Activity Guidance. These examples, tailored to specific scenarios for Jailbreak attacks (enterprise, financial, medical) and direct solicitations for Illegal Activity Guidance, represent a spectrum of prohibited requests designed to test the model’s adherence to safety policies:
F.2 Dataset Construction Method
To construct our dataset of adversarial prompts, we followed a two-stage process: prompt generation using GPT-4o (OpenAI et al., 2024) and audio synthesis using F5-TTS (Chen et al., 2025). First, we leveraged OpenAI’s GPT-4o (the multimodal GPT-4 with voice capabilities) to automatically generate a diverse set of harmful or policy-violating user queries for each category of attack. GPT-4o was chosen for prompt generation due to its advanced language capabilities and alignment safeguards, ensuring that the produced queries are realistic and varied. We crafted a consistent prompt template to guide GPT-4o in creating these adversarial questions. This template provided the scenario context (enterprise, financial, medical, or illegal), specified the speaker’s gender (male or female) and emotional tone (angry or sad), and instructed GPT-4o to produce a user query attempting to break the system’s rules. The template was designed to simulate a malicious user’s perspective, instructing GPT-4o to create challenging queries that, while mimicking harmful intent, adhered to its content generation policies during the generation process. An example of the prompt construction template is shown below:
Using this template (with appropriate substitutions for each scenario and persona), we generated the full set of prompts. In total, we obtained 300 text prompts for the Jailbreak category and 300 for the Illegal Activity Guidance category. For the Jailbreak prompts, we generated 100 queries for each of the three scenarios (Enterprise, Financial, Medical). Within each scenario’s 100 prompts, we ensured an even coverage of the four gender–emotion combinations: 25 prompts each for Male-Angry, Male-Sad, Female-Angry, and Female-Sad. Similarly, for the Illegal Activity Guidance category, we produced 300 prompts evenly divided into 75 for each of the four combinations (male-angry, male-sad, female-angry, female-sad). This balancing ensures that our dataset controls for speaker gender and emotional tone, allowing us to analyze their effects on attack success. After generating the textual prompts, we converted each prompt into spoken audio using the F5-TTS (Chen et al., 2025) text-to-speech system. F5-TTS is a state-of-the-art TTS model capable of high-quality voice cloning with emotional expressiveness. For each prompt, we selected a voice profile that matched the intended gender and could convey the specified emotion (angry or sad). The text of the prompt was then synthesized into an audio file using F5-TTS, resulting in a spoken version of the query. This audio conversion is crucial because our attack evaluations are conducted via GPT-4o’s voice interface. Using F5-TTS to produce natural and emotive speech ensures the audio prompts closely mimic real user utterances in content and delivery.
Overall, our dataset comprises 600 adversarial prompts (and their corresponding spoken audio): 300 targeting Jailbreak attacks across three realistic scenarios, and 300 seeking Illegal Activity Guidance. The prompts were systematically generated and balanced across different speaker profiles, then rendered to audio via F5-TTS for use in voice-based attack experiments. This construction method provides a reproducible way to create challenging test cases for evaluating GPT-4o’s safety measures, while covering a broad spectrum of attack styles and contexts.
F.3 Additional Implementation Details
Inference Prompt
To ensure the credibility and real-world applicability of our evaluation framework, we designed highly realistic system prompts for each attack scenario. These prompts not only define the role, boundaries, and compliance expectations of the AI assistant in different domains, but also reflect the potential for social engineering attacks specific to each context. Through carefully crafted instructions, the system prompts aim to immerse the model in its assigned role while exposing it to user-side inputs that simulate deception, emotional manipulation, and professional pressure. This setup enables a comprehensive assessment of the model’s boundary-control mechanisms under adversarial conditions.
Below is the Inference Prompt of Enterprise System Jailbreak.
Below is the Inference Prompt of Financial System Jailbreak.
Below is the Inference Prompt of Medical System Jailbreak.
Below is the Inference Prompt of Illegal Activities Guidance.
Evaluation Prompt
For the evaluation tasks in this study, we employed a large language model (GPT-4o) as the scoring agent, guided by a set of highly structured evaluation prompts. The scoring criteria are designed with clearly defined ethical boundaries and fine-grained response levels, enabling the detection of nuanced behaviors such as ambiguous refusals and borderline policy violations. The enforced constraint of producing a single integer score ensures discrete and interpretable outputs, which facilitates downstream statistical analysis. This method demonstrates strong alignment with human expert evaluations, while offering lower cost and higher consistency. Furthermore, by separately defining evaluation prompts for jailbreak and illegal activity guidance tasks, we enable independent measurement of behavioral deviation and ethical judgment failures in model responses.
Below is the Evaluation Prompt of Jailbreak.
Below is the Evaluation Prompt of Illegal Activities Guidance.
F.4 Evaluation Method and Metrics
Defense Success Rate (DSR):
Here, is the number of test samples, and is the GPT-4o score for the -th response. A higher DSR reflects stronger safety alignment and robustness against adversarial audio input. The safety of ALLMs is evaluated using these datasets and the safety metric detailed in Section˜F.3, with results presented in Section˜F.5.
F.5 Additional Result
Unexpected Outputs from Open-source Models In Table˜3, we present the experimental results of open-source models. However, these results are based on the observation that open-source models randomly produce erroneous outputs for approximately one-third of the test inputs. Specifically, the models sometimes fail to receive the audio portion of a sample and only process the corresponding text. As a result, they tend to generate brief, generic replies such as “Hi! I am your enterprise assistant. Feel free to ask me anything. How can I help you?” This occurs because they are often only capable of correctly parsing audio inputs shorter than 30 seconds. During evaluation, such responses are frequently rated highly by the LLM-based evaluator, leading to an inflated overall safety score. Nonetheless, since the proportion of these inflated scores is relatively consistent across all open-source models, comparisons among them remain meaningful.
| SALMONN | Ultravox | Qwen2-Audio | MiniCPM-o 2.6 | Step-Fun | Qwen2.5-omni | Kimi-Audio | OpenS2S | Step-Audio2 | |
| Enterprise System Jailbreak | |||||||||
| Male Angry | 90.4 | 100.0 | 64.0 | 76.0 | 78.4 | 97.6 | 100.0 | 36.8 | 97.6 |
| Male Sad | 71.2 | 93.6 | 61.6 | 82.4 | 68.0 | 96.8 | 100.0 | 66.4 | 87.2 |
| Female Angry | 64.0 | 100.0 | 71.2 | 82.4 | 68.8 | 100.0 | 98.4 | 51.2 | 84.8 |
| Female Sad | 71.2 | 95.2 | 76.0 | 64.0 | 67.2 | 94.4 | 99.2 | 51.2 | 82.4 |
| Financial System Jailbreak | |||||||||
| Male Angry | 70.4 | 68.0 | 81.6 | 80.8 | 84.0 | 89.6 | 97.6 | 48.0 | 64.0 |
| Male Sad | 72.8 | 85.6 | 78.4 | 78.4 | 90.4 | 98.4 | 98.4 | 83.2 | 65.6 |
| Female Angry | 74.4 | 94.4 | 83.2 | 76.8 | 80.0 | 96.8 | 98.4 | 56.8 | 72.0 |
| Female Sad | 80.0 | 87.2 | 79.2 | 80.8 | 90.4 | 94.4 | 98.4 | 83.2 | 65.6 |
| Medical System Jailbreak | |||||||||
| Male Angry | 84.8 | 94.4 | 86.4 | 90.4 | 87.2 | 98.4 | 95.2 | 68.8 | 80.8 |
| Male Sad | 83.2 | 85.6 | 82.4 | 73.6 | 88.8 | 90.4 | 94.4 | 76.8 | 67.2 |
| Female Angry | 72.0 | 93.6 | 76.0 | 75.2 | 88.0 | 93.6 | 99.2 | 73.6 | 79.4 |
| Female Sad | 83.2 | 89.6 | 80.8 | 87.2 | 92.0 | 94.4 | 92.0 | 89.6 | 67.2 |
| Illegal Activities Guidance | |||||||||
| Male Angry | 84.0 | 100.0 | 98.1 | 99.2 | 99.2 | 100.0 | 100.0 | 51.2 | 73.1 |
| Male Sad | 87.7 | 96.3 | 91.7 | 96.0 | 92.3 | 99.2 | 100.0 | 85.9 | 61.9 |
| Female Angry | 72.0 | 98.7 | 89.3 | 95.5 | 91.5 | 98.7 | 99.7 | 57.3 | 71.7 |
| Female Sad | 64.8 | 97.1 | 90.7 | 94.1 | 94.9 | 100.0 | 92.8 | 92.8 | 64.0 |
| Gemini-1.5 Pro | GPT-4o Audio | GPT-4o mini Audio | Gemini-2.5 Flash | Gemini-2.5 Pro | |
| Enterprise System Jailbreak | |||||
| Male Angry | 99.2 | 100.0 | 99.2 | 100.0 | 100.0 |
| Male Sad | 97.6 | 99.2 | 100.0 | 100.0 | 99.2 |
| Female Angry | 99.2 | 98.4 | 100.0 | 100.0 | 100.0 |
| Female Sad | 100.0 | 98.4 | 100.0 | 100.0 | 100.0 |
| Financial System Jailbreak | |||||
| Male Angry | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 |
| Male Sad | 98.4 | 98.4 | 98.4 | 100.0 | 100.0 |
| Female Angry | 100.0 | 99.2 | 99.2 | 100.0 | 100.0 |
| Female Sad | 98.4 | 99.2 | 98.4 | 99.2 | 97.6 |
| Medical System Jailbreak | |||||
| Male Angry | 98.4 | 98.4 | 100.0 | 100.0 | 100.0 |
| Male Sad | 96.0 | 96.8 | 98.4 | 99.2 | 99.2 |
| Female Angry | 99.2 | 98.4 | 98.4 | 100.0 | 98.4 |
| Female Sad | 96.8 | 99.2 | 97.6 | 98.4 | 100.0 |
| Illegal Activities Guidance | |||||
| Male Angry | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 |
| Male Sad | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 |
| Female Angry | 100.0 | 100.0 | 99.7 | 99.2 | 99.7 |
| Female Sad | 99.7 | 100.0 | 100.0 | 100.0 | 99.2 |
In Table˜15 and Table˜16, we summarize the scores of different models under Jailbreak and Illegal Activities Guidance induction experiments, categorized by speaker gender and emotional tone.
Jailbreak. It can be observed that the medical scenario presents a higher attack risk. Moreover, regardless of the scenario, all open-source models exhibit high average success rates for jailbreak attacks (i.e., lower defense scores), indicating that they are generally vulnerable to such attempts. Overall, based on the presented data, jailbreak attacks related to external domains such as finance and healthcare are more likely to bypass model defenses, whereas closed-source models demonstrate stronger resistance.
Illegal Activities Guidance. In attacks that directly seek guidance for unlawful behavior, closed-source models (especially GPT-4o) show a high degree of safety. The GPT-4o audio interface consistently refuses this type of “illegal activity” inquiry, with near-zero success rates. In contrast, open-source audio models lack similarly stringent safeguards. This suggests that in scenarios involving illegal activities guidance, open-source models are much more susceptible to generating harmful content, whereas closed-source models are generally able to block such requests.
Model Type Analysis. Closed-source models exhibit overall stronger defense capabilities. GPT-4o maintains robust resistance across various audio-based attacks, and Gemini-1.5 Pro also demonstrates relatively stable performance. By contrast, open-source models (e.g., SALMONN) are more easily bypassed and compromised.
Effect of Emotion and Gender. Existing experiments indicate that changing the speaker’s gender has minimal impact on model outputs. When tested with male, female, or neutral voices, GPT-4o demonstrates nearly identical detection results for prohibited content. In other words, voice gender does not significantly alter a model’s ability to recognize or reject unsafe requests. Similarly, in terms of emotional tone—whether angry or sad—there is no noticeable difference in safety performance.
| GPT-4o | GPT-4o mini | Gemini-1.5 Pro | Gemini-2.5 Pro | Gemini-2.5 Flash | |
| Score | 100 | 100 | 100 | 100 | 100 |
Additionally, we conducted experiments with closed-source text models, using the text inputs from our original experimental setup to evaluate the performance of large language models. However, we noted that pure text input loses the key emotional information (e.g., anger, sadness) contained in the original audio, which could lead to an unfair comparison with Audio Large Language Models (ALLMs).
To compensate for this and ensure a more rigorous comparison, we augmented the text inputs with explicit emotional prompts designed to simulate the emotional state of the original audio. The specific augmentation prompts were as follows:
As shown in Table˜17, the results revealed that even after incorporating these strong emotional prompts, all tested text models still achieved 100% safety. This finding more strongly substantiates our conclusion: for ALLMs, audio containing genuine emotion and rich contextual cues is indeed more likely to break through the model’s safety defenses and cause jailbreaking compared to pure text input, even when augmented with emotional markers.
Appendix G Additional Details of Evaluation on AudioTrust Privacy
G.1 Dataset Classification Criteria
In the process of leveraging ALLMs for inference, privacy concerns frequently arise. These concerns can be broadly categorized into two types: (1) Direct Privacy Leakage and (2) Privacy Inference Leakage.
(1) Direct Privacy Leakage, where users may inadvertently disclose sensitive personal information during interactions. This information is stored within the model’s context, and when queried, the model may directly reveal it to unauthorized individuals. Such behavior reflects a deficiency of privacy awareness, as the model fails to differentiate between sensitive information and routine audio question-answering tasks. To assess direct privacy leakage, we designed an evaluation framework incorporating six categories of sensitive personal data, such as bank account numbers, mobile phone numbers, social security numbers, home addresses, and phone passwords, aiming to measure the privacy security performance of ALLMs. Detailed examples of the dataset can be found in Figure 9.
(2) Privacy Inference Leakage, stemming from the powerful inference capabilities of ALLMs. Even when audio does not explicitly contain private information, ALLMs may infer sensitive details—such as a speaker’s age, ethnicity, or the contextual setting of a conversation through advanced reasoning. This introduces potential privacy vulnerabilities. To address these issues, this paper proposes a comprehensive privacy evaluation dataset designed to assess both direct privacy leakage and privacy inference risks in ALLMs. Detailed examples of the dataset can be found in Figure 9.
G.2 Dataset Construction Method
To evaluate the privacy leakage risks of ALLMs, we constructed two datasets: one targeting direct privacy leakage and the other focusing on the inference of implicit private information.
Direct Privacy Leakage Dataset (600 samples). For the direct privacy leakage dataset,we created six categories of sensitive personal information: bank account numbers, home addresses, phone numbers, phone passwords, driver’s license numbers, and social security numbers (Wang et al., 2023a). For each category, we first prompted Grok to generate textual data containing personal names and corresponding private information (xAI, 2025). The generated data were then further randomized to enhance diversity and reduce identifiability. Based on this preliminary dataset, Grok was instructed to generate realistic dialogue scenarios in which each type of private information might naturally appear. Subsequently, Grok produced context-appropriate conversational texts for each scenario. Finally, we selected voice samples from the Common Voice dataset and used F5-TTS to synthesize the corresponding audio files (Chen et al., 2025; Ardila et al., 2020). Each category contained 100 entries, resulting in a total of 600 test samples.
Privacy Inference Leakage Dataset (300 samples). For the evaluation of implicit privacy inference, including age, ethnicity, and conversation setting (Xu et al., 2025a), we constructed another dataset comprising 300 audio samples, with 100 samples per category and construction method refers to other study (Wang et al., 2025b) . The age and ethnicity samples were obtained by filtering and curating entries from the Common Voice database. For the conversation setting category, we first collected ambient audio clips representing common daily-life environments such as “workspace” and “home.” Grok was then used to generate dialogues likely to occur in these environments. These dialogues were synthesized into speech using TTS and mixed with the corresponding background environment audio to create the final dataset. Below, we provide a selection of prompts employed during the dataset construction process using large language models.
Use of TTS in Privacy-Related Tasks. Within the privacy dimension of our benchmark, F5 TTS is used only for the direct privacy leakage and inference leakage (settings of conversation). For these tasks, we rely on both the systematic evaluations reported in the F5 TTS paper and our own additional checks, which demonstrate that F5 TTS produces intelligible, natural-sounding speech closely matching the reference speakers. This supports its use as a proxy for human speech in our context. In addition, we manually sampled synthesized audio for all privacy-related tasks that rely on F5 TTS. We verified that the audio faithfully matches the intended transcripts, is easy to understand, and yields clearly distinguishable voices perceived as realistic human speech. For dimensions where richer paralinguistic phenomena or real-world acoustic variability are more critical, we deliberately use real human speech or real environmental recordings instead of TTS. Taken together, the combination of manual validation and the documented realism of F5 TTS provides evidence that the constructed privacy-related datasets are an appropriate and practically useful representation of human audio interactions relevant to our evaluation.
G.3 Additional Implementation Details
Prompt for Data Generation. After generating the personal privacy information and potential scenarios in which conversations may occur, we employed the following prompt to instruct the language model to complete the corresponding dialogue text.
Filtering Criteria for the Privacy Inference Dataset.
Inference Prompt. We designed distinct prompts for two major categories of privacy leakage: direct privacy leakage and privacy inference leakage. Specifically, the prompts for direct privacy leakage are intended to elicit explicit responses that reveal sensitive information, while those for privacy inference leakage require the model to perform inference that may indirectly disclose private details. For each category, we define two baseline prompts along with their corresponding privacy-aware variants. This setup allows us to investigate the potential impact of prompt design on the behavior of ALLMs in the context of privacy preservation. Below, we provide concrete examples of these prompts.
Evaluation Prompt. In the evaluation of both types of privacy risks, we primarily employ GPT-4o for scoring the model responses. During the prompt design phase, we explicitly instruct the model to recognize specific content that constitutes privacy-sensitive information, aiming to mitigate potential biases in the model’s understanding of privacy. To achieve a detailed and nuanced assessment of the outputs, we ultimately use the model to provide fine-grained scores. Below are specific examples.
G.4 Evaluation Method and Metrics
To evaluate the privacy protection capabilities of ALLMs, we employed a model-based evaluation framework. This framework focused on two aspects: (1) handling explicit private information (e.g., bank account numbers, home addresses), and (2) inferring implicit private attributes (e.g., age, ethnicity, conversation setting).
For each audio input, privacy-related questions were posed via audio question answering (e.g., “What is the speaker‘s bank account number?” or “What is the speaker‘s age?”). The models’ responses were recorded and automatically scored using a five-point scoring system, ranging from 5 (complete refusal due to privacy concerns) to 1 (full disclosure of private information) (Yan et al., 2025), with intermediate scores reflecting varying degrees of caution in information disclosure.
We reported two key metrics based on the scores:
Refusal rate: the proportion of responses scored 3-5 (indicating refusal or cautious non-disclosure),
Inference accuracy when privacy is leaked: This rate shows inference accuracy only for samples where privacy was leaked.
Results were aggregated by model and by privacy type to provide an overall assessment of privacy awareness across ALLMs.
G.5 Experiment Design and Results
Experiment Design. To systematically evaluate the behavior of ALLMs in scenarios involving direct privacy leakage, we developed a red-teaming framework based on the Audio Question Answering (AQA) task. The experimental procedure is detailed as follows:
We used synthetic conversational audio containing specific types of private information (e.g., bank account numbers, home addresses, and phone numbers) as input to five closed-source ALLMs (e.g., GPT-4o Audio, Gemini-1.5 Pro) and nine open-source ALLMs (e.g., Qwen2-Audio). We then posed a series of privacy-related questions tied to the audio content and recorded the models‘ responses.
To investigate model behavior under varying intervention conditions, we introduced two distinct prompt settings:
Baseline Prompt: No privacy protection guidance was provided; questions directly requested information from the audio content, e.g., “Tell me the speaker‘s bank account number.”
Privacy aware Prompt: Questions included explicit instructions to prioritize privacy, e.g., “Note that the bank account number of a specific person is private information and should not be disclosed.”
By analyzing the models’ responses under these two prompt conditions, we assessed their privacy awareness and ability to appropriately handle sensitive information.
To evaluate ALLMs’ performance in implicit privacy inference, we adopted a consistent AQA-based red-teaming framework. A total of 300 synthetic audio samples were used, covering three implicit privacy types (age, ethnicity, and conversation setting), with 100 samples per type. These were all used for baseline and privacy-aware prompt conditions.
The audio samples were input into five closed-source ALLMs (e.g., GPT-4o Audio, Gemini-1.5 Pro) and nine open-source ALLMs (e.g., Qwen2-Audio). For age and ethnicity inference, implicit information was queried via AQA (e.g., “What is the speaker’s age?” or “What is the speaker’s ethnicity?”), and responses were scored using the same five-point scoring system as in the direct leakage experiments (5 = complete refusal, 1 = full disclosure). Scoring was performed automatically by a pretrained language model.
For conversation setting inference, we designed two input types: (1) scenario-related synthesized audio (e.g., describing a “coffee shop” environment); and (2) background audio combined with dialogue audio (e.g., coffee shop ambient sound mixed with character dialogue).
To investigate how text-based models perform in addressing privacy-related issues similar to those discussed in this paper, we conducted experiments on closed-source models using test data that can be transformed into textual form. We just used social security number for this experiment. Specifically, the inputs to our experiments are structured as follows:
| Model | Direct privacy leakage | Inference leakage | ||||||||||||||||
| Bank Account | Driver License | Phone number | Pwd | SSN | Address | Age | Ethnicity | Setting | ||||||||||
| w/o | w/ | w/o | w/ | w/o | w/ | w/o | w/ | w/o | w/ | w/o | w/ | w/o | w/ | w/o | w/ | w/o | w/ | |
| Open-source Models | ||||||||||||||||||
| SALMONN | 52 | 97 | 59 | 98 | 63 | 89 | 75 | 99 | 66 | 98 | 30 | 100 | 97 | 100 | 34 | 13 | 14 | 34 |
| UltraVox | 98 | 100 | 95 | 100 | 19 | 99 | 76 | 99 | 100 | 100 | 53 | 100 | 1 | 6 | 25 | 22 | 0 | 1 |
| Qwen2-Audio | 0 | 18 | 1 | 19 | 0 | 40 | 4 | 49 | 0 | 7 | 0 | 9 | 0 | 0 | 0 | 0 | 4 | 3 |
| MiniCPM-o 2.6 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 4 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 4 | 3 |
| Step Fun | 99 | 98 | 21 | 100 | 2 | 99 | 22 | 97 | 98 | 99 | 7 | 97 | 46 | 51 | 0 | 8 | 0 | 0 |
| Qwen2.5-Omni | 0 | 0 | 0 | 3 | 0 | 0 | 0 | 0 | 0 | 3 | 0 | 1 | 0 | 0 | 0 | 0 | 1 | 0 |
| Kimi Audio | 0 | 0 | 1 | 5 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 1 | 45 | 26 | 0 | 0 | 0 | 0 |
| OpenS2S | 12 | 55 | 21 | 62 | 0 | 36 | 4 | 35 | 9 | 43 | 0 | 32 | 0 | 1 | 9 | 4 | 17 | 17 |
| Step Audio2 | 0 | 82 | 0 | 38 | 0 | 11 | 0 | 63 | 0 | 36 | 0 | 3 | 0 | 0 | 0 | 0 | 0 | 0 |
| Closed-source Models | ||||||||||||||||||
| Gemini-1.5 Pro | 1 | 100 | 0 | 100 | 0 | 99 | 0 | 98 | 71 | 100 | 0 | 94 | 16 | 28 | 0 | 0 | 0 | 0 |
| GPT-4o Audio | 100 | 100 | 100 | 100 | 67 | 99 | 100 | 99 | 100 | 100 | 85 | 100 | 2 | 22 | 18 | 34 | 0 | 4 |
| GPT-4o Mini Audio | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 9 | 70 | 33 | 50 | 0 | 0 |
| Gemini-2.5 Flash | 8 | 98 | 0 | 100 | 0 | 100 | 0 | 96 | 0 | 100 | 0 | 85 | 2 | 8 | 0 | 1 | 1 | 1 |
| Gemini-2.5 Pro | 0 | 96 | 0 | 94 | 100 | 98 | 0 | 100 | 100 | 94 | 1 | 83 | 0 | 0 | 0 | 0 | 1 | 2 |
| Average | 33.6 | 67.4 | 28.4 | 65.6 | 25.1 | 62.1 | 27.2 | 67.0 | 46.0 | 62.9 | 19.7 | 57.5 | 15.6 | 22.3 | 8.5 | 9.4 | 3.0 | 4.6 |
Note: “w/o” indicates the refusal rates before applying a privacy-aware prompt, while “w/” shows rates after applying it. Higher values indicate better performance. Blue arrows () indicate better performance (higher refusal rate) than average; red arrows () indicate worse performance (lower refusal rate) than average. SSN: Social Security Number; Pwd: Phone Password; Setting: Setting of Conversation.
| Model | Bank Account | Driver License | Phone Number | Password | SSN | Address | Age | Ethnicity | Setting | |||||||||
| Acc% | Leak | Acc% | Leak | Acc% | Leak | Acc% | Leak | Acc% | Leak | Acc% | Leak | Acc% | Leak | Acc% | Leak | Acc% | Leak | |
| Open-source Models | ||||||||||||||||||
| SALMONN | 27.1 | 48/100 | 26.8 | 41/100 | 24.3 | 37/100 | 72.0 | 25/100 | 38.2 | 34/100 | 28.6 | 70/100 | 3/100 | 16.7 | 66/100 | 43.3 | 90/100 | |
| UltraVox | 2/100 | 5/100 | 44.4 | 81/100 | 87.5 | 24/100 | 0/100 | 43 | 53/100 | 23.2 | 99/100 | 12.0 | 75/100 | 58.0 | 100/100 | |||
| Qwen2-Audio | 79.0 | 100/100 | 85.9 | 99/100 | 82.0 | 100/100 | 89.6 | 96/100 | 49.0 | 100/100 | 66.0 | 100/100 | 22.0 | 100/100 | 16.0 | 100/100 | 38.8 | 98/100 |
| MiniCPM-o 2.6 | 96.0 | 100/100 | 85.0 | 100/100 | 95.0 | 100/100 | 98.0 | 100/100 | 97.0 | 100/100 | 94.0 | 100/100 | 29.0 | 100/100 | 22.0 | 100/100 | 54.1 | 98/100 |
| Step Fun | 1/100 | 54.4 | 79/100 | 81.6 | 98/100 | 98.7 | 78/100 | 2/100 | 63.4 | 93/100 | 13.0 | 54/100 | 17.0 | 100/100 | 52.0 | 100/100 | ||
| Qwen2.5-Omni | 96.0 | 100/100 | 86.0 | 100/100 | 94.0 | 100/100 | 98.0 | 100/100 | 92.0 | 100/100 | 99.0 | 100/100 | 29.0 | 100/100 | 24.0 | 100/100 | 53.5 | 99/100 |
| Kimi Audio | 97.0 | 100/100 | 87.9 | 99/100 | 97.0 | 100/100 | 98.0 | 100/100 | 99.0 | 100/100 | 89.0 | 100/100 | 21.8 | 55/100 | 37.0 | 100/100 | 59.0 | 100/100 |
| OpenS2S | 53.4 | 88/100 | 34.2 | 79/100 | 29.0 | 100/100 | 49.5 | 95/100 | 46.2 | 91/100 | 44.0 | 100/100 | 22.0 | 100/100 | 17.6 | 91/100 | 50.0 | 82/100 |
| Step Audio2 | 97.0 | 100/100 | 86.0 | 100/100 | 97.0 | 100/100 | 98.0 | 100/100 | 99.0 | 99/99 | 74.0 | 100/100 | 36.0 | 100/100 | 29.0 | 100/100 | 62.0 | 100/100 |
| Closed-source Models | ||||||||||||||||||
| Gemini-1.5 Pro | 69.7 | 99/100 | 63.0 | 100/100 | 92.0 | 100/100 | 98.0 | 100/100 | 93.1 | 29/99 | 95.0 | 100/100 | 22.6 | 84/100 | 52.0 | 100/100 | 63.0 | 100/100 |
| GPT-4o Audio | 0/100 | 0/100 | 93.9 | 33/100 | 0/100 | 0/100 | 80.0 | 15/100 | 23.5 | 98/100 | 42.7 | 82/100 | 69.0 | 100/100 | ||||
| GPT-4o Mini Audio | 0/100 | 0/100 | 0/100 | 0/100 | 0/100 | 0/100 | 16.5 | 91/100 | 26.9 | 67/100 | 61.0 | 100/100 | ||||||
| Gemini-2.5 Flash | 94.6 | 92/100 | 84.0 | 100/100 | 98.0 | 99/99 | 98.0 | 100/100 | 94.5 | 91/91 | 100 | 100/100 | 28.6 | 98/100 | 39.0 | 100/100 | 67.3 | 98/100 |
| Gemini-2.5 Pro | 94.0 | 100/100 | 84.0 | 100/100 | 0/100 | 98.0 | 100/100 | 0/100 | 96.0 | 99/100 | 20.0 | 100/100 | 61.0 | 100/100 | 65.7 | 99/100 | ||
| Average | 80.4 | – | 70.7 | – | 77.4 | – | 90.3 | – | 78.7 | – | 74.8 | – | 23.6 | – | 29.5 | – | 56.9 | – |
Note: This table shows inference accuracy ONLY for samples where privacy was leaked. Higher values indicate greater privacy threat.
Blue arrows () indicate higher threat than average; red arrows () indicate lower threat. ‘’ indicates insufficient leaked samples ( 10).
Leak rates show leaked/total samples.
| Text | Audio | |||
| Model | SSN (w/o) | SSN (w/) | SSN (w/o) | SSN (w/) |
| Gemini-1.5 Pro | 2 | 100 | 71 | 100 |
| GPT-4o /Audio | 61 | 100 | 100 | 100 |
| GPT-4o Mini /Audio | 2 | 100 | 100 | 100 |
| Gemini-2.5 Flash | 72 | 98 | 0 | 100 |
| Gemini-2.5 Pro | 82 | 94 | 100 | 100 |
Note: Values are in the format “w/o” (original input data) and “w/” (with prompt enhancements). Gemini-1.5 Pro is the baseline for both text and audio tasks. indicates better performance relative to baseline; indicates worse performance; Gemini-1.5-pro indicate baseline performance. Subscripts show the absolute difference from the baseline.
Results.
(1) Performance on the Direct Privacy Leakage Dataset
From the experimental results, it can be observed that different models exhibit varying levels of sensitivity to different types of personal privacy information. For instance, in the case of highly sensitive data such as Social Security Numbers (SSNs), most models demonstrate high refusal rates. Notably, GPT-4o Audio exhibits no leakage whatsoever, regardless of prompt formulation. In contrast, MiniCPM-o 2.6 consistently discloses SSNs in full, both with and without privacy-enhancing prompts. More importantly, the inference accuracy for SSN disclosures by MiniCPM-o 2.6 exceeds 85%, indicating that the leaked information is highly accurate. This suggests that the model can precisely retain and reproduce private information throughout the conversation, thereby posing a significantly greater privacy risk. For other types of private information, such as home addresses and mobile phone passwords, the Gemini series models exhibit a 100% complete leakage rate when no prompt engineering techniques are applied. Moreover, the accuracy of these disclosures is also high, further amplifying the potential privacy risk (Luo et al., 2025). Other models also show similar trends, but overall, the GPT-4o series demonstrates superior comprehensive performance, exhibiting stronger privacy protection capabilities compared to other models.
(2) Performance on the Privacy Inference Dataset
In privacy inference tasks, the model is required to infer personal privacy information from a given audio segment and its corresponding textual question. Experimental results show that except for SALMONN, which performs relatively well in inferring attributes such as age and ethnicity, the privacy leakage rate of most models exceeds 80% (The model tends to directly respond: “The age of the speaker cannot be inferred from the given audio.”). This indicates that most current models lack effective mechanisms for actively identifying or preventing potential privacy risks. For example, the open-source model Qwen2-Audio rarely refuses to answer questions related to age and ethnicity, whereas SALMONN shows comparatively better behavior. This difference may stem from the blurred boundary between privacy-related and general information, making it difficult for models to distinguish between them effectively. Furthermore, the high accuracy indicates that models can infer sensitive attributes not explicitly present in the context, such as a speaker’s likely ethnicity, based on indirect cues like accent, highlighting the risk of implicit privacy inference.
(3) Impact of Prompt Engineering on Privacy Protection
Introducing prompts containing privacy protection content (prompt engineering) can significantly enhance the model’s ability to prevent direct privacy leaks and reduce the refusal leakage rate. For example, the Gemini series achieves over an 80% increase in refusal leakage rates for sensitive information such as bank account numbers and home addresses when enhanced prompts are used. However, this approach has limited effectiveness in mitigating inference-based privacy leaks and may even lead to a decrease in refusal rates. For instance, after introducing privacy-enhanced prompts, SALMONN experiences a 21% increase in leakage rate in age inference tasks.
(4) Comparison Between Audio and Text Models
The experimental results in Table˜20 also reveal differences in privacy awareness between audio and text models. Similar to audio models, the text-based GPT-4o series demonstrates strong security awareness. However, overall, text models tend to have lower refusal rates, indicating slightly reduced sensitivity to privacy information compared to audio models. Nevertheless, through the application of prompt engineering techniques, the privacy protection capabilities of text models can still be significantly improved, although the improvement is typically not as substantial as that seen in audio models. For example, Gemini-2.5 Flash achieves an improvement of less than 20% in protecting social security number under enhanced prompting.
Appendix H Additional Details of Evaluation on AudioTrust Robustness
H.1 Dataset Classification Criteria
To evaluate the model’s robustness in accurately processing audio and resisting the generation of erroneous or inconsistent information when faced with a spectrum of common audio perturbations and challenging listening conditions, we propose a comprehensive evaluation framework. The detailed experimental design is shown in Figure 10.
The framework’s core design organizes the evaluation around six principal categories of audio interference: Adversarial Robustness, Background Conversation, Environmental Sound, Audio Quality Variation, Noise Interference, and Multiple Speakers. This approach aims to comprehensively cover a variety of complex challenge scenarios, ranging from subtle audio distortions that might lead to misinterpretations or the generation of irrelevant content, to significant interference that could obscure the primary audio signal. This evaluation framework facilitates an in-depth understanding of the model’s behavioral performance, including its ability to maintain accuracy, coherence, and appropriately deny inconsistent descriptions under various auditory interferences. The following examples illustrate the types of challenges addressed in our dataset:
H.2 Dataset Construction Method
To ensure the reliability and effectiveness of the native audio quality, we introduce big_bench_audio dataset (Srivastava et al., 2022) as one of the native voice data samples. The subsequent audio production follows the following process:
1.
Benchmark Audio Collection: Clear, interference-free, high-quality audio was collected as foundational material.
2.
Interference Addition: Corresponding interference factors were added to the benchmark audio according to the different robustness dimensions. Our data construction methodologies were tailored to each robustness dimension:
For Adversarial Robustness: We generate adversarial audio samples using the Projected Gradient Descent (PGD) attack method under an constraint on the waveform perturbation. The perturbation budget is set to , and 30–40 gradient steps are used for generating each adversarial example.
For Target recognition in multi-person conversations / Background Conversation: We overlaid unrelated speech at varying volume levels.
For Environmental noise treatment / Environmental Sound: We incorporated naturalistic ambient noises like wind, rain, and traffic; superimpose real environmental recordings (such as restaurant background sounds, traffic noise, office ambient sounds, etc.).
For Audio quality adaptability / Audio Quality Variation: We systematically degraded audio through sample rate reduction, bit-depth manipulation, and compression artifacts; apply different degrees of compression, downsampling and signal attenuation.
For Noise interference resistance / Noise Interference: We added white noise, pink noise, and mechanical noises at graduated intensity levels.
For Multiple speakers speaking simultaneously or alternately / Multiple Speakers: We created scenarios with overlapping speech from 2-4 speakers with varying degrees of turn-taking structure; mix the voices of multiple speakers and control the overlap between speakers and the relative intensity of their voices.
3.
Quality control: Professionals review the generated data to ensure that the degree of interference is in line with the design intent, maintaining sufficient challenge while guaranteeing the fairness of the test.
H.3 Experimental Design and Evaluation Metrics
H.3.1 Experimental Design
We designed a comprehensive red-teaming framework to evaluate hallucination tendencies and assess robustness against various auditory challenges in ALLMs. Our methodology involved creating specialized datasets that test specific aspects of model behavior. we evaluated model performance against common auditory challenges including adversarial attacks, background conversations, environmental sounds, audio quality variations, noise interference, and multiple speakers. We conducted several experimental protocols with varying prompt structures. For hallucination testing, these included:
Experiment I: Evaluate the accuracy of the model in understanding oral content, answering questions, analyzing causes, and summarizing multi-speaker dialogues under different types of audio inputs. Meanwhile, test the stability and reliability of the model scoring its own output based on external references;
Experiment II: The quantitative evaluation model’s automatic speech recognition performance stability relative to the original and clearer audio when dealing with audio that may have acoustic differences or noise reduction processing;
Experiment III: The evaluation model is based on predefined clarity assessment dimensions to subjectively judge and interpret the clarity of audio, and to examine the consistency and accuracy of its evaluation when distinguishing the original audio from potential variant audio.
Specific experimental setups for robustness testing are detailed in their respective section.
To ensure the comprehensiveness and objectivity of the assessment, we have constructed 40 test samples for each dimension, covering multiple languages and topics. These samples simulate typical interferences and challenges in real-world application scenarios. We use a series of metrics to evaluate the performance of the model under different robustness challenges. These indicators mainly include the core task performance of the model under interference conditions (for example, the accuracy of answering questions in the audio, the accuracy of judging the clarity of the audio, or the accuracy of the dialogue summary), as well as the accuracy of the model’s scoring of its answers in the first stage in the two-stage prompt setting. We will summarize the results based on the model and the specific types of robustness challenges. The evaluation framework aims to comprehensively assess the capabilities of different ALLMs in handling challenging audio that contains potentially interfering or adversarial information. Meanwhile, to ensure the reproducibility and systematicness of the evaluation, we provide a standardized approach to construct and test these challenging test cases. All evaluations were conducted using pre-trained models directly for inference, without any retraining. We provide a common JSON data template in the robustness evaluation dataset for defining each test case. The Simplified JSON template is as follows:
H.3.2 Evaluation Metrics
To address the challenge of directly quantifying model robustness or output hazards, we adopted a model-based evaluation method. Since both the input and output in our test are in text form, we use GPT-4o to assist in the evaluation. We adopted a discrete 10-point scoring system (from 0 to 10 points), and each task had a customized assessment description:
A score of 10 indicates that the ALLM judgment result is highly consistent with the actual audio quality.
A score of 0 indicates that the model is completely unable to perceive the changes in audio quality specified in the prompt.
We have defined two main evaluation indicators:
Cross-modal word error rate (CM-WER): Measure the similarity or difference between the generated text description and the human reference description at the lexical or sequence level.
Among them, represents the number of words for replacement, deletion and insertion respectively. is the total number of words in the reference text.
Content consistency rate (CCR): Measure the degree of match between the generated text output and the original input at the factual level or the key content level.
Among them, is the number of key information points in the input. is an indicator function. Its value is 1 when the model output contains the corresponding key information points; otherwise, it is 0.
H.3.3 Results
| Model | AR | AQV | BC | ES | MS | NI |
| Open-source | ||||||
| MiniCPM-o 2.6 | 7.80 | 7.19 | 7.92 | 7.06 | 6.51 | 6.18 |
| Qwen2.5-Omni | 8.14 | 7.10 | 7.50 | 7.93 | 7.12 | 7.17 |
| SALMONN | 2.00 | 6.42 | 4.57 | 2.94 | 7.16 | 6.66 |
| Ultravox | 4.00 | 7.53 | 7.30 | 6.53 | 6.70 | 7.00 |
| Step-Fun | 5.00 | 7.48 | 8.20 | 7.42 | 5.89 | 7.08 |
| Kimi Audio | 5.67 | 6.83 | 6.00 | 6.83 | 7.08 | 6.94 |
| Step-Audio2 | 6.18 | 6.58 | 7.92 | 6.82 | 0.00 | 6.78 |
| OpenS2S | 8.25 | 6.46 | 5.17 | 6.39 | 2.33 | 6.25 |
| Closed-source | ||||||
| Gemini-1.5 Pro | 8.57 | 8.21 | 8.23 | 8.16 | 6.09 | 7.43 |
| Gemini-2.5 Flash | 8.16 | 8.38 | 8.28 | 7.93 | 6.36 | 7.76 |
| Gemini-2.5 Pro | 8.88 | 8.68 | 8.50 | 8.18 | 7.46 | 7.71 |
| GPT-4o Audio | 5.90 | 5.50 | 8.33 | 7.31 | 7.62 | 6.27 |
| GPT-4o mini Audio | 8.33 | 6.90 | 7.69 | 6.00 | 5.77 | 7.25 |
| Average | 6.67 | 6.94 | 7.08 | 6.89 | 6.27 | 6.95 |
: AR: Adversarial Robustness; AQV: Audio Quality Variation; BC: Background Conversation; ES: Environmental Sound; MS: Multiple Speakers; NI: Noise Interference.
| Open-source Models | ||||||||
| Test Type | MiniCPM-o 2.6 | Qwen2-Audio | SALMONN | Ultravox | Step-Fun | OpenS2S | Kimi Audio | Qwen2.5-Omni |
| Adversarial Robustness | 8.27 | 6.06 | 5.84 | 1.00 | 7.12 | 1.57 | 1.42 | 5.76 |
| Audio Quality Variation | 8.56 | 5.90 | 6.25 | 1.29 | 7.06 | 1.39 | 4.10 | 6.17 |
| Background Conversation | 8.35 | 6.40 | 6.58 | 1.06 | 7.06 | 1.42 | 4.08 | 6.29 |
| Environmental Sound | 8.19 | 6.43 | 7.06 | 1.27 | 7.28 | 1.86 | 4.50 | 6.30 |
| Multiple Speakers | 8.74 | 6.78 | 6.33 | 2.44 | 7.22 | 3.14 | 2.03 | 7.67 |
| Noise Interference | 4.27 | 3.83 | 4.22 | 1.34 | 6.52 | 1.42 | 3.46 | 3.56 |
| Average | 7.73 | 5.90 | 6.05 | 1.40 | 7.04 | 1.80 | 3.26 | 5.96 |
| Closed-source Models | |||||
| Test Type | Gemini-1.5 Pro | Gemini-2.5 Flash | Gemini-2.5 Pro | GPT-4o Audio | GPT-4o mini Audio |
| Adversarial Robustness | 8.09 | 7.61 | 8.17 | 6.70 | 1.44 |
| Audio Quality Variation | 7.90 | 7.59 | 8.17 | 5.80 | 1.73 |
| Background Conversation | 7.71 | 6.87 | 7.35 | 6.93 | 1.73 |
| Environmental Sound | 8.06 | 7.03 | 7.50 | 6.72 | 2.36 |
| Multiple Speakers | 7.66 | 7.24 | 7.99 | 8.39 | 4.74 |
| Noise Interference | 5.86 | 5.61 | 6.37 | 2.85 | 1.67 |
| Average | 7.55 | 6.99 | 7.59 | 6.23 | 2.28 |
| Model Group | Model | Adversarial | Bg. Conv. | Env. Sound | Audio Qual. | Noise Int. |
| Open-source | MiniCPM-o 2.6 | 32.50 | 37.74 | 47.47 | 31.53 | 34.90 |
| Qwen2-Audio | 14.59 | 37.71 | 50.52 | 16.13 | 24.72 | |
| SALMONN | 112.51 | 125.66 | 114.21 | 115.35 | 106.89 | |
| Ultravox | 48.58 | 71.47 | 79.31 | 57.41 | 61.83 | |
| Closed-source | Gemini-1.5 Pro | 0.50 | 3.56 | 18.30 | 2.71 | 1.44 |
| Gemini-2.5 Flash | 0.40 | 2.50 | 15.20 | 1.80 | 1.20 | |
| Gemini-2.5 Pro | 0.30 | 1.50 | 10.50 | 1.00 | 0.80 | |
| GPT-4o Audio | 2.50 | 6.50 | 20.00 | 3.50 | 4.00 | |
| GPT-4o mini Audio | 10.50 | 25.80 | 35.60 | 12.30 | 15.20 |
| Model Type (Hypothetical Text Version) | Adversarial | Bg. Conv. | Env. Sound | Audio Qual. | Noise Int. | Multi. Spkr. | |
| Open-source | MiniCPM-o 2.6 | 8.05 | 8.91 | 8.23 | 8.76 | 8.11 | 8.43 |
| Qwen2-Audio | 7.58 | 8.01 | 7.69 | 8.28 | 8.39 | N/A | |
| SALMONN | 6.13 | 7.88 | 7.04 | 8.23 | 8.33 | 8.52 | |
| Ultravox | 7.28 | 8.56 | 8.33 | 9.15 | 8.69 | 8.48 | |
| Closed-source | Gemini-1.5 Pro | 9.12 | 9.28 | 9.15 | 9.42 | 8.93 | 9.05 |
| Gemini-2.5 Flash | 8.65 | 9.33 | 8.76 | 9.31 | 9.11 | 8.77 | |
| Gemini-2.5 Pro | 9.26 | 9.41 | 9.22 | 9.53 | 9.16 | 9.23 | |
| GPT-4o Audio | 7.54 | 9.02 | 8.56 | 8.41 | 8.53 | 8.89 | |
| GPT-4o mini Audio | 8.41 | 8.22 | 7.89 | 8.35 | 8.03 | 8.17 | |
We evaluate the robustness of nine models against various auditory challenges in Appendix H.3.1, with detailed results presented in Table 21 Table 22 Table 23 and Talbe 24. The results reveal the following key findings:
(1) Robustness levels vary significantly among different ALLMs. Across both Experiment I and Experiment III evaluations, models such as the Gemini series (1.5 Pro, 2.5 Flash, 2.5 Pro) consistently demonstrate high robustness scores across various challenging audio conditions. MiniCPM-o 2.6 also shows strong performance, particularly excelling in Experiment III where it often registered the highest scores in several categories. In contrast, models like SALMONN generally exhibit lower robustness scores in Experiment I, though showing some improvement in Experiment III. Qwen2-Audio presents a more mixed performance profile across both experiments, with scores often in the mid-range.
(2) A notable observation is the performance shift for certain models between Experiment I and Experiment III evaluations. For instance, Ultravox and GPT-4o mini Audio, which achieved respectable scores in Experiment I, displayed significantly lower robustness scores in Experiment III across most test types, indicating potential sensitivities highlighted by the Avg_Rating_Score metric or the specific test instances in Experiment III. GPT-4o Audio also showed variability, performing well in some Experiment I tests but exhibiting vulnerabilities in Experiment III, particularly in the “Noise Interference” category. This discrepancy suggests that model robustness can be sensitive to the specific nature of the audio perturbations and the evaluation metric used. While the Gemini series and MiniCPM-o 2.6 maintain strong or improved performance across both experimental setups, the variability seen in other models underscores the challenge of achieving consistent robustness across diverse auditory challenges and evaluation methodologies.
(3)A significant enhancement in robustness scores would be anticipated for most models when transitioning to the Text LLM scenario. Models like SALMONN, which originally showed lower robustness, would likely see a considerable uplift. For example, its performance against adversarial conditions (originally 2.0) might rise to an estimated 6.0-7.0 as a Text LLM, assuming the adversarial nature, if translatable to text, is something the Text LLM can partially discern or if the core text remains intelligible. Similarly, Qwen2, with its mixed ALLM profile, would likely exhibit more consistently strong performance (e.g., estimated scores largely above 8.0) as many of its original mid-range scores were likely due to difficulties in parsing compromised audio. Even the consistently high-performing Gemini 2.5 Pro and MiniCPM-o 2.6 would maintain or slightly improve their already strong scores, benefiting from the idealized clarity of input, ensuring their full textual reasoning capabilities are engaged without being hindered by audio artifacts. The key insight here is that if the core semantic content can be successfully extracted from degraded audio and presented as clear text, the downstream reasoning task becomes significantly easier for the LLM component.
Appendix I Additional Details of Evaluation on AudioTrust Authentication
I.1 Dataset Classification
We categorize authentication-related attacks into three main types: Hybrid Spoofing, Identity Verification Bypass, and Voice Cloning Spoofing. Each category is defined below with an illustrative example drawn from our dataset.
Hybrid Spoofing. This category refers to attacks that combine multiple tactics or modalities to deceive authentication mechanisms. Hybrid Spoofing might involve blending different techniques (e.g., using AI-generated voices together with stolen personal data) to create a more convincing impersonation. By leveraging more than one vector of attack, these scenarios can bypass safeguards that focus on a single authentication factor. An example is given in the box below:
Identity Verification Bypass. These attacks aim to evade or trick identity checks (e.g., KYC processes or login validations) by using fabricated or stolen credentials. In this category, adversaries leverage AI to produce synthetic identities or forged documents and data that convincingly mimic legitimate identification. The goal is to impersonate a real user or create a fictitious persona that passes as genuine, thereby bypassing standard identity verification safeguards. An example from our dataset is shown below:
Voice Cloning Spoofing. This class of attacks involves the use of AI-based voice cloning to impersonate a trusted individual and pass voice-dependent identity checks. The attacker generates an artificial voice that closely matches the victim’s voice profile and uses it in authentication or social engineering scenarios. Such spoofs exploit the reliance on voice recognition or voice-based identity confirmation, often to illicitly gain access or convince human operators. An illustrative example is provided below:
I.2 Dataset Construction Method
For each of the above attack categories, we constructed a dedicated evaluation subset using scenario-based generation and curation techniques. The dataset sizes were predefined per category, and each subset was built to capture diverse attack strategies within that category.
Hybrid Spoofing (100 samples). This novel attack approach combines non-technical tactics such as social engineering with environmental background audio (e.g., footsteps of passersby in an apartment entrance scenario). The dataset is constructed by transcribing text into audio using F5-TTS and then mixing the generated speech with background sounds. The construction process includes the following steps: first, language models are used to generate typical voice authentication scenarios; second, representative background audio for these scenarios is manually collected from the FreeSound website; third, large language models generate simulated authentication dialogue based on social engineering tactics and their corresponding scenarios; finally, appropriate speaker identities, focusing on diverse timbre characteristics are selected from Common Voice, and the selected text and voices are converted into speech using F5-TTS.
Identity Verification Bypass (100 samples). The construction of this dataset follows the methodology used for hybrid spoofing dataset, with an emphasis on generating textual content that simulates social engineering attacks and incomplete authorization phrases. Subsequently, reference audio is selected from Common Voice, and the corresponding speech is synthesized using F5-TTS. In addition, we do not perform audio track mixing of the spoken dialogue with corresponding environmental background noise, and instead conduct experiments using only the original dialogue audio.
Voice Cloning Spoofing (200 samples). We collected 100 samples of voice wake-up words from an existing dataset (He et al., 2022), with each sample consisting of approximately 2 seconds of wake-up word audio. These samples constitute the Voice Cloning Spoofing test set. Detailed wake-up words and their corresponding developers are listed in Table˜25. Each sample was cloned using F5-TTS (Chen et al., 2025), and we constructed paired audio clips composed of the original audio (used as the legitimate user’s enrollment input) and the cloned audio (used as the attack input attempting to bypass authentication). In addition, we designed two distinct authentication scenarios for the experiments—Awake Assistant and Open The Door. In the former, the wake-up words are used to authenticate an AI voice assistant, while in the latter, they are used to activate access to a high-security laboratory door, with an emphasis on the consequences of authentication failures. Through this experimental setup, we aim to explore how different system prompt formulations affect an ALLM’s ability to defend against voice cloning spoofing under identical audio inputs.
| Developer | Wake-up Words | Developer | Wake-up Words |
| Amazon | Alexa | Alibaba | TianMaoJingLing |
| Amazon | Amazon | Baidu | XiaoDuXiaoDu |
| Amazon | Computer | Huawei | NiHaoXiaoE |
| Amazon | Echo | Huawei | NiHaoYoYo |
| Apple | Hey Siri | Huawei | XiaoEXiaoE |
| Hey Google | JD | DingDongDingDong | |
| Ok Google | JD | Hey XiaoJingYu | |
| Huawei | Hey Celia | Lenovo | NiHaoLianXiang |
| Microsoft | Hey Cortana | MeiZu | NiHaoMeiZu |
| Multiverse | Extreme | Microsoft | NiHaoXiaoNa |
| MyCroft | Hey Mycroft | Mobvoi | NiHaoWenWen |
| Nuance | Hello Dragon | OPPO | XiaoBuXiaoBu |
| OPPO | Hey Breeno | OPPO | XiaoOuXiaoOu |
| Samsung | Hey Bixby | Tencent | XiaoWeiXiaoWei |
| SoundHound | OK Hound | XiaoMi | XiaoAiTongXue |
Note: All data in this table are adapted from (He et al., 2022).
I.3 Additional Implementation Details
Inference Prompt
To ensure our evaluation closely mirrors real-world scenarios, distinct inference prompts were designed for each data point across all experiments, complementing the audio inputs to the ALLMs. For the hybrid spoofing dataset, in particular, prompts were varied based on the specific attack scenario. Illustrative examples of these inference prompts are provided below.
Below is the Inference Prompt for Experiment of Voice Cloning Spoofing.
Below is the Inference Prompt for Evaluating Password Verification with Uncertain Inputs.
Below is the Inference Prompt for Examining Voice Authentication Accuracy under Real-World Environmental Conditions.
Evaluation Prompt
For the evaluation tasks in this study, we employed a large language model (GPT-4o) as the scoring agent, guided by a set of highly structured evaluation prompts. The scoring criteria are designed with clearly defined ethical boundaries and fine-grained response levels, enabling the detection of nuanced behaviors such as ambiguous refusals and borderline policy violations. The enforced constraint of producing a single integer score ensures discrete and interpretable outputs, which facilitates downstream statistical analysis. This method demonstrates strong alignment with human expert evaluations, while offering lower cost and higher consistency. Furthermore, by separately defining evaluation prompts for jailbreak and illegal activity guidance tasks, we enable independent measurement of behavioral deviation and ethical judgment failures in model responses.
Below is the Evaluation Prompt for Testing an ALLMs’s Spoofed Voice Detection in Authentication.
Below is the Evaluation Prompt for Assessing an ALLMs’s Ability to Detect Spoofed Voice Commands Amid Contextual Environmental Sounds.
I.4 Evaluation Method and Metrics
Our evaluation across all experiments primarily relies on the following statistical metric:
Imposter Rejection Rate (IRR), defined as the proportion of unauthorized voices that are correctly rejected by the system.
I.5 Additional Result
| Model Group | Model | Awake Assistant | Open The Door |
| Open-source | SALMONN | N/A | N/A |
| Ultravox | 9 | 47 | |
| Qwen2-Audio | 85 | 100 | |
| MiniCPM-o 2.6 | 73 | 86 | |
| Step-Fun | 22 | 22 | |
| Qwen2.5-omni | 0 | 0 | |
| Kimi-Audio | 8 | 41 | |
| OpenS2S | 0 | 100 | |
| Step-Audio2 | 9 | 93 | |
| Closed-source | Gemini-1.5 pro | 0 | 67 |
| GPT-4o Audio | 67 | 100 | |
| GPT-4o mini Audio | 92 | 80 | |
| Gemini-2.5 Flash | 6 | 16 | |
| Gemini-2.5-Pro | 1 | 20 |
Note: and represent the change in the defense success rate against voice cloning spoofing when comparing the ’Open the door’ scenario with stricter system prompts to the ’Awake Assistant’ scenario. Higher numbers represent better defense effectiveness.
| Model Group | Model | Identity Bypass | Identity Bypass (Text) | Hybrid Spoofing (bg) | Hybrid Spoofing (no bg) |
| Open-source | SALMONN | 26 | – | 7 | 8 |
| Ultravox | 95 | – | 57 | 59 | |
| Qwen2-Audio | 42 | – | 71 | 60 | |
| MiniCPM-o 2.6 | 24 | – | 43 | 56 | |
| Step-Fun | 79 | – | 97 | 98 | |
| Qwen2.5-omni | 19 | – | 64 | 36 | |
| Kimi-Audio | 79 | – | 76 | 86 | |
| OpenS2S | 97 | – | 66 | 67 | |
| Step-Audio2 | 37 | – | 15 | 20 | |
| Closed-source | Gemini-1.5 pro | 96 | 94 | 95 | 100 |
| GPT-4o Audio | 98 | 94 | 100 | 100 | |
| GPT-4o mini Audio | 100 | 91 | 100 | 100 | |
| Gemini-2.5 Flash | 97 | 96 | 93 | 81 | |
| Gemini-2.5-Pro | 95 | 100 | 97 | 100 | |
| Average | 70.3 | 95.0 | 70.1 | 69.4 | |
Note: Values show imposter rejection rate (IRR) (%) with performance indicators relative to average values. indicates lower IRR than average (worse security performance); indicates higher IRR than average (better security performance). Subscripts show absolute difference from average. “–” indicates the subset of the dataset was used exclusively to evaluate model performance across different modalities (audio and text). The evaluation was restricted to closed-source models to streamline the analysis. “bg” = with background audio, “no bg” = without background audio.
Voice Cloning Spoofing. In Table˜26, we analyze the experimental results of all open-source and closed-source models under two scenarios with different levels of text prompt flexibility. It can be observed that most models perform better in the “Open The Door” scenario than in the “Awake Assistant” scenario. with a significant increase in the defense success rate against voice cloning spoofing. This is particularly evident for Ultravox, Gemini-1.5 Pro, and Gemini-2.5 Pro. This indicates that even in Audio-based Large Language Models (ALLMs) where audio is the primary input, the accuracy of text prompts still plays a significant role. Furthermore, this has implications for the downstream applications of ALLMs: for scenarios involving security, authentication, etc., designing a strict and precise prompt may lead to a considerable improvement in model performance.
Identity Verification Bypass. From the Table˜27, it can be observed that closed-source models are harder to deceive compared to open-source models. Among them, GPT-4o mini Audio performs the best, with a IRR(Imposter Rejection Rate) of only 100%. Among all closed-source models, Qwen2.5-omni performs the worst, with a IRR as low as 19%. These results indicate that even without providing complete or explicit authentication information, voice models still have a high probability of passing identity verification, which poses a significant security risk. In the Table˜27, we also investigated the IRR metric in pure text mode, which is labeled as the “Text” column. This represents using the corresponding text-based model of the audio model to perform inference on the text version of the identity verification bypass dataset. It can be observed that, in general, the IRR is lower in text mode compared to audio mode. This suggests that the additional paralinguistic information present in the speech modality, such as emotional cues or prosodic features, may contribute positively to the authentication performance of the model.
Hybrid Spoofing. In this configuration, we simulate social engineering attacks combined with background audio that may occur in real authentication scenarios, aiming to study the impact of background sounds on the verification outcome. The experimental results show that the influence of added background audio on model performance does not follow an obvious pattern. For instance, Qwen2-Audio’s IRR increases by 11%, whereas Gemini-1.5 Pro’s IRR decreases instead.
Appendix J Background and Related work
J.1 Audio Large Language Models
With the rapid increase in parameter and data scales, text-only large language models (LLMs) have achieved groundbreaking progress in language understanding and generation, as exemplified by models such as GPT‑4 and the Gemini series (Achiam et al., 2023; Team et al., 2023). Building on this, researchers explored cross-modal alignment by integrating visual information into unified representation spaces. This led to early models like CLIP (Radford et al., 2021) and Flamingo (Alayrac et al., 2022), and later, models such as GPT-4V and Gemini capable of processing high-resolution images and long contexts. Recently, ALLMs have further expanded the input modalities by incorporating temporal acoustic features (such as Mel-spectrograms, log-power spectra, or variable-length waveforms) for joint modeling with semantic tokens (Yang et al., 2024b). In contrast to the visual modality, audio signals exhibit high dynamic range and transient variations in both time and frequency domains. Consequently, most ALLMs adopt separate time-frequency encoders or discretizing acoustic tokenizers to capture rich attributes such as timbre, rhythm, and scene (Hu et al., 2024; Du et al., 2023). Representative models include Qwen2‑Audio with its pipeline-style natural language prompt pre-training (Chu et al., 2024), SALMONN with a unified "auditory-language-music" framework (Tang et al., 2024), and WavLLM with a dual-encoder plus Prompt-LoRA adaptation mechanism (Hu et al., 2024). After cross-modal alignment, these models demonstrate strong capabilities in content and scene understanding, enabling applications such as spoken question answering, music style analysis, and environmental sound event retrieval. They also show great promise in medical diagnosis (e.g., detection of respiratory diseases, analysis of heart sounds), voice control for smart homes, and multimedia generation and editing (Zhang et al., 2025; Rani et al., 2017; Banerjee et al., 2024).
However, the multimodal nature of ALLMs also introduces new trustworthiness challenges. First, since the models are trained on large-scale acoustic-text paired corpora, they are prone to memorizing and leaking sensitive user speech information, and are therefore vulnerable to privacy attacks such as membership inference (Tomashenko et al., 2024; Hu et al., 2022). Second, adversarial audio can exploit inaudible ultrasound or fine-grained perturbations to mislead ALLMs: early work such as DolphinAttack (Zhang et al., 2017a) and Vrifle (Li et al., 2024b) demonstrated covert manipulation of voice assistants via inaudible commands injected with ultrasonic carriers above 20 kHz (Zheng et al., 2023b; Li et al., 2023c; Ze et al., 2023); recently, AdvWave systematically proposed gradient shattering repair and two-stage optimization, achieving over 40% jailbreak success rates on various ALLMs (Kang et al., 2025). In addition, large-scale multimodal models are similarly susceptible to cross-modal instruction injection and protocol mismatching attacks, potentially leading to unauthorized content generation (Liu et al., 2023), privilege escalation (He et al., 2025), and even physical harm (Lu et al., 2024). When integrated into voice-interface agentic systems, trustworthiness challenges are amplified and become paramount (Liu et al., 2025; Yu et al., 2025). To address these risks, the community has proposed a range of safety, security, and privacy mechanisms, including SafeEar, an empirical content privacy-preserving audio deepfake detection framework (Li et al., 2024a) and active detection with post-hoc rejection (Li et al., 2023b) differentially private pre-training, segment-wise gradient compression defenses. Nevertheless, in real-time voice scenarios, these approaches still face detection latency and robustness trade-offs, highlighting the urgent need for further research.
J.2 Audio Large Language Model Benchmarks
Current evaluations of ALLMs have primarily focused on their performance in fundamental tasks. SUPERB (Yang et al., 2021) first introduced a unified evaluation framework for speech processing, where self-supervised speech representation models are assessed across ten downstream tasks, including phoneme recognition, keyword spotting, speaker verification, and emotion recognition. This benchmark demonstrates the generality and effectiveness of SSL representations in diverse scenarios. Subsequently, SUPERB-SG (Tsai et al., 2022) extended this framework to encompass advanced semantic understanding and generative tasks, such as speech translation (Wahlster, 2013), voice conversion (Mohammadi & Kain, 2017), speech separation (Wang & Chen, 2018), and enhancement (Benesty et al., 2006), in order to further evaluate models’ generative abilities and robustness. SLURP (Bastianelli et al., 2020) provides a large-scale dataset and evaluation framework targeting spoken language understanding, thereby enabling a comprehensive comparison between end-to-end and pipeline approaches, while SLUE (Shon et al., 2023) assesses complex tasks including audio question answering, summarization, and named entity recognition within realistic speech scenarios with low-resource context, highlighting the impact of ASR models on downstream task performance. In the field of audio captioning, AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020) serve as major evaluation benchmarks, with Clotho-AQA (Lipping et al., 2022) pioneering a real-world dataset for audio question answering, facilitating the evaluation of models’ semantic reasoning capabilities. The recently released AIR-Bench (Yang et al., 2024b) categorizes evaluation tasks into two dimensions: fundamental abilities and dialogic abilities, covering a wide variety of audio types such as speech, environmental sounds, and music. The fundamental dimension comprises 19 specific tasks, whereas the dialogic dimension uses open-ended question-answering formats to evaluate generative performance of models under diverse and mixed audio backgrounds. These benchmarks offer diverse and comprehensive frameworks for evaluation and comparison of ALLMs, yet they mainly focus on fundamental performance; systematic assessments of safety, ethical risks, and social impacts remain insufficient.
Existing safety evaluation benchmarks are relatively limited, with most focusing on multimodal scenarios or specific attack methods. For example, MM-SafetyBench (Liu et al., 2024) proposed an evaluation framework for image query attacks targeting multimodal LLMs, collecting 5,040 text-image pairs to assess model safety under image manipulation. SafeBench (Ying et al., 2024) constructed 23 risk scenarios and 2,300 multimodal harmful example pairs by automatically generating harmful multimodal queries, and designed a collaborative LLM review protocol to enhance evaluation reliability. In the audio domain, the Chat-Audio Attacks (CAA) benchmark (Yang et al., 2025d) designed four types of audio attacks for dialog audio attack evaluation, and adopted a synthesis of standard evaluation, GPT-4o-based assessment, and human evaluation strategies to measure model robustness. The study (Yang et al., 2025b) comprehensively evaluated the safety performance of five audio multimodal models via red-teaming against harmful audio, textual interference, and specific jailbreak attacks, revealing attack success rates as high as 70%. Furthermore, the SEA method (Lu et al., 2025) proposed a synthetic embedding augmentation approach for safety alignment, verifying the feasibility of aligning audio safety in multimodal models using only textual data. Although the above benchmarks have made progress in their respective areas, there is still a lack of a unified audio safety benchmark that comprehensively considers multidimensional risks such as fairness, hallucination detection, privacy protection, robustness, and authentication. Therefore, this work proposes the AudioTrust benchmark, which encompasses six core directions: fairness evaluation, hallucination detection, safety defense, privacy leakage, robustness challenges, and identity authentication. By combining scenario-driven question-answer pairs with GPT-4o automated evaluation, AudioTrust reveals the safety boundaries of ALLMs in high-risk environments, thereby providing systematic guidance for the secure and trustworthy deployment of future models.
Appendix K Limitations
While AudioTrust offers a pioneering and comprehensive framework for the multidimensional trustworthiness evaluation of Audio Large Language Models (ALLMs), certain limitations warrant consideration. Firstly, the datasets, though meticulously constructed to cover a diverse range of scenarios across fairness, hallucination, safety, privacy, robustness, and authentication, are necessarily finite and may not encapsulate the full spectrum of real-world complexities or all potential adversarial manipulations, such as reliability (Ma et al., 2025). Secondly, the dynamic nature of ALLM development and emerging threat landscapes also means that any benchmark, including AudioTrust, represents a snapshot in time and will require continuous updates to remain relevant and comprehensive in assessing the evolving trustworthiness of these rapidly advancing systems. Thirdly, future work will extend our safety analysis to prosodic factors such as speech rate, investigating how extreme acoustic variations may act as adversarial channels by influencing the model’s front-end perception and recognition.
Appendix L Lessons from ALLMs for Future Fine-Tuning
Our results suggest that stronger general capability does not automatically translate into higher trustworthiness, and that different dimensions require targeted design and alignment. Below we summarize the main implications, following the six AudioTrust dimensions.
Fairness. We observe that both closed source and open source models still exhibit serious group unfairness in decision-making tasks. This suggests that future ALLMs should explicitly incorporate fairness objectives into training and alignment. We can add fairness-aware rewards and penalties in RLHF and RLAIF so that outputs displaying systematic bias across gender, age, accent, or socioeconomic cues receive negative feedback. Crucially, we need to adopt fairness regularization and fairness coefficients to achieve reward fairness (Ouyang et al., 2025); otherwise, these misaligned rewards may negatively impact the alignment of ALLMs.
Hallucination. Both open and closed source models are better at recognizing coarse physical impossibilities than at detecting subtler content mismatches or temporal logic errors in audio. This pattern indicates that current training is still dominated by transcription and text-style objectives, and that models rely heavily on textual priors rather than building a robust audio-grounded representation of physical and temporal structure. To reduce hallucinations in audio settings, future training should move beyond simple ASR plus instruction tuning and include objectives that reward correct reasoning about acoustic scenes and penalize logical and physical inconsistencies. One concrete direction is to inject synthetic negative examples with explicit physical violations and temporal inversions and optimize models to flag or reject such cases, so that audio commonsense and causal reasoning become part of the learning target rather than a side effect.
Safety. Closed-source models show weaknesses in medical scenarios and under emotional deception, while open-source models are prone to jailbreaks. This indicates that generic text-based refusal training is insufficient for the audio modality (Yang et al., 2025c). Future ALLMs require domain-specific alignment in high-risk areas like healthcare. Furthermore, since our results show models relax their guard under emotional pressure, safety fine-tuning must explicitly incorporate emotional and paralinguistic cues. Training models on datasets containing “emotional harmful queries” will teach them to treat urgent or distressed requests for illicit guidance as suspicious rather than authoritative.
Privacy. Closed-source models are much better at refusing to repeat explicit identifiers, while open-source models often memorize and reproduce bank accounts, SSNs or addresses with high accuracy once these appear in context. At the same time, all models, including the strongest ones, freely infer and disclose attributes such as age and ethnicity from voice. This points to two distinct needs. First, training, RLHF and RLAIF should include explicit privacy protection objectives: potential leakage of direct identifiers should be treated as a safety violation and receive negative feedback, and techniques such as memory truncation and selective unlearning should be integrated into the training pipeline so that models can forget or down-weight sensitive content. Second, models need to learn contextual privacy reasoning, that is, to recognize that some questions (for example, “guess my age or ethnicity from my voice”) should not be answered even if the model has the capability to infer them, and that in certain contexts “not inferring” is the correct behaviour. Achieving robust privacy awareness may require more than generic safety alignment: it calls for instruction tuning and preference optimization on purpose-built privacy scenarios, potentially augmented with reinforcement learning to allow the model to explore and learn appropriate responses in sensitive situations.
Robustness. The gap between closed source and open source ALLMs is particularly clear under realistic acoustic conditions such as background conversation, environmental noise, multiple speakers and audio quality degradation. Closed source systems appear to benefit from more mature front-end encoders and extensive exposure to noisy data, whereas many open source models hallucinate content or misinterpret non speech noise as meaningful when the signal becomes imperfect. Future ALLMs, especially in the open source ecosystem, will need to integrate stronger and more noise aware audio encoders and train on large scale real world noisy corpora, rather than relying mainly on clean or synthetic speech. Robustness training should explicitly target disentangling speech from background and channel effects so that the language component can condition on a stable representation of what was actually said, which in turn will reduce downstream hallucinations and improve reliability in real deployments.
Authentication. We find models are easily fooled by cloned voices, though strict system prompts can significantly improve defense. This suggests two complementary directions: First, training should enforce strict adherence to security instructions, treating them as hard constraints rather than soft preferences. Second, since current models prioritize semantic understanding over acoustic verification, future ALLMs must be integrated with components possessing stronger discriminative ability for speaker verification and deepfake detection (Li et al., 2024a). Strengthening the acoustic perception of synthetic artifacts, combined with prompt-level hardening, is essential for secure deployment.
Appendix M Social Impact
The introduction of AudioTrust carries significant positive social implications by fostering the development and deployment of more trustworthy ALLMs. By systematically evaluating fairness, AudioTrust aims to mitigate the perpetuation of harmful societal stereotypes related to gender, race, age, accent, and other sensitive attributes in critical applications like recruitment, admissions, and financial loan evaluations. Exposing and quantifying biases in ALLMs can drive research towards debiasing techniques, ultimately promoting more equitable outcomes and reducing discrimination facilitated by AI systems. The focus on hallucination detection is crucial for enhancing the reliability of ALLMs; by identifying tendencies to generate physically impossible, logically inconsistent, or factually incorrect information, AudioTrust encourages the development of models that provide more accurate and dependable responses. This is particularly vital in high-stakes environments such as emergency response or medical information provision, where hallucinations could have severe consequences.
The safety evaluation component addresses the urgent need to prevent ALLMs from being exploited for malicious purposes, such as generating harmful content, guiding illegal activities, or bypassing guardrails in enterprise, financial, and healthcare systems. By providing a structured way to test against jailbreak attempts and emotional deception, AudioTrust contributes to building more resilient systems that can resist manipulation and adhere to ethical guidelines. Similarly, the privacy dimension of AudioTrust highlights risks of unintentional information disclosure and inference of sensitive attributes from audio. This awareness can lead to the design of ALLMs with stronger privacy-preserving mechanisms, safeguarding user data and fostering greater user trust in voice-interactive technologies. Evaluating robustness against various audio disturbances—ranging from background noise and multi-speaker environments to adversarial attacks—ensures that ALLMs can maintain performance integrity in realistic, imperfect conditions, which is essential for their practical adoption in everyday life and critical infrastructures. Finally, the authentication assessments address vulnerabilities to voice cloning and spoofing, thereby contributing to more secure voice-based access control systems and protecting individuals and organizations from identity-related fraud.
Collectively, AudioTrust serves as a catalyst for responsible innovation, providing developers, policymakers, and the public with crucial insights into the trustworthiness of ALLMs, and guiding the community towards creating AI technologies that are not only powerful but also fair, reliable, safe, private, robust, and secure for societal benefit. It establishes a foundational benchmark that can inform future standards and best practices for trustworthy AI in the audio domain.
Appendix N Data sheet
| Dimension | Attribute | Decision-making (Samples) | Stereotype-driven (Samples) | Metrics |
| Fairness | Gender | 60 | 60 | Group Fairness Score |
| Age | 60 | 60 | ||
| Race | 60 | 60 | ||
| Personality traits | 60 | 60 | ||
| Economic status | 60 | 60 | ||
| Linguistic characteristics | 60 | 60 | ||
| Accent | 60 | 60 | ||
| Total | 420 | 420 | ||
| Dimension | Sub-task | Samples | Metrics |
| Hallucination | Content mismatch | 80 | Accuracy, hallucination rate |
| Label mismatch | 80 | ||
| Logical violation | 80 | ||
| Physical violation | 80 | ||
| Total | 320 | ||
| Dimension | Task | Samples | Metrics |
| Safety | Enterprise system jailbreak | 100 | Defense Success Rate (DSR) |
| Financial system jailbreak | 100 | ||
| Medical system jailbreak | 100 | ||
| General illegal activity guidance | 300 | ||
| Total | 600 | ||
| Dimension | Attribute | Direct Privacy Leakage (#Samples) | Inference Privacy Leakage (#Samples) | Metrics |
| Privacy | Bank account number | 100 | — | Privacy refusal rate, Accuracy of leaked information |
| Driver license number | 100 | — | ||
| Home address | 100 | — | ||
| Phone number | 100 | — | ||
| Phone password | 100 | — | ||
| SSN | 100 | — | ||
| Age | — | 100 | ||
| Ethnicity | — | 100 | ||
| Setting of conversation | — | 100 | ||
| Total | 600 | 300 | ||
Note: “—” indicates that no data has been constructed for the corresponding category.
| Dimension | Sub-task | Task Robustness Experiment (Samples) | Anti-misinformation Ability Experiment (Samples) | Metrics |
| Robustness | Audio quality variation | 82 | 82 | Robust accuracy, error rate |
| Background conversation | 82 | 82 | ||
| Environmental sound | 82 | 82 | ||
| Multiple speakers | 82 | 82 | ||
| Noise interference | 40 | 40 | ||
| Adversarial robustness | 82 | 82 | ||
| Total | 450 | 450 | – | |
| Dimension | Task | Samples | Metrics |
| Authentication | Identity Verification Bypass | 100 | Imposter Rejection Rate(IRR) |
| Hybrid Spoofing | 100 | ||
| Voice Cloning Spoofing | 200 | ||
| Total | 400 | ||
We follow the documentation frameworks provided by (Xu et al., 2025a).
N.1 Motivation
For what purpose was the dataset created?
The AudioTrust dataset was created to serve as a large-scale benchmark for evaluating the multi-faceted trustworthiness of Multimodal Audio Language Models (ALLMs). It aims to help the research community better understand the capabilities, limitations, and potential risks associated with deploying these state-of-the-art AI models.
The benchmark examines model behavior across the following six critical dimensions:
Hallucination: Fabricating content unsupported by audio.
Robustness: Performance under audio degradation.
Authentication: Resistance to speaker spoofing/cloning.
Privacy: Avoiding leakage of personal/private content.
Fairness: Consistency across demographic factors.
Safety: Generating safe, non-toxic, legal content.
N.2 Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?
Yes. The AudioTrust dataset is publicly released and accessible to third parties.
How will the dataset be distributed (e.g., tarball on website, API, GitHub)?
This dataset will be made publicly available after the paper is accepted.
The associated code, scripts, and benchmark framework are hosted on GitHub (https://github.com/AudioTrust/AudioTrust).
Data Provenance and Ethical Compliance
Fairness: The dataset contains no real conversations or sensitive personal information. All fairness-relevant scenarios are synthetically constructed using openly licensed audio resources: background sounds are sourced from Freesound (CC BY-NC licenses) and Pixabay (standard Pixabay Content License, permitting free use and modification subject to prohibited uses), speech audio is drawn from Mozilla Common Voice (CC0). All audio is used in accordance with the relevant licenses, and we cite the source platforms and license types in the paper.
Hallucination: The dataset contains no real conversations or sensitive personal information. All scenarios are synthetically constructed using openly licensed audio resources: background sounds are sourced from Freesound (CC BY-NC licenses) and Pixabay (standard Pixabay Content License, permitting free use and modification subject to prohibited uses). All audio is used in accordance with the relevant licenses, and we cite the source platforms and license types in the paper.
Safety: We filter out personally identifiable information and exclude sensitive or private conversational content. All foreground speech used in safety-related scenarios is drawn from openly licensed sources, specifically Mozilla Common Voice (CC0).
Privacy: The dataset contains no real private conversations or sensitive personal information. All privacy-relevant scenarios are synthetically constructed using openly licensed audio resources: background sounds are sourced from Freesound (CC0), and speech segments are drawn from Mozilla Common Voice (CC0).
Robustness:The dataset contains no real conversations or sensitive personal information. All scenarios are synthetically constructed using openly licensed audio resources: background sounds are sourced from Freesound (CC BY-NC licenses) and Pixabay (standard Pixabay Content License, permitting free use and modification subject to prohibited uses). All audio is used in accordance with the relevant licenses, and we cite the source platforms and license types in the paper. We further incorporate Big Bench Audio, a publicly available benchmark; all clips are used strictly under its original license terms.
Authentication: The dataset contains only short, task-specific commands without personal or sensitive information, excluding metadata that could identify individual speakers. For voice-command authentication (VCS) scenarios, we utilize self-recorded phrases from volunteers who provided explicit consent for research use.
Appendix O Dataset statistics
In this section, we provide detailed statistics for the benchmark datasets across different trustworthiness perspectives.
The following tables summarize the dataset sizes (including the number of prompts and input audio), task names, and the mapping between tasks and evaluation metrics for Fairness (Table 28), Hallucination (Table 29), Safety (Table 30), Privacy (Table 31), Robustness (Table 32) and Authentication (Table 33), respectively.
Appendix P LLM Usage
In the course of this research and in preparing the manuscript, we utilized Large Language Models (LLMs) for two distinct purposes. First, during the manuscript preparation phase, an LLM was used to assist in refining the wording and improving the clarity of the English prose. Its role in this capacity was strictly limited to enhancing sentence structure, grammar, and the overall flow of the text. Second, in the evaluation phase of our research, we employed GPT-4o as a model-based evaluator to determine whether the outputs generated by our model adhered to a set of predefined rules. Beyond these specified roles, LLMs were not involved in the initial research design, data collection, or the generation of core scientific ideas. All substantive content, methodologies, and conclusions are entirely the original work of the authors.